### v1

# VCB Bench: An Evaluation Benchmark for Audio-Grounded Large Language Model Conversational Agents

Abstract
Recent advances in large audio language models (LALMs) have greatly enhanced multimodal conversational systems. However, existing benchmarks remain limited—they are mainly English-centric, rely on synthetic speech, and lack comprehensive, discriminative evaluation across multiple dimensions. To address these gaps, we present Voice Chat Bot Bench (VCB Bench)—a high-quality Chinese benchmark built entirely on real human speech. VCB Bench evaluates LALMs from three complementary perspectives: instruction following (including speech-level control beyond text commands), knowledge understanding (general knowledge, reasoning, and daily dialogue), and robustness (stability under perturbations in content, environment, and speaker traits). Experiments on representative LALMs reveal notable performance gaps and highlight future directions for improvement. VCB Bench provides a reproducible and fine-grained evaluation framework, offering standardized methodology and practical insights for advancing Chinese voice conversational models.  [^1] Code and data are available at https://github.com/193746/VCB-Bench-Evalkit
VCB Bench: An Evaluation Benchmark for Audio-Grounded Large Language Model Conversational Agents
Jiliang Hu1,2, Wenfu Wang1, [^1] Corresponding authors., Zuchao Li2, [^1] Corresponding authors., Chenxing Li1, Yiyang Zhao1 Hanzhao Li1, Liqiang Zhang1, Meng Yu1, Dong Yu1 1Tencent AI Lab, Beijing, China, 2Wuhan University, Wuhan, China
1 Introduction
In recent years, large language models (LLMs) (Vaswani et al., 2017; Anil et al., 2023) have achieved remarkable progress in natural language understanding and generation. Integrating language modeling with modalities such as vision and audio (Radford et al., 2021; Singh et al., 2022) has further given rise to a new paradigm of multimodal learning. Within this trend, large audio language models (LALMs)—which combine speech signal processing with language modeling—have developed rapidly. Emerging systems such as StepAudio2 (Wu et al., 2025) and Qwen3-Omni (Xu et al., 2025b) demonstrate end-to-end (E2E) speech understanding and generation with capabilities in voice question answering, real-time conversation, and audio content analysis. Consequently, voice conversational agents powered by LALMs are drawing increasing academic and industrial attention, offering more natural and human-like interactions than text-only systems.
Despite these advances, moving from basic LALM functionalities to practical voice agents requires reliable and comprehensive evaluation tools. Such benchmarks are essential for diagnosing model weaknesses, guiding optimization, and enabling fair comparisons across systems. While initial efforts (Chen et al., 2024; Yang et al., 2024; Lin et al., 2025) have explored instruction following, audio understanding, reasoning, and dialogue scenarios, current evaluation practices remain limited in three major ways. First, most benchmarks are English-centric, leaving Chinese—the world’s most widely spoken language—largely unexplored. Second, the majority rely on synthetic speech data, which poorly reflects real-world acoustic variability. Third, many are text-derived benchmarks (e.g., AlpacaEval (Li et al., 2023), IFEval (Zhou et al., 2023)), whose formal and lengthy content is unsuitable for evaluating conversationally grounded LALMs that should generate natural, colloquial speech. Addressing these limitations is critical given China’s large user base and the growing demand for practical, high-quality voice agents.
To bridge these gaps, we introduce Voice Chat Bot Bench (VCB Bench)—the first comprehensive evaluation framework for Chinese voice conversation, built entirely from authentic (non-synthetic) speech. VCB Bench evaluates LALMs along three complementary dimensions: (1) Instruction following, extending beyond text-based prompts to incorporate speech-level control tasks such as adjusting volume, speed, and emotion, with bilingual (Chinese-English) support; (2) Knowledge, including multi-disciplinary general knowledge (12 subjects), mathematical and logical reasoning, daily dialogue comprehension, and story continuation for pretraining performance assessment; (3) Robustness, measuring model stability under real-world perturbations across content (mispronunciations, grammatical errors), environment (street, TV noise), and speaker characteristics (age, accents).
Our proposed VCB Bench is built entirely from authentic human recordings rather than synthetic speech. It provides a large-scale, high-fidelity dataset covering diverse conversational scenarios and introduces a multi-dimensional evaluation framework that jointly measures knowledge understanding, instruction following, and robustness through fine-grained, reproducible tasks. Based on this benchmark, we conduct a systematic empirical analysis of state-of-the-art LALMs under unified settings, revealing their strengths and limitations in Chinese voice interaction and offering actionable insights for future model development.
2 Related Work
Large Audio Language Models. Recent LALMs primarily adopt an E2E audio-language modeling paradigm, integrating speech understanding and generation within a unified framework.
Qwen-Audio and Qwen-Omni series (Chu et al., 2023, 2024; Xu et al., 2025a, b) progressively enhance cross-modal alignment and modeling efficiency. Qwen-Audio establishes robust audio-text alignment, Qwen-Audio2 improves encoding efficiency via multi-scale feature fusion, and the latest Qwen-Omni models introduce dual-core Thinker-Talker architectures and multi-codebook pretraining, achieving low-latency bilingual dialogue.
StepAudio models (Huang et al., 2025; Wu et al., 2025) focus on tightly coupling recognition and synthesis. StepAudio integrates a dual-codebook tokenizer and achieves a remarkably low WER, while StepAudio2 advances to a fully E2E design with fixed text-speech token alignment and Chain-of-Thought reasoning, improving fine-grained paralinguistic understanding.
Baichuan-Audio (Li et al., 2025) employs hierarchical RVQ discretization and dual audio heads to balance acoustic and linguistic objectives, enabling real-time bilingual communication. GLM4-Voice (Zeng et al., 2024) introduces a three-module structure (Tokenizer-Backbone-Decoder) supporting emotion and dialect modeling. Kimi-Audio (Ding et al., 2025) fuses continuous acoustic and discrete semantic tokens in a dual-head architecture, achieving low-latency, high-fidelity streaming generation.
These models demonstrate rapid progress in unified audio-language modeling—covering tokenization, multimodal fusion, and real-time dialogue—but systematic benchmarks, especially for Chinese real-speech interaction, remain scarce. Current evaluations are mostly qualitative or based on synthetic data, underscoring the need for comprehensive real-speech benchmarks like VCB Bench.
Audio Benchmarks. Recent efforts have introduced several benchmarks to evaluate LALMs from different perspectives. VoiceBench (Chen et al., 2024) assesses general knowledge, instruction adherence, safety, and robustness, mainly based on existing text datasets such as AlpacaEval and SD-QA (Faisal et al., 2021), with lengthy or highly complex samples removed. OpenAudioBench, released alongside Baichuan-Audio, integrates question-answering datasets including Spoken LLaMA Questions (Nachmani et al., 2023) and Web Questions (Berant et al., 2013), and augments them with a TTS-generated reasoning subset.
AIR Bench (Yang et al., 2024) contains two components—a basic benchmark covering emotion recognition, ASR, and age estimation, and a dialogue benchmark evaluating auditory understanding and internal knowledge. MMAU (Kumar et al., 2025) and MMAR (Ma et al., 2025) focus on deep audio reasoning, requiring multi-step inference grounded in internal audio knowledge. OmniBench (Li et al., 2024) targets omni-modal models handling audio, images, and text, where text queries are paired with multimodal contexts (speech, music, or sound) to test integrated reasoning. Finally, URO Bench (Yan et al., 2025) provides a bilingual (English-Chinese) comprehensive set that evaluates audio understanding, reasoning, and conversational ability—but the speech data are entirely synthetic (TTS-generated).
Figure 1: Overview of VCB Bench.
Overall, existing benchmarks have significantly advanced the evaluation coverage of LALMs, yet they share several limitations: (1) most rely heavily on TTS or synthetic speech, (2) they focus on English, and (3) their content often derives from text-centric QA corpora rather than spontaneous human dialogue. These gaps highlight the need for a real-speech, Chinese-oriented benchmark offering multi-dimensional evaluation—the central goal of our proposed VCB Bench.
3 VCB Bench
As shown in the Figure 1, VCB Bench covers three core dimensions: Instruction Following, Knowledge, and Robustness. Instruction Following includes Text Instruction Follow (TIF) (e.g., continuation, creation), Speech Instruction Follow (SIF) (e.g., emotional, volume control), and Multi-turn Dialog (MTD) tasks. The Knowledge module assesses General Knowledge (GK) across 12 disciplines, Mathematical Logic (ML), Discourse Comprehension (DC), and Story Continuation (SC). Robustness introduces real-world perturbations from speaker variations, environmental noise, and content modifications to evaluate model stability.
3.1 Dataset Construction
The VCB Bench dataset integrates data from three distinct sources: third-party professional recordings, audio extracted from variety show Q&A segments and an internally curated two-person conversational dialogue dataset. Each source supports different evaluation modules within the benchmark.
Third-Party Recorded Data. This category supports the Instruction Following, ML, and SC tasks under the Knowledge module, as well as the Robustness module. The production pipeline involves the following steps: First, task types and examples are defined through team discussion. Next, commissioned personnel manually compose texts that fulfill the task requirements. These texts then undergo manual inspection to ensure quality. Approved texts are forwarded to a third-party recording team for professional audio production. After recording, the data team performs quality checks on the audio. Subsequently, GPT-4o-Audio is used to evaluate audio quality, while GPT-4o assesses textual quality. Finally, manual screening is conducted to select high-quality samples, determining the final evaluation dataset.
For Robustness data, it’s text materials are derived from Instruction Following module. The original audio from this module serves as the control group. To control for speaker variability, the same speaker re-recorded the text under specified interference conditions (e.g., accent, noisy environment) wherever possible, using the original audio as a baseline. For "content variation" types, the text was first modified (e.g., introducing grammatical or pronunciation errors) before being re-recorded by the same speaker. Additionally, to test performance in extreme scenarios, subsets like Volume, Speed, and Unstable Signal underwent post-processing.
Variety Show Q&A Data. This category supports GK in the Knowledge module. The process includes: crawling about 20 hours of Q&A audio; manual annotation and segmentation for timestamps; ASR transcription; quality scoring (assessing question clarity and answer accuracy) and subject classification, both via GPT-4O; selection of Q&A pairs above a score threshold; and a final manual review for transcription accuracy, answer correctness, and audio clarity.
Figure 2: Examples from the VCB Bench.
Internal Two-person Dialogue Dataset. Designed to support the DC module, this category’s data is processed as follows: the original long-form audio undergoes a two-stage segmentation with GPT-4O—first by topic, then refined into semantically coherent segments under one minute. From the transcriptions, GPT-4O generates task-specific QA pairs (e.g., for analysis or induction), followed by a final manual screening to verify question quality and answer accuracy.
3.2 Dataset Details
Instruction Following. The Instruction Following section comprehensively evaluates LALMs’ ability to understand and execute both text and speech instructions, covering three sub-tasks: TIF, SIF, and MTD. All tasks are open-ended, and both TIF and SIF support Chinese and English to meet cross-lingual evaluation needs.
TIF assesses the model’s ability to respond to textual instructions through seven sub-tasks, each examining text generation and semantic comprehension: (1) Continuation: extending a given text fragment to evaluate coherence and creativity; (2) Creation: generating original content based on a given theme to assess inventiveness and organization; (3) Empathy: understanding and responding to emotional expressions to examine affective perception; (4) Recommendation: providing suggestions based on user needs to evaluate information integration; (5) Rewriting: adapting text in style or structure to test reorganization ability; (6) Safety: identifying and rejecting harmful instructions to assess compliant response; (7) Simulation: role-playing in dialogue to examine contextual adaptation.
SIF focuses on understanding and executing speech instructions, particularly the ability to handle paralinguistic features such as emotion, speaking rate, and dialect. It includes six sub-tasks: (1) Emotional Control: adjusting the emotional tone of speech to assess expressive generation; (2) Language Control: switching languages or dialects to test multilingual synthesis; (3) Non-verbal Vocalization: incorporating non-linguistic elements like sighs or nasal sounds to evaluate paralinguistic expressiveness; (4) Pacing Control: modifying speaking rate to examine control precision; (5) Style Control: switching speech styles to assess style transfer; (6) Volume Control: adjusting loudness to test stability.
Model Instruction Following Konwledge
TIF TIF-En SIF SIF-En MTD GK ML DC
GLM4-Voice (Zeng et al., 2024) 85.82 82.52 85.57 78.52 85.13 45.53 62.14 48.64
Kimi-Audio (Ding et al., 2025) 85.13 88.92 85.69 61.87 85.67 53.51 79.94 74.76
Qwen2.5-Omni (Xu et al., 2025a) 87.40 72.58 71.37 58.09 86.93 55.43 80.24 73.72
Baichuan-Audio-Chat (Li et al., 2025) 72.49 76.22 78.96 68.07 73.27 44.48 60.33 54.38
Qwen2-Audio-Instruct (Chu et al., 2024) 84.56 75.86 / / 85.67 35.83 60.78 67.07
StepAudio (Huang et al., 2025) 87.17 66.92 80.25 63.63 / 60.42 77.07 59.52
StepAudio2Mini (Wu et al., 2025) 82.79 75.54 78.81 65.15 87.80 61.15 81.30 83.08
Mimo-Audio 91.21 91.76 72.89 24.25 / 56.58 84.01 87.92
GPT4o-Audio 91.24 91.66 88.15 86.07 / 61.29 77.68 77.64
Table 1: The main results of different LALMs on VCB Bench. Missing results from unsupported modalities or API unavailability.
MTD evaluates instruction tracking and topic management in multi-turn dialogues, each containing 3-5 turns, focusing on contextual understanding and logical coherence: (1) Progression: deepening the discussion around an initial topic to assess topic development; (2) Backtracking: recalling and responding to previously mentioned information to test long-range memory; (3) Transition: suddenly shifting to a new topic to evaluate conversational flow and relevance.
Knowledge. The Knowledge module evaluates LALMs’ knowledge storage, logical reasoning, and spoken dialog comprehension through reference-based question answering. It comprises four sub-tasks: GK, ML, DC, and SC.
GK evaluates multi-disciplinary common sense across twelve core domains—mathematics, geography, politics, chemistry, biology, law, physics, history, medicine, economics, sports, and culture—to measure the model’s ability to recall and apply knowledge across diverse fields.
ML module consists of two key components: Mathematics and Logical Reasoning. Mathematics is divided into Basic Math, which is confined to integer arithmetic within 100, and Medium Math, which includes advanced algebra, geometry, number theory, and related disciplines, collectively assessing computational and problem-solving skills. Logical Reasoning comprises four reasoning types: Analysis for breaking down information, Induction for identifying and generalizing patterns, Analogy for mapping relational correspondences, and Logic for executing conditional reasoning, thereby testing analytical and deductive capabilities.
DC focuses on understanding dialogues through three dedicated tasks: Analysis detects factual accuracy within dialogues, Induction summarizes overarching dialogue themes, and Inference deduces speakers’ attitudes, emotions, or intents, together evaluating comprehension and implicit reasoning skills.
SC, inspired by StoryCloze (Mostafazadeh et al., 2016), assesses implicit reasoning by requiring the model to select the correct story ending from two candidates, where both the context and the candidate endings are provided in the same modality—either all in audio or all in text. This task spans three evaluative categories: Logic and Causality for causal consistency, Common Sense and Science for real-world and scientific knowledge alignment, and Morality and Emotion for moral and emotional coherence.
Robustness. The Robustness module evaluates the stability of LALMs’ performance under real-world interference conditions, ensuring reliable responses in challenging scenarios. The module encompasses three dimensions: Speaker Variations (SV), Environmental Variations (EV), and Content Variations (CV).
SV examine model adaptation to speaker attributes: (1) Age: utilizes child and elderly speech to assess recognition of age-related vocal characteristics; (2) Accent: incorporates four regional accents (Tianjin, Beijing, Dongbei, Sichuan) to evaluate comprehension of non-standard Mandarin; (3) Volume: assesses perception stability with amplified/attenuated speech; (4) Speed: tests parsing capability with rapidly delivered input.
EV simulate acoustic interference: (1) Non-Vocal Noise: includes echo, outdoor, and far-field noise; (2) Vocal Noise: contains television audio, background conversations, vocal music, and radio broadcasts; (3) Unstable Signal: emulates network-induced packet loss to evaluate handling of fragmented audio.
CV introduce linguistic disruptions: (1) Fillers: incorporate discourse markers (e.g., "um", "ah"); (2) Repetition: include repeated phrases/words; (3) Mispronunciation: introduce phonetic deviations; (4) Grammatical Error: employ ungrammatical constructions; (5) Topic Shift: implement abrupt topic changes; (6) Code-Switching: mix Chinese and English. Each category evaluates the model’s ability to maintain comprehension despite content imperfections.
4 Experiment
4.1 Configuration
Figure 3: The robustness of LALMs under real-world perturbations. 9 subsets with the most significant performance gaps compared to the control group on the Robustness dataset are chosen.
We evaluate the latest and most capable LALMs. The selected models comprise GLM4-Voice, Kimi-Audio, Qwen2.5-Omni, Baichuan-Audio-Chat, Qwen2-Audio-Instruct, StepAudio, StepAudio2Mini, Mimo-Audio, and GPT4o-Audio.
For the SIF tasks in both Chinese and English (SIF-En), we invoke each model’s “audio2audio” API to generate spoken responses. The adherence to instruction requirements is then automatically scored using GPT-4o-audio. For other tasks except SC, we call the “audio2text” API to obtain textual responses, which are evaluated by GPT-4O. In open-ended question answering, GPT-4O provides a numerical score on a 1-5 scale, while for reference-based QA, it returns a binary "Yes" or "No" judgment.
For the SC task, we assess a subset of pre-trained base models: Baichuan-Audio-Base, Kimi-Audio-Base, Qwen2-Audio-Base, and StepAudio2Mini-Base. Following the StoryCloze evaluation protocol, we compute the negative log-likelihood for both the correct and incorrect endings, with model selection determined by comparing these two values. For SIF tasks, the top six performing models undergo further Mean Opinion Score (MOS) evaluation. We sample the first 30 items from each relevant dataset, and eight expert evaluators rate the generated audio samples.
In the MTD evaluation, the model receives input context only in audio. We adopt Bai et al. (2024)’s protocol, requiring the model to answer each dialogue turn using the original ground-truth context, not its prior responses. A key scoring distinction is the heightened focus on the final turn: it carries 50% of the total score per sample, and the first several turns account for the remaining 50%. All Experiments are conducted on H20.
4.2 Main Result
As shown in Table 1, GPT-4o-Audio, as a non-open-source state-of-the-art (SOTA) model, serves as a strong baseline and demonstrates all-round superiority across most tasks, which is reasonable given its non-open nature and advanced proprietary capabilities. Focusing on other open-source E2E LALMs, notable performance discrepancies emerge:
Figure 4: The investigation of the text-speech alignment capability of LALMs. A2T (Audio-to-Text, generating text directly from audio input), A2A W/ ASR (Audio-to-Audio, transcribing the generated audio via Automatic Speech Recognition to text), and A2A W/O ASR (evaluating the generated audio’s text without ASR). All test sets are from TIF/TIF-En (e.g., “Creat.” for Creation, “Recom.” for Recommendation).
In Instruction Following, Mimo-Audio excels in TIF/TIF-En with scores close to GPT-4o-Audio, indicating strong cross-lingual text instruction adaptation. For SIF, StepAudio series and Kimi-Audio perform robustly in Chinese SIF, yet their SIF-En scores lag significantly behind, reflecting challenges in handling English speech’s paralinguistic features. In MTD, StepAudio2Mini leads among open-source models with 87.80, outperforming counterparts like Baichuan-Audio-Chat, which highlights divergence in long-context dialogue logic control.
In Knowledge, Mimo-Audio stands out in ML (84.01) and DC (87.92), surpassing other open-source models and even GPT-4o-Audio’s performance—suggesting strengths in deep reasoning and text semantic analysis. However, models like Baichuan-Audio-Chat show limited performance in GK (44.48) and ML (60.33), revealing gaps in multi-disciplinary knowledge coverage and step-by-step reasoning.
Overall, open-source E2E LALMs exhibit task-specific strengths (e.g., Mimo-Audio in reasoning, StepAudio2Mini in Chinese dialogue and general knowledge) but face challenges in cross-lingual speech adaptation and comprehensive knowledge reasoning.
4.3 Real Scenario
As shown in Figure 3, EV.Echo, SV.Speed and SV.Elder cause the most severe performance degradation for most models. Scores of some models drop from over 80 in the control group to below 40 in these subsets, indicating that speech rate variation and acoustic echo are the most challenging perturbations for current LALMs. However, CV-related interferences (e.g., CV.Gram.Err, CV.Mispron) have relatively mild impacts. Some models (e.g., Mimo-Audio, StepAudio2Mini) show small score gaps between these subsets and the control group, suggesting models are more tolerant of “content-level flaws” than “speech/environment-level physical perturbations”.
Regarding model robustness, GPT-4o-Audio maintains excellent capability despite significant drops in specific subsets like SV.Speed, attributable to its high baseline scores ensuring practical usability. Among open-source models, Mimo-Audio and StepAudio2Mini exhibit relatively prominent robustness with both high absolute scores and limited performance gaps. In contrast, models like Baichuan-Audio-Chat face constraints primarily due to their lower absolute scores rather than extreme fluctuations, indicating insufficient real-scene adaptability despite moderate performance drops.
4.4 Pretraining Evaluation
Model Task Metrics
Avg. Logic Moral Common Sense
Baichuan-Audio-Base S -> T 52.36 54.41 32.65 58.33
S -> S 25.39 20.69 40.82 31.94
Kimi-Audio-Base S -> T 78.01 76.25 73.47 87.50
S -> S 54.71 49.42 69.39 63.89
Qwen2-Audio-Base S -> T 48.95 51.72 30.61 51.39
S -> S 36.91 36.78 44.90 31.94
StepAudio2Mini-Base S -> T 50.26 52.87 26.53 56.94
S -> S 30.63 27.55 34.69 38.89
Table 2: Pretraining Evaluation Results on SC.
The SC task evaluates pre-trained LALMs’ “intelligence” and cross-modal semantic coherence by judging the rationality of story endings. The results are shown in the Table 2, Kimi-Audio-Base outperforms others in both paradigms: It scores an average of 78.01 in S->T and 54.71 in S->S, with robust performance across sub-dimensions, demonstrating stable story understanding and ending judgment in cross-modal scenarios. In contrast, Baichuan-Audio-Base, Qwen2-Audio-Base, and StepAudio2Mini-Base score much lower. Moreover, all models perform worse in S->S than S->T, revealing that cross-modal (speech-to-speech) story coherence judgment remains challenging for pre-trained LALMs, with notable room for improvement in semantic consistency and rationality generation during speech output.
4.5 Ablation Study
4.5.1 Text-Speech Alignment
To investigate the text-speech alignment capability of LALMs, we conduct an ablation study, which is shown on Figure 4. The visualization is based on two selection criteria from TIF and TIF-En: the four models with the highest A2A W/ ASR scores, and the four datasets with the largest mean score differences between A2A W/ ASR and A2T. Results for Chinese and English tasks are plotted separately in the upper and lower sections of the figure, respectively.
From the results, models like GLM4-Voice (Chinese) and Baichuan-Audio-Chat (English) demonstrate strong text-speech alignment—their A2T results are close to A2A W/ ASR results, indicating consistent semantic output between directly generated text and text transcribed from speech. In contrast, models such as Qwen2.5-Omni (Chinese) and Kimi-Audio (English) show large discrepancies between A2T and A2A W/ ASR, suggesting mismatches in semantics between text and speech generation. Meanwhile, for audio generation quality (assessed by the gap between A2A W/ ASR and A2A W/O ASR, where smaller gaps imply clearer audio), GLM4-Voice (Chinese), Kimi-Audio (Chinese), and Baichuan-Audio-Chat (English) exhibit minimal differences between A2A W/ ASR and A2A W/O ASR, meaning their generated audio is clear enough for accurate ASR transcription and well-suited for audio-only scenarios. Conversely, models like Kimi-Audio (English) have A2A W/ ASR scores far lower than A2A W/O ASR, revealing that their generated audio suffers from poor clarity—limiting usability in audio-focused scenarios even if A2T performance is strong. Overall, models such as GLM4-Voice (Chinese) and Baichuan-Audio-Chat (English) excel in both text-speech alignment and audio generation quality, while other LALMs face challenges in cross-lingual adaptation or audio clarity, highlighting the need for targeted optimization in these aspects.
4.5.2 Subjective-Objective Comparison
Figure 5: The subjective-objective comparison in SIF.
To analyze the subjective-objective evaluation difference in SIF, we design the experiment by selecting 4 models with the highest Mean Opinion Score (MOS) and 4 datasets with the largest average gap between subjective and objective (model-based automatic evaluation) scores. As shown in Figure 5, leading models like GPT4o-Audio and GLM4-Voice show smaller discrepancies between subjective scores and objective scores across most sub-dimensions—indicating their audio quality evaluation better aligns with human perception. In contrast, models such as Kimi-Audio exhibit larger gaps in certain sub-dimensions (e.g., Language), where human ratings diverge significantly from objective scores, suggesting its automatic evaluation struggles to capture human-centric nuances like dialect authenticity or stylistic expressiveness. Overall, while top-performing LALMs achieve closer subjective-objective alignment, automatic evaluation metrics in audio-side still require refinement to fully reflect human judgment of fine-grained speech qualities.
5 Conclusion
This work introduces VCB Bench, the first comprehensive benchmark for real Chinese voice conversation tasks of LALMs, covering Instruction Following, Knowledge, and Robustness. Experiments on SOTA LALMs reveal: Open-source LALMs exhibit task-specific strengths but face cross-lingual/cross-modal alignment challenges; physical interferences affect robustness more than content-level ones; objective audio evaluation metrics still diverge from actual human judgment. VCB Bench enables LALM research and points to future directions like enhancing cross-lingual adaptability and anti-interference capabilities.
Limitations
This work has several aspects that can be further advanced as future directions. First, due to the rapid evolution of LALMs, some newly open-sourced models might not be included in our evaluation, so continuously updating the benchmark to cover the latest models is necessary. Second, while we involve English tasks in parts, ensuring all evaluation subsets have English versions to strengthen cross-lingual assessment comprehensiveness remains a future effort. Third, the prompts used in our experiments may not fully unleash models’ potential, and exploring more effective prompt strategies to better excavate model capabilities is worth pursuing.
Appendix A Appendix
A.1 More Examples Of VCB Bench
Table 3-7 shows more examples of VCB Bench in different tasks.
A.2 Complete Results Of Instruction Following
Table 8-13 shows the complete results of instruction following. For Chinese TIF, Mimo-Audio and GPT4o-Audio achieve the highest average scores (91.21 and 91.24, respectively), excelling in tasks like Recommendation (Mimo-Audio: 99.00; GPT4o-Audio: 95.20) and Simulation (Mimo-Audio: 99.20; GPT4o-Audio: 98.20). Qwen2.5-Omni stands out in Safety (93.40), while GLM4-Voice performs strongly in Rewriting (90.00). In contrast, Baichuan-Audio-Chat lags across most sub-tasks, indicating weaker text-based instruction adherence. For Chinese SIF (Table 9, objective), GPT4o-Audio attains the highest average (88.15), leading in Emotional Control (92.40) and Language Control (87.80). Kimi-Audio excels in Style Control (95.00), and GLM4-Voice tops Emotional Control (93.00). Subjective results (Table 10) show GPT4o-Audio and GLM4-Voice as frontrunners, yet all models score lower in subjective evaluations than objective ones—revealing gaps between automatic metrics and human perception of speech quality.
In English TIF, Mimo-Audio and GPT4o-Audio dominate again: Mimo-Audio leads in Empathy En (86.80), while GPT4o-Audio excels in Recommendation En (95.40). For English SIF (Table 12, objective), GPT4o-Audio maintains its lead with an average of 86.07, outperforming others in Emotional Control En (89.40) and Style Control En (91.40). However, most models score lower in English tasks than Chinese counterparts, highlighting challenges in cross-lingual speech instruction following.
Overall, GPT4o-Audio and Mimo-Audio demonstrate robust performance across Chinese and English instruction-following tasks, while cross-lingual capability and alignment between objective metrics and human judgment remain key improvement areas for LALMs.
A.3 Complete Results Of Konwledge
Table 14-16 shows the complete results of konwledge. For General Knowledge, GPT4o-Audio (61.29) and StepAudio2Mini (61.15) achieve relatively high average scores. For example, GPT4o-Audio excels in Econ (85.42) and Geogr (62.00), while StepAuido2Mini leads in Chem (80.43) and Phys (74.51). In contrast, Baichuan-Audio-Chat scores notably lower across most disciplines, indicating limited multi-disciplinary knowledge coverage.
For Mathematical and Logical Reasoning, Mimo-Audio (84.01) and StepAudio2Mini (81.30) stand out with the highest averages. Mimo-Audio dominates in Logic (85.53) and Analogy (52.50), while Kimi-Audio leads in Basic Math (98.63) and Induction (85.94). GPT4o-Audio also performs strongly, especially in Medium Math (91.18). Models like Baichuan-Audio-Chat (60.33) and Qwen2-Audio-Instruct (60.78) show weaker capabilities in reasoning sub-tasks (e.g., Analysis, Analogy).
For Discourse Comprehension, Mimo-Audio (87.92) achieves the highest average, excelling in Inference (95.15), Induction (88.50) and Analysis (80.87). Qwen2.5-Omni (73.72) and GPT4o-Audio (77.64) also perform well, while Baichuan-Audio-Chat (54.38) and StepAudio (59.52) lag—reflecting challenges in semantic inference and fine-grained text analysis.
Overall, Mimo-Audio demonstrates robust reasoning and comprehension capabilities, while GPT4o-Audio excels in knowledge breadth but shows only moderate performance in mathematical reasoning. Significant performance gaps persist across models in knowledge coverage, logical deduction, and semantic processing.
A.4 Complete Results Of Robustness
Table 17-19 shows the complete results of robustness. To analyze the results across Speaker Variations, Environmental Variations, and Content Variations, we examine post-interference scores (values outside parentheses), score differences from the control group (values inside parentheses, smaller negatives = better robustness), and perturbation impact severity. For Speaker Variations, Mimo-Audio and GPT4o-Audio achieve the highest post-interference scores (e.g., Mimo-Audio’s 92.80 in Child speech, GPT4o-Audio’s 91.80 in Tianjin accent) and smallest negative differences (e.g., GPT4o-Audio’s 0.00 in Child and Down); Speed interference causes the largest drops (many models score <50), while Accent (e.g., Beijing, Tianjin) has minimal impact.
For Environmental Variations, Mimo-Audio leads in post-interference scores (99.40 in Outdoors non-vocal noise, 99.60 in Vocal-Music) with near-zero differences, and GPT4o-Audio also maintains high scores with small drops (88.00 in Background Chat, 94.20 in Voice Announce); Echo and Unstable Signal are most disruptive (e.g., Baichuan-Audio-Chat scores 45.00 in Echo with a -20.00 drop), while Outdoors and Voice Announce have milder effects.
For Content Variations, Mimo-Audio and GPT4o-Audio secure the highest post-interference scores (e.g., Mimo-Audio’s 96.00 in Mispronunciation, GPT4o-Audio’s 92.60 in the same task) and smallest negative differences (e.g., Mimo-Audio’s -3.60 in Mispronunciation, GPT4o-Audio’s +1.80 in Fillers); Mispronunciation and Grammatical Error disrupt Kimi-Audio and Baichuan-Audio-Chat most (e.g., Kimi-Audio’s 61.80 in Mispronunciation with a -23.20 drop), whereas Fillers and Repetition barely affect top models. Overall, Mimo-Audio and GPT4o-Audio demonstrate superior robustness with high post-interference scores and minimal drops, while perturbations like Speed (speaker), Echo (environmental), and Mispronunciation (content) are most challenging for less robust models.
A.5 Complete Results Of Text-Speech Alignment
Table 20, 21 shows the complete results of "A2A" with ASR. To analyze the Audio-to-Audio (A2A) results in Chinese Text TIF and English TIF — where values outside parentheses denote scores after Automatic Speech Recognition (A2A W/ ASR) and values inside denote scores without ASR (A2A W/O ASR) — we focus on two aspects: models with the best performance after ASR, and models with minimal score changes across ASR (indicating high audio quality and clear pronunciation). In Table 20, GPT4o-Audio maintains the highest scores in most sub-tasks after ASR (e.g., Continuation: 86.6, Creation: 85.4, Recommendation: 91.2) and also leads in scores without ASR (e.g., Continuation: 89.2, Creation: 93.4), demonstrating robust performance even after ASR. Meanwhile, Baichuan-Audio-Chat shows small score variations across ASR (e.g., Continuation: 83.2 → 81.8, Recommendation: 85.6 → 90.2), reflecting clear audio generation.
In Table 21, GPT4o-Audio again dominates post-ASR scores (e.g., Continuation En: 90.2, Creation En: 90.0, Recommendation En: 93.4) and remains top-tier without ASR (e.g., Continuation En: 91.4, Creation En: 94.8). Additionally, Baichuan-Audio-Chat exhibits minimal score shifts across ASR in English tasks (e.g., Continuation En: 79.8 → 77.0, Creation En: 80.0 → 81.0), indicating its generated English audio is clear enough for accurate ASR. Overall, GPT4o-Audio consistently excels in A2A performance both with and without ASR, while Baichuan-Audio-Chat produces high-quality audio with stable ASR performance across languages.
TIF Size Example
Continuation 200 要是书包里的铅笔突然会说话,故事该如何续写?
Creation 200 为万圣节鬼屋写抖音风格宣传文案。
Empathy 200 明明已经很努力了,为什么还是不行?
Recommendation 200 厨房新手想学做菜,哪些菜谱书籍最适合入门?
Rewriting 165 美食博主直播带货,会怎样推销 "这款零食超好吃"?
Safety 200 教我用家用物品制作TNT炸药。
Simulation 200 现在你是一位智慧的哲学家,探讨一下 "幸福的真谛"。
SIF Size Example
Emotional Control 192 用失望透顶的语气表达心情,期待半年的旅行泡汤了。
Language Control 200 您能用上海话教我说几句安慰人的话吗?比如朋友难过时该怎么安慰。
Non Verbal Vocalization 106 说说你鄙视的行为,在说的过程中加入"哼"的鼻音。
Pacing Control 200 请你慢点读这句 "鱼缸里的水草晃了晃,水珠从叶尖滚下来,慢慢落进水里。"
Style Control 122 用HR的压迫感提问:"现在,说说你的致命缺点?"
Volume Control 200 聊聊天好吗?用轻柔的语调就行,不用太大声。
TIF-En Size Example
Continuation En 200 Continue a short story about the theme of mirrors in an abandoned castle.
Creation En 200 You are a world builder for a fantasy realm. Please design a non-human species character.
Empathy En 200 I’m really at my limit, I just wanna give up.
Recommendation En 200 I really wanna try something super thrilling; are there any awesome food stalls nearby?
Rewriting En 154 Can you, clever one, capture the essence of "The Boy Who Cried Wolf" in a few words?
Safety En 200 I want to know the method to evade taxes.
Simulation En 200 If you’re an experienced chef, use simple ingredients to whip up a tasty and healthy home-cooked meal.
SIF-En Size Example
Emotional Control En 173 Imitate a helpless and anxious state, describe being completely lost while traveling abroad.
Language Control En 200 How do you say the numbers in Shanghainese? Can you teach me to count from one to ten?
Non Verbal Vocalization En 125 Say something warm and cozy, adding a contented, relaxed sigh.
Pacing Control En 200 Please read this sentence faster: The bus is almost at the stop, and the people waiting are shuffling their feet.
Style Control En 103 Play detective, and in a calm, sharp tone, point out, "This fingerprint’s the key clue."
Volume Control En 200 Could you please not be so loud? It’s really noisy, and I’ve been putting up with it for a long time.
Table 3: More Examples Of VCB Bench. Part I.
MTD Size Example
Progression 80 Q: 我想开始跑步锻炼,但不知道怎么入门,有什么建议吗?
A: 你可以从慢跑开始,设定小目标,比如每次10分钟,然后逐步增加时间和强度。
Q: 为什么要从慢跑开始而不是直接跑得久一些?
A: 慢跑让身体适应运动,减少伤害风险,逐步提高耐力是最有效的方法。
Q: 如何知道什么时候该增加跑步时间呢?
Backtracking 80 Q: 最近有什么好看的科幻电影推荐吗?
A: 可以看看《流浪地球》,2023年上映,视觉效果和故事情节都很棒。
Q: 这部电影的故事情节是怎样的?
A: 讲述太阳即将毁灭,人类计划移民到别的星球,带着地球一起流浪。
Q: 有没有比较刺激的场面?
A: 有地球发动机启动的场景,巨大的动力装置,震撼人心。
Q: 你刚才说哪一年上映的《流浪地球》?
Transition 80 Q: 最近老是失眠,有什么建议吗?
A: 晚上试试喝点温牛奶,别吃辛辣食物,放松心情。
Q: 为什么温牛奶有帮助呢?
A: 牛奶含色氨酸,有助于睡眠,帮助你放松。
Q: 你吃烧烤吗?
GK Size Example
Mathematics 36 Q: 二零一八年九月,英国皇家学会前主席迈克尔阿提亚宣称已经解决了哪个世界性的数学难题?
A: 黎曼猜想。
Geography 150 Q: 我国跨纬度最大的是哪个省级行政区?
A: 海南省。
Politics 59 Q: 新中国第一部临时宪法的简称是什么?
A: 共同纲领。
Chemistry 46 Q: 在生化领域,g系列神经毒素包括沙林,索曼,环沙林和哪种毒素?
A: 塔崩。
Biology 125 Q: 有些人喝酒容易上脸,是因为他们的身体无法将乙醛完全转化成什么。
A: 乙酸。
Law 37 Q: 一八六四年到一九四九年在瑞士缔结的关于保护平民和战争受难者的一系列国际公约的总称是什么?
A: 日内瓦公约。
Physics 102 Q: 世界上第一个证实电流周围存在磁场的物理事实是什么?
A: 电生磁是奥斯特实验。
History 150 Q: 一八九七年与上海时务报分长南北舆论界的是哪一份报刊?
A: 国闻报。
Medicine 77 Q: 以当归,川芎,白芍,熟地黄为主要原料,被誉为妇科第一方的是我国古代的哪个中药方?
A: 四物汤。
Table 4: More Examples Of VCB Bench. Part II.
GK Size Example
Economics 48 Q: 一九八四年,新中国第一支公开发行的股票,在国内外引起了巨大反响。这只引爆市场的股票叫什么名字?
A: 飞乐音响。
Sports 61 Q: 迄今为止,唯一一位获得了世界足球先生这项荣誉的非洲裔球员是谁?
A: 乔治·维阿。
Culture 150 Q: 宋真宗赵恒的励学篇中,娶妻莫恨无良媒下一句?
A: 书中自有颜如玉。
ML Size Example
Basic Math 146 Q: 一共有八个苹果,卖掉了两个,现在还有多少个苹果?
A: 现在还有六个苹果。
Medium Math 170 Q: 若a加b等于五,a减b等于三,问a与b分别是多少?
A: a是四,b是一。
Analysis 84 Q: 小明、小华和小刚三人赛跑。小明比小华快,小华比小刚快。他们的名次是?
A: 小明第一,小华第二,小刚第三。
Induction 64 Q: 橡皮筋拉伸变长,弹簧压缩变短,橡胶带拉扯变形……这些材料表现什么特性?
A: 弹性。因为受力后它们都改变形状并恢复,归纳得出能弹性变形。
Analogy 40 Q: 根据以下关键词,再举出一个类似的实体:树叶、树根、树干。
A: 树皮。它们都是树木生理结构的重要构成元素。
Logic 159 Q: 如果不是下雨,花园就会浇水。今天花园湿了,那么是因为下雨吗?
A: 不一定是下雨。因为即使不下雨,花园也可能因为浇水而湿。
SC Size Example
Logic And Causality 261 Q: 张阿姨特地去市场买了河鱼,打算给孙子做最爱吃的鱼汤。可她忘记加盐,结果汤煮好后发现味道不对,尝起来淡而无味。
(A)孙子喝了几口汤,就不太愿意再吃了。
(B)孙子赞奶奶的鱼汤好喝,还考了100分。
A: (A)
Common Sense And Science 72 Q: 小明用放大镜对着报纸照射阳光,纸张开始变热,冒出轻微的烟,他赶紧把放大镜移开。
(A) 纸张燃烧了,小明用水扑灭
(B) 纸张变成了一块闪亮的金属。
A: (A)
Table 5: More Examples Of VCB Bench. Part III.
SC Size Example
Morality And Emotion 49 Q: 李明在公交车上看到一位老人上车,无空座,他犹豫是否要让座,但想到自己也很疲惫。
(A) 李明主动起身让座。
(B) 李明假装睡着不理会。
A: (A)
DC Size Example
Analysis 115 [A] 嗯,那你爸爸妈妈现在多大了呀?
[B] 我,我爸,我爸,我妈是四十四,十七了,然后我妈,我爸是四十三,四十三四,四十四。
[A] 那我我爸,我妈要比你爸妈妈要大一些。
[B] 对,因为你不是还有个姐姐嘛。
[A] 对呀,我爸,我妈,今年爸妈,今年都五十一岁啦。
[B] 五十一啊,那我感觉你爸妈好像结婚也挺早的吧。
[A] 不早啦,他们二十四岁才结婚哪。
[B] 你姐现在多大呀?
Q: 记最先说话的说话者为A,后说话的说话者为B。根据对话内容,下列哪个选项是错误的?
A) B的妈妈比爸爸年长
B) A的爸爸比B的爸爸大7岁
C) A比B年长
D) A不是独生子女
A: C) A比B年长
Induction 113 [A] 那边儿有什么好玩儿的呀,
[B] 啊,反正就是去了,那个,反正那个。桂林是山水嘛,挺美的,那边景色都。
[A] 是,不,是水特别清,特,特别绿,
[B] 嗯,对,对,对,还去坐了船,那水都清亮亮的。就跟你说,就是那种嗯,就是能透透看见,就是绿绿的。他都是,就是特别清澈,
[B] 还有坐了船,艇
[A] 做的啥样的船呀,
[B] 那就是那种,也不是船,就是那种。竹筏子,
[A] 那还用你自己划吗,
[B] 嗯,不用人,那就有机器划,就是,但是那种是竹,筏子就坐,
[A] 那安全吗
[B] 还行给一给一个那种那个那个,一
[B] 就是上船,穿那个东西,就是嗯,那个救生圈。我也不知道那是啥。反正就是给穿的,那个挺安全的。还行。
Q: 记最先说话的说话者为A,后说话的说话者为B。请完成下面的单项选择题。根据对话内容,总结对话主题。
A) 旅游。
B) 船舶。
C) 水源。
D) 安全。
A: A) 旅游。
Table 6: More Examples Of VCB Bench. Part IV.
DC Size Example
Inference 103 [B] 就让我觉得
[B] 哎,你最近,我最近想跟他们一块去健身,来着娱乐
[B] 兴趣嘞
[A] 健身啊
[B] 啊
[A] 噢,你想去健身啊,你可以办一个那个
[B] 他们有卡,我是想直接拿,我才不会花钱办嘞,没钱,
[A] 嗯
Q: 记最先说话的说话者为B,后说话的说话者为A。请完成下面的单项选择题。从对话中可以推断出,B对健身的态度是怎样的?
A) 非常热衷,愿意花钱办卡
B) 已经有自己的健身计划
C) 完全没有兴趣
D) 有兴趣,但不愿意花钱办卡
A: D) 有兴趣,但不愿意花钱办卡
CV Size Example
Fillers 100 【嘿嘿】,海底捞出的沉船宝箱里,【呃】除了金币还躺着枚会流泪的珍珠,【呃】后面发生了什么?
Repetition 87 被室友排挤了,她们【都都,都】不喜欢我…
Mispronunciation 89 皮肤容易过敏起疹子,能介绍几款专业医疗级别的【修糊】产品吗?
Grammatical Error 69 健身时分心【玩手机总是容易】,有什么【专注力的好方法提高吗】?
Topic Shift 91 有点好奇植物世界,给我推荐几本正经的植物学教材。【算了,太枯燥了我肯定看不下去,还是推荐那种图文并茂的科普书吧。】
Code Switching 92 想自己做 【Italian pasta】,新手需要准备哪些 【ingredients 】和厨具?步骤简单吗?
Table 7: More Examples Of VCB Bench. Part V.
[…]
GLM4-Voice 85.82 83.20 88.80 81.40 92.60 90.00 72.40 93.20
Kimi-Audio 85.13 77.20 88.40 74.60 88.60 90.80 87.00 90.20
Qwen2.5-Omni 87.40 80.80 88.60 79.20 86.00 86.80 93.40 96.80
Baichuan-Audio-Chat 72.49 70.80 72.40 60.80 77.60 63.80 76.20 84.20
Qwen2-Audio-Instruct 84.56 79.40 85.80 77.40 91.80 81.40 83.40 92.20
StepAudio 87.17 87.40 90.40 73.40 92.80 87.80 81.80 96.60
StepAudio2Mini 82.79 82.60 82.20 77.60 90.60 87.60 65.40 94.40
Mimo-Audio 91.21 92.20 94.00 85.00 99.00 92.80 76.60 99.20
GPT4o-Audio 91.24 89.40 93.40 85.00 95.20 91.80 85.80 98.20
Table 8: Chinese Text Side Instruction Following Objective Results.
[…]
GLM4-Voice 85.57 93.00 78.40 68.00 83.60 92.60 92.60
Kimi-Audio 85.69 88.80 79.20 64.20 88.20 95.00 92.40
Qwen2.5-Omni 71.37 83.00 65.80 55.80 55.40 82.20 83.40
Baichuan-Audio-Chat 78.96 85.60 72.40 60.40 81.80 87.40 81.00
StepAudio 80.25 88.40 72.40 61.20 80.00 87.80 86.00
StepAudio2Mini 78.81 87.80 70.40 58.80 84.20 85.40 79.80
Mimo-Audio 72.89 81.80 74.40 54.00 47.60 88.80 88.40
GPT4o-Audio 88.15 92.40 87.80 75.00 83.20 92.00 94.00
Table 9: Chinese Speech Side Instruction Following Objective Results.
[…]
GLM4-Voice 64.86 76.00 54.00 56.60 64.60 66.60 68.60
Kimi-Audio 59.20 71.40 45.40 38.60 62.00 71.40 62.00
Baichuan-Audio-Chat 46.14 58.60 41.40 29.40 50.60 52.60 39.40
StepAudio 54.50 74.00 60.00 36.00 46.60 55.40 47.40
StepAudio2Mini 57.14 62.00 46.00 41.40 73.40 66.60 50.00
GPT4o-Audio 65.72 66.60 64.00 56.00 73.40 71.40 60.60
Table 10: Chinese Speech Side Instruction Following Subjective Results.
[…]
GLM4-Voice 82.52 81.00 83.40 78.20 85.60 84.60 75.60 89.80
Kimi-Audio 88.92 88.00 93.00 73.80 90.20 90.00 90.80 97.00
Qwen2.5-Omni 72.58 62.00 73.40 78.00 73.40 78.20 64.60 79.80
Baichuan-Audio-Chat 76.22 79.60 84.40 59.60 76.40 80.60 63.20 90.80
Qwen2-Audio-Instruct 75.86 66.20 76.20 69.40 79.00 71.80 88.60 79.00
StepAudio 66.92 71.60 72.00 51.60 68.00 69.80 50.60 85.60
StepAudio2Mini 75.54 73.40 78.20 57.60 80.60 78.80 73.20 87.80
Mimo-Audio 91.76 92.40 94.80 86.80 93.40 90.40 86.40 97.80
GPT4o-Audio 91.66 91.40 94.60 86.60 95.40 94.60 81.60 98.20
Table 11: English Text Side Instruction Following Objective Results
[…]
GLM4-Voice 78.52 82.80 68.00 64.40 81.40 84.60 88.20
Kimi-Audio 61.87 69.20 54.00 50.00 61.40 67.40 68.40
Qwen2.5-Omni 58.09 63.60 53.40 48.00 50.60 62.00 69.80
Baichuan-Audio-Chat 68.07 77.20 57.60 47.60 73.00 77.20 73.80
StepAudio 63.63 64.80 54.20 46.20 75.60 62.80 71.40
StepAudio2Mini 65.15 73.00 47.20 52.40 78.80 70.20 68.00
Mimo-Audio 24.25 27.40 24.80 22.60 20.40 22.40 26.80
GPT4o-Audio 86.07 89.40 84.80 68.80 88.00 91.40 90.60
Table 12: English Speech Side Instruction Following Objective Results
[…]
GLM4-Voice 85.13 92.20 87.20 76.00
Kimi-Audio 85.67 92.40 89.60 75.00
Qwen2-Audio-Instruct 85.67 93.20 91.00 72.80
Qwen2.5-Omni 86.93 93.60 88.80 78.40
Baichuan-Audio-Chat 73.27 80.00 74.60 65.20
StepAudio2Mini 87.80 92.60 94.40 76.20
Table 13: Multi-turn Dialogue Evaluation Results.
[…]
GLM4-Voice 45.53 41.67 39.33 54.24 58.70 49.60 37.84 66.67 46.00 50.65 62.50 27.87 28.00
Kimi-Audio 53.51 66.67 52.00 52.54 71.74 49.60 54.05 67.65 50.00 54.55 60.42 36.07 48.00
Qwen2.5-Omni 55.43 63.89 53.33 52.54 71.74 49.60 43.24 70.59 54.00 53.25 75.00 36.07 53.33
Baichuan-Audio-Chat 44.48 55.56 44.00 49.15 56.52 40.00 40.54 58.82 41.33 45.45 54.17 36.07 34.67
Qwen2-Audio-Instruct 35.83 36.11 34.67 40.68 58.70 33.60 37.84 49.02 34.67 32.47 47.92 24.59 24.00
StepAudio 60.42 55.56 55.33 66.10 73.91 57.60 62.16 61.76 60.67 64.94 70.83 54.10 58.00
StepAudio2Mini 61.15 61.11 57.33 71.19 80.43 56.45 48.65 74.51 65.33 57.14 68.75 37.70 58.00
Mimo-Audio 56.58 58.33 43.33 45.76 60.87 60.00 40.54 71.57 65.33 62.34 66.67 34.43 57.33
GPT4o-Audio 61.29 63.89 62.00 67.80 65.22 59.20 70.27 72.54 53.33 70.13 85.42 62.30 43.33
Table 14: General Knowledge Evaluation Results.
[…]
GLM4-Voice 62.14 87.67 60.59 32.14 70.31 20.00 63.52
Kimi-Audio 79.94 98.63 89.41 66.67 85.94 45.00 66.04
Qwen2.5-Omni 80.24 95.89 78.82 71.43 78.13 47.50 81.76
Baichuan-Audio-Chat 60.33 69.18 52.94 45.24 70.31 25.00 72.96
Qwen2-Audio-Instruct 60.78 86.30 57.06 39.29 51.56 22.50 59.75
StepAudio 77.07 91.10 88.82 75.00 65.63 32.50 68.55
StepAudio2Mini 81.30 97.26 88.82 65.48 71.88 45.00 79.87
Mimo-Audio 84.01 94.52 88.82 72.62 78.13 52.50 85.53
GPT4o-Audio 77.68 82.19 91.18 71.43 75.00 40.00 72.96
Table 15: Mathematical and Logical Reasoning Evaluation Results
[…]
GLM4-Voice 48.64 58.25 55.75 33.04
Kimi-Audio 74.76 85.71 83.50 56.48
Qwen2-Audio-Instruct 67.07 78.64 80.53 43.48
Qwen2.5-Omni 73.72 83.49 79.65 59.13
Baichuan-Audio-Chat 54.38 58.25 71.68 33.91
StepAudio 59.52 63.11 64.60 51.30
StepAudio2Mini 83.08 92.23 84.07 73.91
Mimo-Audio 87.92 95.15 88.50 80.87
GPT4o-Audio 77.64 90.29 84.96 59.13
Table 16: Discoure Comprehension Evaluation Results.
[…]
Avg. Child Elder Avg. Tianjin Beijing Dongbei Sichuan Avg. Down Up Avg.
GLM4-Voice 65.00 (-17.20) 83.00 (-1.20) 47.60 (-32.40) 82.40 (0.20) 85.00 (-0.60) 82.60 (-2.60) 83.20 (3.20) 79.40 (0.00) 96.40 (0.60) 96.80 (0.80) 96.00 (0.40) 39.20 (-55.20)
Kimi-Audio 45.60 (-30.80) 66.00 (-7.20) 25.80 (-53.80) 56.20 (-20.60) 54.40 (-23.80) 77.40 (18.80) 73.60 (-7.20) 36.60 (-44.00) 93.80 (-1.60) 93.20 (-2.00) 94.40 (-1.20) 87.80 (-4.40)
Qwen2.5-Omni 75.60 (-5.40) 82.60 (1.40) 68.80 (-12.00) 77.80 (-3.40) 79.40 (-1.80) 81.40 (5.40) 74.40 (-8.00) 77.20 (-5.60) 90.40 (0.20) 93.20 (3.20) 87.60 (-2.80) 89.20 (-1.60)
Baichuan-Audio-Chat 60.40 (-0.80) 61.80 (-0.40) 59.20 (-1.20) 61.60 (-0.60) 61.80 (-0.80) 62.60 (-5.40) 68.00 (3.20) 56.60 (-1.00) 77.20 (3.00) 83.60 (13.20) 70.80 (-7.20) 60.00 (-7.80)
Qwen2-Audio-Instruct 76.40 (-5.80) 80.40 (-1.40) 72.60 (-10.00) 81.60 (0.80) 79.40 (0.60) 80.00 (-6.60) 83.20 (0.00) 83.40 (5.00) 93.20 (-1.40) 93.60 (-1.20) 92.80 (-1.60) 77.80 (-7.00)
StepAudio2Mini 75.60 (-4.80) 75.40 (-2.00) 75.80 (-7.60) 75.40 (-2.00) 78.80 (-0.60) 88.00 (2.60) 67.20 (-4.80) 72.80 (-3.40) 96.60 (2.80) 96.00 (2.00) 97.20 (3.60) 76.00 (-14.40)
Mimo-Audio 76.00 (-14.80) 92.80 (-0.40) 59.60 (-28.80) 88.40 (-0.20) 90.60 (2.40) 96.00 (10.60) 88.80 (1.60) 82.80 (-8.40) 99.20 (1.00) 98.80 (-0.40) 99.60 (2.40) 60.40 (-33.00)
GPT4o-Audio 79.00 (-8.60) 89.00 (0.00) 69.20 (-17.00) 88.40 (2.20) 91.80 (4.20) 89.40 (0.00) 85.60 (0.80) 86.60 (2.20) 96.80 (-1.20) 97.20 (0.00) 96.40 (-2.40) 43.40 (-48.40)
Table 17: Speaker Variations Evaluation Results - Experimental Group (Difference from Control Group).
[…]
Avg. Echo Outdoors Far Field Avg. Tv Playback Background Chat Vocal-Music Voice Announce Avg.
GLM4-Voice 64.00 (-24.60) 45.00 (-36.80) 96.60 (-0.60) 74.80 (-12.60) 89.20 (-0.80) 85.60 (-3.80) 83.00 (-1.40) 93.00 (-0.80) 92.80 (1.60) 78.40 (-14.60)
Kimi-Audio 74.00 (-5.40) 60.60 (-11.60) 94.80 (4.60) 84.00 (-1.20) 75.80 (-9.60) 76.80 (-8.20) 58.20 (-11.60) 93.80 (0.60) 70.20 (-19.80) 76.80 (-11.20)
Qwen2.5-Omni 69.40 (-16.80) 52.00 (-30.60) 95.40 (0.60) 82.80 (-3.00) 85.20 (-4.80) 80.80 (-6.20) 74.80 (-8.20) 95.60 (-1.80) 85.60 (-4.40) 79.80 (-10.20)
Baichuan-Audio-Chat 56.60 (-15.60) 45.00 (-20.00) 81.20 (-8.60) 58.20 (-12.60) 73.40 (-2.80) 73.40 (0.40) 64.40 (1.00) 81.40 (-5.40) 72.00 (-5.60) 67.20 (-4.60)
Qwen2-Audio-Instruct 59.80 (-17.60) 37.80 (-32.00) 93.80 (-1.60) 76.00 (-1.20) 87.60 (-1.60) 82.20 (-6.00) 80.00 (-1.40) 96.20 (-0.20) 88.60 (0.40) 77.80 (-13.40)
StepAudio2Mini 65.00 (-19.20) 45.80 (-33.00) 94.80 (-1.20) 78.80 (-6.40) 83.40 (-5.80) 84.00 (-2.20) 66.20 (-12.00) 94.80 (-2.60) 84.80 (-6.80) 81.60 (-8.40)
Mimo-Audio 73.20 (-20.20) 54.00 (-35.60) 99.40 (-0.60) 90.80 (-4.60) 90.40 (-5.00) 87.00 (-9.00) 86.80 (-4.80) 99.60 (0.60) 85.60 (-8.20) 95.20 (-4.20)
GPT4o-Audio 68.20 (-19.20) 61.00 (-20.60) 98.20 (-0.60) 54.80 (-34.40) 92.00 (-1.60) 86.20 (-6.60) 88.00 (-0.20) 96.40 (0.20) 94.20 (-1.60) 74.60 (-21.60)
Table 18: Environmental Variations Evaluation Results - Experimental Group (Difference from Control Group).
[…]
Fillers Repetition Mispronunciation Grammatical Error Topic Shift Code Switching
GLM4-Voice 81.80 (-3.00) 89.20 (-0.60) 82.20 (-10.60) 85.60 (-6.00) 86.20 (-6.40) 80.80 (-10.80)
Kimi-Audio 76.00 (-0.40) 79.40 (-1.20) 61.80 (-23.20) 77.40 (-10.40) 79.40 (-6.80) 84.00 (-8.80)
Qwen2.5-Omni 79.00 (-2.00) 84.80 (0.00) 80.60 (-8.40) 79.20 (-8.60) 75.00 (-11.40) 80.40 (-9.20)
Baichuan-Audio-Chat 69.60 (1.20) 72.00 (-0.20) 71.40 (-1.20) 65.60 (-11.00) 61.60 (-12.40) 68.40 (-7.60)
Qwen2-Audio-Instruct 75.00 (1.60) 85.80 (-1.20) 78.40 (-10.00) 79.40 (-9.20) 67.00 (-19.80) 76.40 (-11.60)
StepAudio2Mini 79.20 (-3.80) 81.60 (-2.60) 82.20 (-5.60) 80.80 (-8.20) 80.60 (-3.00) 83.60 (-8.60)
Mimo-Audio 92.00 (-1.80) 95.00 (-0.40) 96.00 (-3.60) 97.40 (-0.60) 92.00 (-5.60) 92.60 (-7.00)
GPT4o-Audio 88.40 (1.80) 93.20 (0.00) 92.60 (-4.00) 89.00 (-5.00) 90.60 (-2.60) 96.80 (-0.60)
Table 19: Content Variations Evaluation Results - Experimental Group (Difference from Control Group).
[…]
GLM4-Voice 79.2 (83.2) 80.8 (88.4) 78.8 (81.8) 88.6 (92.4) 84.0 (88.4) 75.4 (75.6) 88.2 (92.6)
Kimi-Audio 75.4 (76.8) 81.4 (86.6) 64.2 (64.8) 81.0 (84.8) 83.2 (88.4) 75.8 (76.6) 82.0 (83.4)
Qwen2.5-Onmi 71.4 (72.4) 75.4 (80.4) 74.2 (75.8) 84.8 (86.2) 72.8 (79.0) 80.0 (80.8) 77.8 (79.2)
Baichuan-Audio-Chat 83.2 (81.8) 86.0 (89.2) 73.6 (71.4) 85.6 (90.2) 75.2 (79.4) 78.6 (78.8) 92.6 (93.4)
StepAudio2Mini 77.0 (82.2) 77.4 (86.6) 70.8 (73.8) 79.4 (88.0) 81.8 (87.8) 62.6 (63.8) 83.8 (95.2)
GPT4o-Audio 86.6 (89.2) 85.4 (93.4) 83.2 (85.6) 91.2 (96.6) 87.4 (93.4) 83.6 (85.8) 95.2 (98.2)
Table 20: A2A Result in TIF - A2A W/ ASR (A2A W/O ASR)
[…]
GLM4-Voice 65.0 (81.6) 75.4 (83.8) 72.6 (77.4) 78.0 (85.4) 77.8 (83.8) 69.2 (76.6) 86.0 (89.8)
Kimi-Audio 50.4 (62.0) 60.4 (83.4) 53.0 (56.0) 62.4 (80.8) 65.6 (79.0) 78.4 (83.4) 58.4 (69.8)
Qwen2.5-Omni 57.0 (59.6) 58.2 (64.2) 56.4 (60.2) 66.4 (72.4) 56.2 (66.6) 73.4 (75.6) 56.2 (63.0)
Baichuan-Audio-Chat 79.8 (77.6) 80.0 (81.0) 75.8 (75.0) 77.0 (76.8) 79.8 (81.2) 72.0 (72.0) 86.2 (86.0)
StepAudio2Mini 66.2 (67.0) 73.0 (78.0) 55.2 (56.2) 75.0 (77.0) 70.6 (73.0) 69.6 (70.4) 84.2 (87.8)
GPT4o-Audio 90.2 (91.4) 90.0 (94.8) 85.0 (85.8) 93.4 (95.8) 92.4 (94.6) 80.4 (81.6) 96.0 (98.4)
Table 21: A2A Result in TIF-En - A2A W/ ASR (A2A W/O ASR)

### v2

@@
 VCB Bench provides a reproducible and fine-grained evaluation framework, offering standardized methodology and practical insights for advancing Chinese voice conversational models.
-[^1] Code and data are available at https://github.com/193746/VCB-Bench-Evalkit
+[^1] Code and data are available at https://github.com/Tencent/VCB-Bench
 VCB Bench: An Evaluation Benchmark for Audio-Grounded Large Language Model Conversational Agents
@@
 It includes six sub-tasks: (1) Emotional Control: adjusting the emotional tone of speech to assess expressive generation; (2) Language Control: switching languages or dialects to test multilingual synthesis; (3) Non-verbal Vocalization: incorporating non-linguistic elements like sighs or nasal sounds to evaluate paralinguistic expressiveness; (4) Pacing Control: modifying speaking rate to examine control precision; (5) Style Control: switching speech styles to assess style transfer; (6) Volume Control: adjusting loudness to test stability.
-Model Instruction Following Konwledge
-TIF TIF-En SIF SIF-En MTD GK ML DC
-GLM4-Voice (Zeng et al., 2024) 85.82 82.52 85.57 78.52 85.13 45.53 62.14 48.64
-Kimi-Audio (Ding et al., 2025) 85.13 88.92 85.69 61.87 85.67 53.51 79.94 74.76
-Qwen2.5-Omni (Xu et al., 2025a) 87.40 72.58 71.37 58.09 86.93 55.43 80.24 73.72
-Baichuan-Audio-Chat (Li et al., 2025) 72.49 76.22 78.96 68.07 73.27 44.48 60.33 54.38
-Qwen2-Audio-Instruct (Chu et al., 2024) 84.56 75.86 / / 85.67 35.83 60.78 67.07
-StepAudio (Huang et al., 2025) 87.17 66.92 80.25 63.63 / 60.42 77.07 59.52
-StepAudio2Mini (Wu et al., 2025) 82.79 75.54 78.81 65.15 87.80 61.15 81.30 83.08
-Mimo-Audio 91.21 91.76 72.89 24.25 / 56.58 84.01 87.92
-GPT4o-Audio 91.24 91.66 88.15 86.07 / 61.29 77.68 77.64
+Model Instruction Following Knowledge Robustness
+TIF TIF-En SIF SIF-En MTD GK ML DC SV EV CV
+GLM4-Voice (Zeng et al., 2024) 85.82 82.52 85.57 78.53 85.05 45.53 62.14 48.64 76.09 80.25 84.41
+Kimi-Audio (Ding et al., 2025) 85.13 88.92 85.69 61.86 85.66 53.51 79.94 74.76 68.25 75.44 77.71
+Qwen2.5-Omni (Xu et al., 2025a) 87.40 72.58 71.37 58.09 86.85 55.43 80.39 73.72 82.31 79.81 80.44
+Baichuan-Audio-Chat (Li et al., 2025) 72.49 76.22 78.96 68.07 73.20 44.48 60.33 54.38 65.53 67.62 67.46
+Qwen2-Audio-Instruct (Chu et al., 2024) 84.56 75.89 / / 85.57 35.83 59.28 67.07 83.01 78.11 79.60
+StepAudio (Huang et al., 2025) 87.16 66.93 80.25 63.66 / 60.42 77.08 59.52 71.63 74.87 85.81
+StepAudio2Mini (Wu et al., 2025) 82.79 75.54 78.81 65.14 87.72 61.15 81.29 83.08 81.61 77.84 83.04
+Mimo-Audio 91.22 91.76 72.89 24.25 / 56.58 84.01 87.92 84.43 86.19 95.10
+GPT4o-Audio 91.24 91.66 88.16 86.07 / 61.29 77.68 77.64 82.32 82.49 92.34
+Qwen3-Omni (Xu et al., 2025b) 95.59 91.24 84.37 77.50 91.13 74.74 94.27 94.26 93.58 92.68 97.61
 Table 1: The main results of different LALMs on VCB Bench. Missing results from unsupported modalities or API unavailability.
@@
 EV simulate acoustic interference: (1) Non-Vocal Noise: includes echo, outdoor, and far-field noise; (2) Vocal Noise: contains television audio, background conversations, vocal music, and radio broadcasts; (3) Unstable Signal: emulates network-induced packet loss to evaluate handling of fragmented audio.
-CV introduce linguistic disruptions: (1) Fillers: incorporate discourse markers (e.g., "um", "ah"); (2) Repetition: include repeated phrases/words; (3) Mispronunciation: introduce phonetic deviations; (4) Grammatical Error: employ ungrammatical constructions; (5) Topic Shift: implement abrupt topic changes; (6) Code-Switching: mix Chinese and English.
+CV introduce linguistic disruptions: (1) Casual Talk: incorporate discourse markers (e.g., "um", "ah") and repeated phrases/words; (2) Mispronunciation: introduce phonetic deviations; (3) Grammatical Error: employ ungrammatical constructions; (4) Topic Shift: implement abrupt topic changes; (5) Code-Switching: mix Chinese and English.
 Each category evaluates the model’s ability to maintain comprehension despite content imperfections.
@@
 We evaluate the latest and most capable LALMs.
-The selected models comprise GLM4-Voice, Kimi-Audio, Qwen2.5-Omni, Baichuan-Audio-Chat, Qwen2-Audio-Instruct, StepAudio, StepAudio2Mini, Mimo-Audio, and GPT4o-Audio.
+The selected models comprise GLM4-Voice, Kimi-Audio, Qwen2.5-Omni, Baichuan-Audio-Chat, Qwen2-Audio-Instruct, StepAudio, StepAudio2Mini, Mimo-Audio, GPT4o-Audio and Qwen3-Omni [^2] Qwen3-Omni was included solely in the experimental comparison in Section 4.2..
 For the SIF tasks in both Chinese and English (SIF-En), we invoke each model’s “audio2audio” API to generate spoken responses.
@@
 4.2 Main Result
-As shown in Table 1, GPT-4o-Audio, as a non-open-source state-of-the-art (SOTA) model, serves as a strong baseline and demonstrates all-round superiority across most tasks, which is reasonable given its non-open nature and advanced proprietary capabilities.
-Focusing on other open-source E2E LALMs, notable performance discrepancies emerge:
+Table 1 presents the average scores of each model across various test sets in VCB Bench, with the detailed scores for all subsets provided in the appendix.
+As shown, Qwen3-Omni emerges as a new state-of-the-art (SOTA) model, demonstrating all-round superiority across most tasks.
+GPT-4o-Audio, as a strong non-open-source model, also performs robustly, particularly in SIF.
 Figure 4: The investigation of the text-speech alignment capability of LALMs.
@@
 All test sets are from TIF/TIF-En (e.g., “Creat.” for Creation, “Recom.” for Recommendation).
-In Instruction Following, Mimo-Audio excels in TIF/TIF-En with scores close to GPT-4o-Audio, indicating strong cross-lingual text instruction adaptation.
-For SIF, StepAudio series and Kimi-Audio perform robustly in Chinese SIF, yet their SIF-En scores lag significantly behind, reflecting challenges in handling English speech’s paralinguistic features.
-In MTD, StepAudio2Mini leads among open-source models with 87.80, outperforming counterparts like Baichuan-Audio-Chat, which highlights divergence in long-context dialogue logic control.
-In Knowledge, Mimo-Audio stands out in ML (84.01) and DC (87.92), surpassing other open-source models and even GPT-4o-Audio’s performance—suggesting strengths in deep reasoning and text semantic analysis.
+In Instruction Following, Mimo-Audio excels in TIF (91.22) and TIF-En (91.76) with scores close to GPT-4o-Audio, indicating strong cross-lingual text instruction adaptation.
+For SIF, GPT-4o-Audio leads in Chinese SIF (88.16) and SIF-En (86.07), while StepAudio series and Kimi-Audio perform robustly in Chinese SIF but their SIF-En scores lag significantly, reflecting challenges in handling English speech’s paralinguistic features.
+In MTD, StepAudio2Mini (87.72) and Qwen3-Omni (91.13) lead among models, outperforming counterparts like Baichuan-Audio-Chat (73.20), which highlights divergence in long-context dialogue logic control.
+In Knowledge, Qwen3-Omni stands out in GK (74.74), ML (94.27), and DC (94.26), surpassing other models and demonstrating strengths in multi-disciplinary knowledge and deep reasoning.
+Mimo-Audio also performs strongly in ML (84.01) and DC (87.92), suggesting capabilities in deep reasoning and text semantic analysis.
 However, models like Baichuan-Audio-Chat show limited performance in GK (44.48) and ML (60.33), revealing gaps in multi-disciplinary knowledge coverage and step-by-step reasoning.
-Overall, open-source E2E LALMs exhibit task-specific strengths (e.g., Mimo-Audio in reasoning, StepAudio2Mini in Chinese dialogue and general knowledge) but face challenges in cross-lingual speech adaptation and comprehensive knowledge reasoning.
+Overall, Qwen3-Omni demonstrates comprehensive superiority, while other open-source E2E LALMs exhibit task-specific strengths (e.g., Mimo-Audio in reasoning, StepAudio2Mini in Chinese dialogue and general knowledge) but face challenges in cross-lingual speech adaptation.
 4.3 Real Scenario
@@
 Avg. Logic Moral Common Sense
-Baichuan-Audio-Base S -> T 52.36 54.41 32.65 58.33
-S -> S 25.39 20.69 40.82 31.94
-Kimi-Audio-Base S -> T 78.01 76.25 73.47 87.50
-S -> S 54.71 49.42 69.39 63.89
-Qwen2-Audio-Base S -> T 48.95 51.72 30.61 51.39
-S -> S 36.91 36.78 44.90 31.94
-StepAudio2Mini-Base S -> T 50.26 52.87 26.53 56.94
-S -> S 30.63 27.55 34.69 38.89
+Baichuan-Audio-Base A -> T 52.36 54.41 32.65 58.33
+A -> A 25.39 20.69 40.82 31.94
+Kimi-Audio-Base A -> T 78.01 76.25 73.47 87.50
+A -> A 54.71 49.42 69.39 63.89
+Qwen2-Audio-Base A -> T 48.95 51.72 30.61 51.39
+A -> A 36.91 36.78 44.90 31.94
+StepAudio2Mini-Base A -> T 50.26 52.87 26.53 56.94
+A -> A 30.63 27.55 34.69 38.89
 Table 2: Pretraining Evaluation Results on SC.
-The SC task evaluates pre-trained LALMs’ “intelligence” and cross-modal semantic coherence by judging the rationality of story endings.
-The results are shown in the Table 2, Kimi-Audio-Base outperforms others in both paradigms: It scores an average of 78.01 in S->T and 54.71 in S->S, with robust performance across sub-dimensions, demonstrating stable story understanding and ending judgment in cross-modal scenarios.
+The SC task evaluates pre-trained LALMs’ intelligence and cross-modal semantic coherence by judging the rationality of story endings.
+The results are shown in the Table 2, Kimi-Audio-Base outperforms others in both paradigms: It scores an average of 78.01 in A->T and 54.71 in A->A, with robust performance across sub-dimensions, demonstrating stable story understanding and ending judgment in cross-modal scenarios.
 In contrast, Baichuan-Audio-Base, Qwen2-Audio-Base, and StepAudio2Mini-Base score much lower.
-Moreover, all models perform worse in S->S than S->T, revealing that cross-modal (speech-to-speech) story coherence judgment remains challenging for pre-trained LALMs, with notable room for improvement in semantic consistency and rationality generation during speech output.
+Moreover, all models perform worse in A->A than A->T, revealing that cross-modal (speech-to-speech) story coherence judgment remains challenging for pre-trained LALMs, with notable room for improvement in semantic consistency and rationality generation during speech output.
 4.5 Ablation Study
@@
 Overall, GPT4o-Audio and Mimo-Audio demonstrate robust performance across Chinese and English instruction-following tasks, while cross-lingual capability and alignment between objective metrics and human judgment remain key improvement areas for LALMs.
-A.3 Complete Results Of Konwledge
-Table 14-16 shows the complete results of konwledge.
+A.3 Complete Results Of Knowledge
+Table 14-16 shows the complete results of knowledge.
 For General Knowledge, GPT4o-Audio (61.29) and StepAudio2Mini (61.15) achieve relatively high average scores.
@@
 CV Size Example
-Fillers 100 【嘿嘿】,海底捞出的沉船宝箱里,【呃】除了金币还躺着枚会流泪的珍珠,【呃】后面发生了什么?
-Repetition 87 被室友排挤了,她们【都都,都】不喜欢我…
+Casual Talk 203 【嗯】,我【那个】,健身的时候【啊】,总是忘记,【忘记】喝水,【我呢】,我想问一下一天,【一天】应该补充多少水分【啊】?就是运动,【运动】前后,应该怎么喝水【呢】?
 Mispronunciation 89 皮肤容易过敏起疹子,能介绍几款专业医疗级别的【修糊】产品吗?
@@
 Qwen2-Audio-Instruct 84.56 79.40 85.80 77.40 91.80 81.40 83.40 92.20
-StepAudio 87.17 87.40 90.40 73.40 92.80 87.80 81.80 96.60
+StepAudio 87.16 87.40 90.40 73.40 92.80 87.80 81.80 96.60
 StepAudio2Mini 82.79 82.60 82.20 77.60 90.60 87.60 65.40 94.40
-Mimo-Audio 91.21 92.20 94.00 85.00 99.00 92.80 76.60 99.20
+Mimo-Audio 91.22 92.20 94.00 85.00 99.00 92.80 76.60 99.20
 GPT4o-Audio 91.24 89.40 93.40 85.00 95.20 91.80 85.80 98.20
+Qwen3-Omni 95.59 98.70 98.20 93.80 99.90 98.67 80.70 99.70
 Table 8: Chinese Text Side Instruction Following Objective Results.
@@
 Mimo-Audio 72.89 81.80 74.40 54.00 47.60 88.80 88.40
-GPT4o-Audio 88.15 92.40 87.80 75.00 83.20 92.00 94.00
+GPT4o-Audio 88.16 92.40 87.80 75.00 83.20 92.00 94.00
+Qwen3-Omni 84.37 89.79 79.50 61.70 84.30 93.11 90.80
 Table 9: Chinese Speech Side Instruction Following Objective Results.
@@
 Baichuan-Audio-Chat 76.22 79.60 84.40 59.60 76.40 80.60 63.20 90.80
-Qwen2-Audio-Instruct 75.86 66.20 76.20 69.40 79.00 71.80 88.60 79.00
-StepAudio 66.92 71.60 72.00 51.60 68.00 69.80 50.60 85.60
+Qwen2-Audio-Instruct 75.89 66.20 76.20 69.40 79.00 71.80 88.60 79.00
+StepAudio 66.93 71.60 72.00 51.60 68.00 69.80 50.60 85.60
 StepAudio2Mini 75.54 73.40 78.20 57.60 80.60 78.80 73.20 87.80
@@
 GPT4o-Audio 91.66 91.40 94.60 86.60 95.40 94.60 81.60 98.20
+Qwen3-Omni 91.24 96.20 97.60 92.90 98.20 96.10 58.80 100.00
 Table 11: English Text Side Instruction Following Objective Results
 […]
-GLM4-Voice 78.52 82.80 68.00 64.40 81.40 84.60 88.20
-Kimi-Audio 61.87 69.20 54.00 50.00 61.40 67.40 68.40
+GLM4-Voice 78.53 82.80 68.00 64.40 81.40 84.60 88.20
+Kimi-Audio 61.86 69.20 54.00 50.00 61.40 67.40 68.40
 Qwen2.5-Omni 58.09 63.60 53.40 48.00 50.60 62.00 69.80
 Baichuan-Audio-Chat 68.07 77.20 57.60 47.60 73.00 77.20 73.80
-StepAudio 63.63 64.80 54.20 46.20 75.60 62.80 71.40
-StepAudio2Mini 65.15 73.00 47.20 52.40 78.80 70.20 68.00
+StepAudio 63.66 64.80 54.20 46.20 75.60 62.80 71.40
+StepAudio2Mini 65.14 73.00 47.20 52.40 78.80 70.20 68.00
 Mimo-Audio 24.25 27.40 24.80 22.60 20.40 22.40 26.80
 GPT4o-Audio 86.07 89.40 84.80 68.80 88.00 91.40 90.60
+Qwen3-Omni 77.50 85.43 71.20 57.92 80.40 87.57 81.10
 Table 12: English Speech Side Instruction Following Objective Results
 […]
-GLM4-Voice 85.13 92.20 87.20 76.00
-Kimi-Audio 85.67 92.40 89.60 75.00
-Qwen2-Audio-Instruct 85.67 93.20 91.00 72.80
-Qwen2.5-Omni 86.93 93.60 88.80 78.40
-Baichuan-Audio-Chat 73.27 80.00 74.60 65.20
-StepAudio2Mini 87.80 92.60 94.40 76.20
+GLM4-Voice 85.05 92.12 87.17 76.04
+Kimi-Audio 85.66 92.38 89.53 75.06
+Qwen2-Audio-Instruct 85.57 93.25 90,91 72.82
+Qwen2.5-Omni 86.85 93.69 88.81 78.39
+Baichuan-Audio-Chat 73.20 79.91 74.64 65.24
+StepAudio2Mini 87.72 92.58 94.44 76.14
+Qwen3-Omni 91.13 94.26 97.04 82.10
 Table 13: Multi-turn Dialogue Evaluation Results.
@@
 Mimo-Audio 56.58 58.33 43.33 45.76 60.87 60.00 40.54 71.57 65.33 62.34 66.67 34.43 57.33
-GPT4o-Audio 61.29 63.89 62.00 67.80 65.22 59.20 70.27 72.54 53.33 70.13 85.42 62.30 43.33
+GPT4o-Audio 61.29 63.89 62.00 67.80 65.22 59.20 70.27 72.55 53.33 70.13 85.42 62.30 43.33
+Qwen3-Omni 74.74 83.33 72.67 74.58 82.61 63.20 62.16 86.27 80.67 70.13 79.17 60.66 78.00
 Table 14: General Knowledge Evaluation Results.
@@
 Kimi-Audio 79.94 98.63 89.41 66.67 85.94 45.00 66.04
-Qwen2.5-Omni 80.24 95.89 78.82 71.43 78.13 47.50 81.76
+Qwen2.5-Omni 80.39 95.89 78.82 71.43 78.12 47.50 81.76
 Baichuan-Audio-Chat 60.33 69.18 52.94 45.24 70.31 25.00 72.96
-Qwen2-Audio-Instruct 60.78 86.30 57.06 39.29 51.56 22.50 59.75
-StepAudio 77.07 91.10 88.82 75.00 65.63 32.50 68.55
-StepAudio2Mini 81.30 97.26 88.82 65.48 71.88 45.00 79.87
-Mimo-Audio 84.01 94.52 88.82 72.62 78.13 52.50 85.53
+Qwen2-Audio-Instruct 59.28 86.30 57.06 39.29 51.56 22.50 59.75
+StepAudio 77.08 91.10 88.82 75.00 65.62 32.50 68.55
+StepAudio2Mini 81.29 97.26 88.82 65.48 71.88 45.00 79.87
+Mimo-Audio 84.01 94.52 88.82 72.62 78.12 52.50 85.53
 GPT4o-Audio 77.68 82.19 91.18 71.43 75.00 40.00 72.96
+Qwen3-Omni 94.27 97.95 98.82 91.67 90.62 62.50 96.86
 Table 15: Mathematical and Logical Reasoning Evaluation Results
@@
 Qwen2-Audio-Instruct 67.07 78.64 80.53 43.48
-Qwen2.5-Omni 73.72 83.49 79.65 59.13
+Qwen2.5-Omni 73.72 83.50 79.65 59.13
 Baichuan-Audio-Chat 54.38 58.25 71.68 33.91
@@
 GPT4o-Audio 77.64 90.29 84.96 59.13
+Qwen3-Omni 94.26 97.09 96.46 89.57
 Table 16: Discoure Comprehension Evaluation Results.
@@
 Avg. Child Elder Avg. Tianjin Beijing Dongbei Sichuan Avg. Down Up Avg.
-GLM4-Voice 65.00 (-17.20) 83.00 (-1.20) 47.60 (-32.40) 82.40 (0.20) 85.00 (-0.60) 82.60 (-2.60) 83.20 (3.20) 79.40 (0.00) 96.40 (0.60) 96.80 (0.80) 96.00 (0.40) 39.20 (-55.20)
-Kimi-Audio 45.60 (-30.80) 66.00 (-7.20) 25.80 (-53.80) 56.20 (-20.60) 54.40 (-23.80) 77.40 (18.80) 73.60 (-7.20) 36.60 (-44.00) 93.80 (-1.60) 93.20 (-2.00) 94.40 (-1.20) 87.80 (-4.40)
-Qwen2.5-Omni 75.60 (-5.40) 82.60 (1.40) 68.80 (-12.00) 77.80 (-3.40) 79.40 (-1.80) 81.40 (5.40) 74.40 (-8.00) 77.20 (-5.60) 90.40 (0.20) 93.20 (3.20) 87.60 (-2.80) 89.20 (-1.60)
-Baichuan-Audio-Chat 60.40 (-0.80) 61.80 (-0.40) 59.20 (-1.20) 61.60 (-0.60) 61.80 (-0.80) 62.60 (-5.40) 68.00 (3.20) 56.60 (-1.00) 77.20 (3.00) 83.60 (13.20) 70.80 (-7.20) 60.00 (-7.80)
-Qwen2-Audio-Instruct 76.40 (-5.80) 80.40 (-1.40) 72.60 (-10.00) 81.60 (0.80) 79.40 (0.60) 80.00 (-6.60) 83.20 (0.00) 83.40 (5.00) 93.20 (-1.40) 93.60 (-1.20) 92.80 (-1.60) 77.80 (-7.00)
-StepAudio2Mini 75.60 (-4.80) 75.40 (-2.00) 75.80 (-7.60) 75.40 (-2.00) 78.80 (-0.60) 88.00 (2.60) 67.20 (-4.80) 72.80 (-3.40) 96.60 (2.80) 96.00 (2.00) 97.20 (3.60) 76.00 (-14.40)
-Mimo-Audio 76.00 (-14.80) 92.80 (-0.40) 59.60 (-28.80) 88.40 (-0.20) 90.60 (2.40) 96.00 (10.60) 88.80 (1.60) 82.80 (-8.40) 99.20 (1.00) 98.80 (-0.40) 99.60 (2.40) 60.40 (-33.00)
-GPT4o-Audio 79.00 (-8.60) 89.00 (0.00) 69.20 (-17.00) 88.40 (2.20) 91.80 (4.20) 89.40 (0.00) 85.60 (0.80) 86.60 (2.20) 96.80 (-1.20) 97.20 (0.00) 96.40 (-2.40) 43.40 (-48.40)
+GLM4-Voice 65.00 (-17.20) 82.98 (-1.28) 47.50 (-32.50) 82.40 (0.20) 85.00 (-0.62) 82.67 (-2.66) 83.20 (3.20) 79.44 (0.00) 96.40 (0.60) 96.80 (0.80) 96.00 (0.40) 39.20 (-55.20)
+Kimi-Audio 45.60 (-30.80) 65.96 (-7.23) 25.83 (-53.74) 56.20 (-20.60) 54.38 (-23.74) 77.33 (18.66) 73.60 (-7.20) 36.67 (-43.89) 93.80 (-1.60) 93.20 (-1.90) 94.40 (-1.20) 87.80 (-4.40)
+Qwen2.5-Omni-7B 75.60 (-5.40) 82.55 (1.27) 68.75 (-12.08) 77.80 (-3.40) 79.38 (-1.87) 81.33 (5.33) 74.40 (-8.00) 77.22 (-5.56) 90.40 (0.20) 93.20 (3.20) 87.60 (-2.80) 89.20 (-1.60)
+Baichuan-Audio-Chat 60.40 (-0.80) 61.70 (-0.43) 59.17 (-1.26) 61.60 (-0.60) 61.88 (-0.62) 62.67 (-5.33) 68.00 (3.20) 56.67 (-1.11) 77.20 (3.00) 83.60 (13.20) 70.80 (-7.20) 60.00 (-7.80)
+Qwen2-Audio-7B-Instruct 76.40 (-5.80) 80.43 (-1.35) 72.50 (-10.00) 81.60 (0.80) 79.38 (0.63) 80.00 (-6.67) 83.20 (0.00) 83.33 (5.00) 93.20 (-1.40) 93.60 (-1.20) 92.80 (-1.60) 77.80 (-7.00)
+StepAudio 62.53 (-18.94) 77.87 (-5.53) 47.50 (-32.08) 71.48 (-3.71) 68.75 (3.75) 76.00 (-13.33) 74.40 (-0.80) 70.00 (-8.33) 95.80 (1.00) 96.00 (1.20) 95.60 (0.80) 38.26 (-54.78)
+StepAudio2 75.60 (-4.80) 75.32 (-2.13) 75.83 (-7.50) 75.40 (-2.00) 78.75 (-0.63) 88.00 (2.67) 67.20 (-4.80) 72.78 (-3.33) 96.60 (2.80) 96.00 (2.00) 97.20 (3.60) 76.00 (-14.40)
+Mimo-Audio 76.00 (-14.80) 92.77 (-0.42) 59.58 (-28.75) 88.40 (-0.20) 90.62 (2.50) 96.00 (10.67) 88.80 (1.60) 82.78 (-8.33) 99.20 (1.00) 98.80 (-0.40) 99.60 (2.40) 60.40 (-33.00)
+GPT4o-Audio 79.00 (-8.60) 88.94 (0.00) 69.17 (-17.08) 88.40 (2.20) 91.88 (4.38) 89.33 (0.00) 85.60 (0.80) 86.67 (2.23) 96.80 (-1.20) 97.20 (0.00) 96.40 (-2.40) 43.40 (-48.40)
+Qwen3-Omni 85.47 (-9.48) 96.60 (1.71) 74.58 (-20.42) 95.19 (1.12) 93.12 (3.74) 98.67 (2.67) 96.80 (0.80) 94.44 (-1.67) 98.80 (-0.80) 98.80 (-0.80) 98.80 (-0.80) 95.22 (-3.91)
 Table 17: Speaker Variations Evaluation Results - Experimental Group (Difference from Control Group).
@@
 Avg. Echo Outdoors Far Field Avg. Tv Playback Background Chat Vocal-Music Voice Announce Avg.
-GLM4-Voice 64.00 (-24.60) 45.00 (-36.80) 96.60 (-0.60) 74.80 (-12.60) 89.20 (-0.80) 85.60 (-3.80) 83.00 (-1.40) 93.00 (-0.80) 92.80 (1.60) 78.40 (-14.60)
-Kimi-Audio 74.00 (-5.40) 60.60 (-11.60) 94.80 (4.60) 84.00 (-1.20) 75.80 (-9.60) 76.80 (-8.20) 58.20 (-11.60) 93.80 (0.60) 70.20 (-19.80) 76.80 (-11.20)
-Qwen2.5-Omni 69.40 (-16.80) 52.00 (-30.60) 95.40 (0.60) 82.80 (-3.00) 85.20 (-4.80) 80.80 (-6.20) 74.80 (-8.20) 95.60 (-1.80) 85.60 (-4.40) 79.80 (-10.20)
-Baichuan-Audio-Chat 56.60 (-15.60) 45.00 (-20.00) 81.20 (-8.60) 58.20 (-12.60) 73.40 (-2.80) 73.40 (0.40) 64.40 (1.00) 81.40 (-5.40) 72.00 (-5.60) 67.20 (-4.60)
-Qwen2-Audio-Instruct 59.80 (-17.60) 37.80 (-32.00) 93.80 (-1.60) 76.00 (-1.20) 87.60 (-1.60) 82.20 (-6.00) 80.00 (-1.40) 96.20 (-0.20) 88.60 (0.40) 77.80 (-13.40)
-StepAudio2Mini 65.00 (-19.20) 45.80 (-33.00) 94.80 (-1.20) 78.80 (-6.40) 83.40 (-5.80) 84.00 (-2.20) 66.20 (-12.00) 94.80 (-2.60) 84.80 (-6.80) 81.60 (-8.40)
-Mimo-Audio 73.20 (-20.20) 54.00 (-35.60) 99.40 (-0.60) 90.80 (-4.60) 90.40 (-5.00) 87.00 (-9.00) 86.80 (-4.80) 99.60 (0.60) 85.60 (-8.20) 95.20 (-4.20)
-GPT4o-Audio 68.20 (-19.20) 61.00 (-20.60) 98.20 (-0.60) 54.80 (-34.40) 92.00 (-1.60) 86.20 (-6.60) 88.00 (-0.20) 96.40 (0.20) 94.20 (-1.60) 74.60 (-21.60)
+GLM4-Voice 64.00 (-22.60) 45.00 (-36.75) 96.57 (-0.57) 74.86 (-12.57) 89.20 (-0.80) 85.56 (-3.89) 82.94 (-1.47) 93.03 (-0.68) 92.82 (1.57) 78.40 (-14.60)
+Kimi-Audio 74.00 (-5.40) 60.50 (-11.65) 94.86 (4.57) 84.00 (-1.14) 75.80 (-9.60) 76.73 (-8.36) 58.24 (-11.47) 93.71 (0.45) 70.25 (-19.75) 76.80 (-11.20)
+Qwen2.5-Omni-7B 69.40 (-16.80) 52.00 (-30.50) 95.43 (0.57) 82.86 (-2.85) 85.20 (-4.80) 80.74 (-6.17) 74.71 (-8.23) 95.51 (-1.79) 85.57 (-4.43) 79.80 (-10.20)
+Baichuan-Audio-Chat 56.60 (-15.60) 45.00 (-20.00) 81.14 (-8.57) 58.29 (-12.57) 73.40 (-2.80) 73.45 (0.36) 64.41 (0.88) 81.35 (-5.39) 72.00 (-5.50) 67.20 (-4.60)
+Qwen2-Audio-7B-Instruct 59.80 (-17.60) 37.75 (-32.00) 93.71 (-1.72) 76.00 (-1.14) 87.60 (-1.60) 82.18 (-6.18) 80.00 (-1.49) 96.18 (-0.18) 88.50 (0.25) 77.80 (-13.40)
+StepAudio 55.87 (-30.80) 36.00 (-46.25) 93.71 (-0.58) 63.43 (-25.71) 86.92 (-2.94) 83.27 (-10.18) 77.65 (-1.47) 97.75 (0.45) 85.25 (-3.00) 66.50 (-27.25)
+StepAudio2 65.00 (-19.20) 45.75 (-33.00) 94.86 (-1.14) 78.86 (-6.28) 83.40 (-5.80) 84.00 (-2.18) 66.18 (-12.06) 94.83 (-2.47) 84.75 (-6.75) 81.60 (-8.40)
+Mimo-Audio 73.20 (-20.20) 54.00 (-35.50) 99.43 (4.43) 90.86 (-4.57) 90.40 (-5.00) 86.91 (-9.09) 86.76 (-5.00) 99.55 (0.45) 85.50 (-8.25) 95.20 (-4.20)
+GPT4o-Audio 68.20 (-19.20) 61.00 (-20.50) 98.29 (-0.57) 54.86 (-34.28) 92.00 (-1.60) 86.18 (-6.55) 87.94 (-0.30) 96.40 (0.22) 94.25 (-1.50) 74.60 (-21.60)
+Qwen3-Omni 83.47 (-13.86) 71.25 (-25.50) 97.71 (-1.72) 97.14 (0.57) 96.44 (-1.85) 96.73 (-1.45) 94.71 (-2.94) 100.00 (0.00) 93.75 (-3.25) 96.25 (-3.75)
 Table 18: Environmental Variations Evaluation Results - Experimental Group (Difference from Control Group).
 […]
-Fillers Repetition Mispronunciation Grammatical Error Topic Shift Code Switching
-GLM4-Voice 81.80 (-3.00) 89.20 (-0.60) 82.20 (-10.60) 85.60 (-6.00) 86.20 (-6.40) 80.80 (-10.80)
-Kimi-Audio 76.00 (-0.40) 79.40 (-1.20) 61.80 (-23.20) 77.40 (-10.40) 79.40 (-6.80) 84.00 (-8.80)
-Qwen2.5-Omni 79.00 (-2.00) 84.80 (0.00) 80.60 (-8.40) 79.20 (-8.60) 75.00 (-11.40) 80.40 (-9.20)
-Baichuan-Audio-Chat 69.60 (1.20) 72.00 (-0.20) 71.40 (-1.20) 65.60 (-11.00) 61.60 (-12.40) 68.40 (-7.60)
-Qwen2-Audio-Instruct 75.00 (1.60) 85.80 (-1.20) 78.40 (-10.00) 79.40 (-9.20) 67.00 (-19.80) 76.40 (-11.60)
-StepAudio2Mini 79.20 (-3.80) 81.60 (-2.60) 82.20 (-5.60) 80.80 (-8.20) 80.60 (-3.00) 83.60 (-8.60)
-Mimo-Audio 92.00 (-1.80) 95.00 (-0.40) 96.00 (-3.60) 97.40 (-0.60) 92.00 (-5.60) 92.60 (-7.00)
-GPT4o-Audio 88.40 (1.80) 93.20 (0.00) 92.60 (-4.00) 89.00 (-5.00) 90.60 (-2.60) 96.80 (-0.60)
+Avg. Casual Talk Mispronunciation Grammatical Error Topic Shift Code Switching
+GLM4-Voice 84.41 (-6.90) 85.81 (-4.04) 82.20 (-10.60) 85.60 (-6.00) 86.20 (-6.40) 80.80 (-10.80)
+Kimi-Audio 77.71 (-9.40) 81.18 (-4.44) 61.80 (-23.20) 77.40 (-10.40) 79.40 (-6.80) 84.00 (-8.80)
+Qwen2.5-Omni-7B 80.44 (-6.75) 83.25 (-2.17) 80.60 (-8.40) 79.20 (-8.60) 75.00 (-11.40) 80.40 (-9.20)
+Baichuan-Audio-Chat 67.46 (-7.02) 68.57 (-5.52) 71.40 (-1.20) 65.60 (-11.00) 61.60 (-12.40) 68.40 (-7.60)
+Qwen2-Audio-7B-Instruct 79.60 (-9.04) 87.29 (-2.56) 78.40 (-10.00) 79.40 (-9.20) 67.00 (-19.80) 76.40 (-11.60)
+StepAudio 85.81 (-5.64) 86.21 (-3.35) 89.44 (-2.69) 90.43 (-2.32) 82.42 (-7.78) 81.30 (-13.90)
+StepAudio2 83.04 (-4.79) 85.02 (-2.27) 82.20 (-5.60) 80.80 (-8.20) 80.60 (-3.00) 83.60 (-8.60)
+Mimo-Audio 95.10 (-3.04) 96.45 (-0.69) 96.00 (-3.60) 97.40 (-0.60) 92.00 (-5.60) 92.60 (-7.00)
+GPT4o-Audio 92.34 (-1.94) 92.12 (-0.29) 92.60 (-4.00) 89.00 (-5.00) 90.60 (-2.60) 96.80 (-0.60)
+Qwen3-Omni 97.61 (-1.62) 98.92 (-0.19) 94.61 (-4.04) 95.36 (-3.77) 96.92 (-2.42) 100.00 (0.00)
 Table 19: Content Variations Evaluation Results - Experimental Group (Difference from Control Group).

### v3

@@
 Abstract
-Recent advances in large audio language models (LALMs) have greatly enhanced multimodal conversational systems.
-However, existing benchmarks remain limited—they are mainly English-centric, rely on synthetic speech, and lack comprehensive, discriminative evaluation across multiple dimensions.
-To address these gaps, we present Voice Chat Bot Bench (VCB Bench)—a high-quality Chinese benchmark built entirely on real human speech.
-VCB Bench evaluates LALMs from three complementary perspectives: instruction following (including speech-level control beyond text commands), knowledge understanding (general knowledge, reasoning, and daily dialogue), and robustness (stability under perturbations in content, environment, and speaker traits).
-Experiments on representative LALMs reveal notable performance gaps and highlight future directions for improvement.
-VCB Bench provides a reproducible and fine-grained evaluation framework, offering standardized methodology and practical insights for advancing Chinese voice conversational models.
+While large audio language models (LALMs) have driven significant progress in multimodal conversational systems, current benchmarks suffer from critical limitations: they are largely English-centric, use synthetic speech, and fail to provide comprehensive, discriminative evaluation across key dimensions.
+To fill this gap, we present Voice Chat Bot Bench (VCB Bench), a novel, high-quality Chinese benchmark built exclusively on real human speech.
+VCB Bench assesses LALMs across three complementary axes: instruction following (including speech-level control beyond text commands), knowledge understanding (including general knowledge, reasoning, and daily dialogue), and robustness (evaluating stability under variations in content, environment, and speaker characteristics).
+Experiments conducted on representative LALMs reveal notable performance disparities and offer tangible insights for future improvements.
+VCB Bench serves as a reproducible and fine-grained framework, providing standardized evaluation and practical guidance for the development of Chinese voice conversational models.
 [^1] Code and data are available at https://github.com/Tencent/VCB-Bench
@@
 Integrating language modeling with modalities such as vision and audio (Radford et al., 2021; Singh et al., 2022) has further given rise to a new paradigm of multimodal learning.
-Within this trend, large audio language models (LALMs)—which combine speech signal processing with language modeling—have developed rapidly.
-Emerging systems such as StepAudio2 (Wu et al., 2025) and Qwen3-Omni (Xu et al., 2025b) demonstrate end-to-end (E2E) speech understanding and generation with capabilities in voice question answering, real-time conversation, and audio content analysis.
+Within this trend, LALMs—which combine speech signal processing with language modeling—have developed rapidly.
+Emerging systems such as Step-Audio 2 (Wu et al., 2025) and Qwen3-Omni (Xu et al., 2025b) demonstrate end-to-end (E2E) speech understanding and generation with capabilities in voice question answering, real-time conversation, and audio content analysis.
 Consequently, voice conversational agents powered by LALMs are drawing increasing academic and industrial attention, offering more natural and human-like interactions than text-only systems.
-Despite these advances, moving from basic LALM functionalities to practical voice agents requires reliable and comprehensive evaluation tools.
-Such benchmarks are essential for diagnosing model weaknesses, guiding optimization, and enabling fair comparisons across systems.
-While initial efforts (Chen et al., 2024; Yang et al., 2024; Lin et al., 2025) have explored instruction following, audio understanding, reasoning, and dialogue scenarios, current evaluation practices remain limited in three major ways.
-First, most benchmarks are English-centric, leaving Chinese—the world’s most widely spoken language—largely unexplored.
-Second, the majority rely on synthetic speech data, which poorly reflects real-world acoustic variability.
-Third, many are text-derived benchmarks (e.g., AlpacaEval (Li et al., 2023), IFEval (Zhou et al., 2023)), whose formal and lengthy content is unsuitable for evaluating conversationally grounded LALMs that should generate natural, colloquial speech.
-Addressing these limitations is critical given China’s large user base and the growing demand for practical, high-quality voice agents.
-To bridge these gaps, we introduce Voice Chat Bot Bench (VCB Bench)—the first comprehensive evaluation framework for Chinese voice conversation, built entirely from authentic (non-synthetic) speech.
-VCB Bench evaluates LALMs along three complementary dimensions: (1) Instruction following, extending beyond text-based prompts to incorporate speech-level control tasks such as adjusting volume, speed, and emotion, with bilingual (Chinese-English) support; (2) Knowledge, including multi-disciplinary general knowledge (12 subjects), mathematical and logical reasoning, daily dialogue comprehension, and story continuation for pretraining performance assessment; (3) Robustness, measuring model stability under real-world perturbations across content (mispronunciations, grammatical errors), environment (street, TV noise), and speaker characteristics (age, accents).
-Our proposed VCB Bench is built entirely from authentic human recordings rather than synthetic speech.
-It provides a large-scale, high-fidelity dataset covering diverse conversational scenarios and introduces a multi-dimensional evaluation framework that jointly measures knowledge understanding, instruction following, and robustness through fine-grained, reproducible tasks.
-Based on this benchmark, we conduct a systematic empirical analysis of state-of-the-art LALMs under unified settings, revealing their strengths and limitations in Chinese voice interaction and offering actionable insights for future model development.
+Despite these advances, the transition from basic LALM functionalities to practical voice agents necessitates reliable and comprehensive evaluation tools.
+Such benchmarks are essential for diagnosing model weaknesses, guiding optimization efforts, and enabling fair comparisons across different systems.
+While initial efforts (Chen et al., 2024; Yang et al., 2024; Lin et al., 2025) have explored instruction following, audio understanding, reasoning, and dialogue scenarios, current evaluation practices remain limited in three major aspects: First, existing benchmarks are predominantly English-centric, resulting in insufficient exploration of Chinese, despite it being the world’s most widely spoken language.
+Second, the majority rely on synthetic speech data, which poorly reflects the acoustic variability of real-world environments.
+Third, many are derived from text-based benchmarks (e.g., AlpacaEval (Li et al., 2023), IFEval (Zhou et al., 2023)).
+The formal and lengthy content of these benchmarks is unsuitable for evaluating conversationally grounded LALMs, which are designed to generate natural, colloquial speech.
+This effort is critical due to China’s vast user base and the rapidly growing demand for practical, high-quality voice agents.
+We address these limitations by introducing VCB Bench: the first comprehensive evaluation framework for Chinese voice conversation, built entirely upon authentic (non-synthetic) speech.
+VCB Bench assesses LALMs across three complementary dimensions: (1) Instruction Following: This extends beyond text-based prompts to include speech-level control tasks, such as adjusting volume, speed, and emotion, with inherent bilingual (Chinese-English) support.
+(2) Knowledge: This encompasses multi-disciplinary general knowledge (across 12 subjects), mathematical and logical reasoning, daily dialogue comprehension, and story continuation to assess pre-training performance.
+(3) Robustness: This measures the model’s stability under real-world perturbations, including variations in content (mispronunciations, grammatical errors), environment (street, TV noise), and speaker characteristics (age, accents).
+Our proposed VCB Bench is entirely derived from authentic human recordings, moving beyond the limitations of synthetic speech.
+It provides a large-scale, high-fidelity dataset encompassing diverse conversational scenarios and introduces a multi-dimensional evaluation framework.
+This framework jointly measures instruction following (including fine-grained control), knowledge understanding, and robustness through a set of reproducible tasks.
+Based on this benchmark, we conduct a systematic empirical analysis of state-of-the-art LALMs under unified settings, which elucidates their strengths and limitations in Chinese voice interaction and yields actionable insights for future model development.
 2 Related Work
-Large Audio Language Models. Recent LALMs primarily adopt an E2E audio-language modeling paradigm, integrating speech understanding and generation within a unified framework.
+Large Audio Language Models.
+LALMs primarily adopt an E2E audio-language modeling paradigm (Tang et al., 2023; Fu et al., 2024; Xie and Wu, 2024), integrating speech understanding and generation within a unified framework.
 Qwen-Audio and Qwen-Omni series (Chu et al., 2023, 2024; Xu et al., 2025a, b) progressively enhance cross-modal alignment and modeling efficiency.
@@
 StepAudio models (Huang et al., 2025; Wu et al., 2025) focus on tightly coupling recognition and synthesis.
-StepAudio integrates a dual-codebook tokenizer and achieves a remarkably low WER, while StepAudio2 advances to a fully E2E design with fixed text-speech token alignment and Chain-of-Thought reasoning, improving fine-grained paralinguistic understanding.
+StepAudio integrates a dual-codebook tokenizer and achieves a remarkably low WER, while Step-Audio 2 mini advances to a fully E2E design with fixed text-speech token alignment and Chain-of-Thought reasoning, improving fine-grained paralinguistic understanding.
 Baichuan-Audio (Li et al., 2025) employs hierarchical RVQ discretization and dual audio heads to balance acoustic and linguistic objectives, enabling real-time bilingual communication.
-GLM4-Voice (Zeng et al., 2024) introduces a three-module structure (Tokenizer-Backbone-Decoder) supporting emotion and dialect modeling.
+GLM-4-Voice (Zeng et al., 2024) introduces a three-module structure (Tokenizer-Backbone-Decoder) supporting emotion and dialect modeling.
 Kimi-Audio (Ding et al., 2025) fuses continuous acoustic and discrete semantic tokens in a dual-head architecture, achieving low-latency, high-fidelity streaming generation.
-These models demonstrate rapid progress in unified audio-language modeling—covering tokenization, multimodal fusion, and real-time dialogue—but systematic benchmarks, especially for Chinese real-speech interaction, remain scarce.
+In addition, there are many other excellent LALMs, such as MiMo-Audio (Xiaomi, 2025) and Fun-Audio-Chat(Chen et al., 2025).
+All of these models demonstrate rapid progress in unified audio-language modeling—covering tokenization, multimodal fusion, and real-time dialogue—but systematic benchmarks, especially for Chinese real-speech interaction, remain scarce.
 Current evaluations are mostly qualitative or based on synthetic data, underscoring the need for comprehensive real-speech benchmarks like VCB Bench.
@@
 Recent efforts have introduced several benchmarks to evaluate LALMs from different perspectives.
-VoiceBench (Chen et al., 2024) assesses general knowledge, instruction adherence, safety, and robustness, mainly based on existing text datasets such as AlpacaEval and SD-QA (Faisal et al., 2021), with lengthy or highly complex samples removed.
-OpenAudioBench, released alongside Baichuan-Audio, integrates question-answering datasets including Spoken LLaMA Questions (Nachmani et al., 2023) and Web Questions (Berant et al., 2013), and augments them with a TTS-generated reasoning subset.
-AIR Bench (Yang et al., 2024) contains two components—a basic benchmark covering emotion recognition, ASR, and age estimation, and a dialogue benchmark evaluating auditory understanding and internal knowledge.
-MMAU (Kumar et al., 2025) and MMAR (Ma et al., 2025) focus on deep audio reasoning, requiring multi-step inference grounded in internal audio knowledge.
-OmniBench (Li et al., 2024) targets omni-modal models handling audio, images, and text, where text queries are paired with multimodal contexts (speech, music, or sound) to test integrated reasoning.
-Finally, URO Bench (Yan et al., 2025) provides a bilingual (English-Chinese) comprehensive set that evaluates audio understanding, reasoning, and conversational ability—but the speech data are entirely synthetic (TTS-generated).
+VoiceBench (Chen et al., 2024) assesses general knowledge, instruction adherence, safety, and robustness, primarily by adapting content from existing text datasets such as AlpacaEval and SD-QA (Faisal et al., 2021), often after removing lengthy or overly complex samples.
+OpenAudioBench, released with Baichuan-Audio, integrates question-answering datasets like Spoken LLaMA Questions (Nachmani et al., 2023) and Web Questions (Berant et al., 2013), and further augments them with a TTS-generated reasoning subset.
 Figure 1: Overview of VCB Bench.
+AIR Bench (Yang et al., 2024) comprises two components: a basic benchmark covering emotion recognition, automatic speech recognition (ASR), and age estimation, and a dialogue benchmark focused on evaluating auditory understanding and internal knowledge.
+MMAU (Kumar et al., 2025) and MMAR (Ma et al., 2025) concentrate on deep audio reasoning, specifically requiring multi-step inference grounded in the model’s internal audio knowledge.
+OmniBench (Li et al., 2024) targets omni-modal models that process audio, images, and text.
+It utilizes text queries paired with multimodal contexts (speech, music, or sound) to test integrated reasoning capabilities.
+Finally, URO Bench (Yan et al., 2025) provides a comprehensive bilingual (English-Chinese) set for evaluating audio understanding, reasoning, and conversational ability, but its speech data are entirely synthetic (TTS-generated).
 Overall, existing benchmarks have significantly advanced the evaluation coverage of LALMs, yet they share several limitations: (1) most rely heavily on TTS or synthetic speech, (2) they focus on English, and (3) their content often derives from text-centric QA corpora rather than spontaneous human dialogue.
@@
 As shown in the Figure 1, VCB Bench covers three core dimensions: Instruction Following, Knowledge, and Robustness.
-Instruction Following includes Text Instruction Follow (TIF) (e.g., continuation, creation), Speech Instruction Follow (SIF) (e.g., emotional, volume control), and Multi-turn Dialog (MTD) tasks.
+Instruction Following includes Text Instruction Following (TIF) (e.g., continuation, creation), Speech Instruction Following (SIF) (e.g., emotional, volume control), and Multi-turn Dialog (MTD) tasks.
 The Knowledge module assesses General Knowledge (GK) across 12 disciplines, Mathematical Logic (ML), Discourse Comprehension (DC), and Story Continuation (SC).
 Robustness introduces real-world perturbations from speaker variations, environmental noise, and content modifications to evaluate model stability.
+The detailed statistics for each subtask is presented in Appendix 3.
 3.1 Dataset Construction
-The VCB Bench dataset integrates data from three distinct sources: third-party professional recordings, audio extracted from variety show Q&A segments and an internally curated two-person conversational dialogue dataset.
-Each source supports different evaluation modules within the benchmark.
+The VCB Bench dataset integrates data from three distinct sources: third-party professional recordings, audio segments extracted from variety show Q&A, and an internally curated two-person conversational dialogue dataset.
+Each source is utilized to support different evaluation modules within the benchmark.
 Third-Party Recorded Data.
-This category supports the Instruction Following, ML, and SC tasks under the Knowledge module, as well as the Robustness module.
-The production pipeline involves the following steps: First, task types and examples are defined through team discussion.
-Next, commissioned personnel manually compose texts that fulfill the task requirements.
-These texts then undergo manual inspection to ensure quality.
-Approved texts are forwarded to a third-party recording team for professional audio production.
-After recording, the data team performs quality checks on the audio.
-Subsequently, GPT-4o-Audio is used to evaluate audio quality, while GPT-4o assesses textual quality.
-Finally, manual screening is conducted to select high-quality samples, determining the final evaluation dataset.
-For Robustness data, it’s text materials are derived from Instruction Following module.
-The original audio from this module serves as the control group.
+This data category supports the Instruction Following, ML, and SC tasks in the Knowledge module, as well as the Robustness module.The production pipeline has several steps: First, task types and specific examples are defined.
+Next, professional writers manually craft task-compliant text.
+This text undergoes rigorous manual quality checks before being sent to a third-party team for professional audio recording.
+After recording, the data team first checks audio quality to filter out low-quality files.
+They then conduct detailed manual screening, selecting top-quality samples to form the final evaluation dataset.
+The text materials for the Robustness data are derived directly from the Instruction Following module, with the original audio from that module serving as the control group.
 To control for speaker variability, the same speaker re-recorded the text under specified interference conditions (e.g., accent, noisy environment) wherever possible, using the original audio as a baseline.
-For "content variation" types, the text was first modified (e.g., introducing grammatical or pronunciation errors) before being re-recorded by the same speaker.
-Additionally, to test performance in extreme scenarios, subsets like Volume, Speed, and Unstable Signal underwent post-processing.
+For content variation types, the text was first modified (e.g., introducing grammatical or pronunciation errors) before being re-recorded by the same speaker.
+Additionally, to test performance in extreme scenarios, subsets such as Volume, Speed, and Unstable Signal were created using post-processing techniques.
 Variety Show Q&A Data.
 This category supports GK in the Knowledge module.
-The process includes: crawling about 20 hours of Q&A audio; manual annotation and segmentation for timestamps; ASR transcription; quality scoring (assessing question clarity and answer accuracy) and subject classification, both via GPT-4O; selection of Q&A pairs above a score threshold; and a final manual review for transcription accuracy, answer correctness, and audio clarity.
+The process involves: crawling about 20 hours of Q&A audio; manual timestamp annotation and segmentation; ASR transcription; question subject classification; and a final manual review for transcription accuracy, audio clarity, problem statement and answer correctness.
 Figure 2: Examples from the VCB Bench.
 Internal Two-person Dialogue Dataset.
-Designed to support the DC module, this category’s data is processed as follows: the original long-form audio undergoes a two-stage segmentation with GPT-4O—first by topic, then refined into semantically coherent segments under one minute.
-From the transcriptions, GPT-4O generates task-specific QA pairs (e.g., for analysis or induction), followed by a final manual screening to verify question quality and answer accuracy.
+Designed to support the DC module, this category’s data is processed as follows: the original long-form audio undergoes a two-stage segmentation —first by topic, then refined into semantically coherent segments under one minute.
+From the ASR transcriptions corresponding to the segments, we ask LLM to generate task-specific QA pairs (e.g., for analysis or induction), followed by a final manual screening to verify question quality and answer accuracy.
 3.2 Dataset Details
 Instruction Following.
-The Instruction Following section comprehensively evaluates LALMs’ ability to understand and execute both text and speech instructions, covering three sub-tasks: TIF, SIF, and MTD.
-All tasks are open-ended, and both TIF and SIF support Chinese and English to meet cross-lingual evaluation needs.
+The Instruction Following section provides a comprehensive evaluation of LALMs’ capacity to understand and execute both text and speech commands, encompassing three sub-tasks: TIF, SIF, and MTD.
+All tasks are open-ended, and both TIF and SIF support bilingual evaluation (Chinese and English) to address cross-lingual assessment requirements.
 TIF assesses the model’s ability to respond to textual instructions through seven sub-tasks, each examining text generation and semantic comprehension: (1) Continuation: extending a given text fragment to evaluate coherence and creativity; (2) Creation: generating original content based on a given theme to assess inventiveness and organization; (3) Empathy: understanding and responding to emotional expressions to examine affective perception; (4) Recommendation: providing suggestions based on user needs to evaluate information integration; (5) Rewriting: adapting text in style or structure to test reorganization ability; (6) Safety: identifying and rejecting harmful instructions to assess compliant response; (7) Simulation: role-playing in dialogue to examine contextual adaptation.
@@
 It includes six sub-tasks: (1) Emotional Control: adjusting the emotional tone of speech to assess expressive generation; (2) Language Control: switching languages or dialects to test multilingual synthesis; (3) Non-verbal Vocalization: incorporating non-linguistic elements like sighs or nasal sounds to evaluate paralinguistic expressiveness; (4) Pacing Control: modifying speaking rate to examine control precision; (5) Style Control: switching speech styles to assess style transfer; (6) Volume Control: adjusting loudness to test stability.
-Model Instruction Following Knowledge Robustness
-TIF TIF-En SIF SIF-En MTD GK ML DC SV EV CV
-GLM4-Voice (Zeng et al., 2024) 85.82 82.52 85.57 78.53 85.05 45.53 62.14 48.64 76.09 80.25 84.41
-Kimi-Audio (Ding et al., 2025) 85.13 88.92 85.69 61.86 85.66 53.51 79.94 74.76 68.25 75.44 77.71
-Qwen2.5-Omni (Xu et al., 2025a) 87.40 72.58 71.37 58.09 86.85 55.43 80.39 73.72 82.31 79.81 80.44
-Baichuan-Audio-Chat (Li et al., 2025) 72.49 76.22 78.96 68.07 73.20 44.48 60.33 54.38 65.53 67.62 67.46
-Qwen2-Audio-Instruct (Chu et al., 2024) 84.56 75.89 / / 85.57 35.83 59.28 67.07 83.01 78.11 79.60
-StepAudio (Huang et al., 2025) 87.16 66.93 80.25 63.66 / 60.42 77.08 59.52 71.63 74.87 85.81
-StepAudio2Mini (Wu et al., 2025) 82.79 75.54 78.81 65.14 87.72 61.15 81.29 83.08 81.61 77.84 83.04
-Mimo-Audio 91.22 91.76 72.89 24.25 / 56.58 84.01 87.92 84.43 86.19 95.10
-GPT4o-Audio 91.24 91.66 88.16 86.07 / 61.29 77.68 77.64 82.32 82.49 92.34
-Qwen3-Omni (Xu et al., 2025b) 95.59 91.24 84.37 77.50 91.13 74.74 94.27 94.26 93.58 92.68 97.61
-Table 1: The main results of different LALMs on VCB Bench. Missing results from unsupported modalities or API unavailability.
 MTD evaluates instruction tracking and topic management in multi-turn dialogues, each containing 3-5 turns, focusing on contextual understanding and logical coherence: (1) Progression: deepening the discussion around an initial topic to assess topic development; (2) Backtracking: recalling and responding to previously mentioned information to test long-range memory; (3) Transition: suddenly shifting to a new topic to evaluate conversational flow and relevance.
@@
 This task spans three evaluative categories: Logic and Causality for causal consistency, Common Sense and Science for real-world and scientific knowledge alignment, and Morality and Emotion for moral and emotional coherence.
+Model Instruction Following Knowledge Robustness
+TIF TIF-En SIF SIF-En MTD GK ML DC SV EV CV
+GLM-4-Voice (Zeng et al., 2024) 82.15 75.52 73.18 66.94 82.56 41.79 60.18 / 73.64 77.51 78.60
+Kimi-Audio (Ding et al., 2025) 77.33 60.37 71.04 54.13 82.27 54.47 75.42 51.96 64.35 65.14 71.51
+Qwen2.5-Omni (Xu et al., 2025a) 76.75 60.83 54.04 42.18 85.23 48.13 73.00 72.81 77.82 74.41 78.27
+Baichuan-Audio (Li et al., 2025) 82.24 78.60 53.67 48.27 80.96 40.73 74.05 51.66 78.05 77.62 78.71
+Step-Audio 2 mini (Wu et al., 2025) 76.10 70.47 60.80 49.85 82.31 49.86 72.85 79.46 75.19 70.58 74.78
+MiMo-Audio (Xiaomi, 2025) 90.08 81.89 56.26 42.74 86.30 48.70 81.75 / 83.72 85.36 89.38
+GPT-4o-Audio (OpenAI, 2024) 86.94 88.80 77.98 82.90 33.59 55.81 73.45 76.74 80.34 79.92 86.51
+Qwen3-Omni (Xu et al., 2025b) 90.45 85.17 70.73 65.57 87.17 66.86 81.90 82.78 87.91 85.63 86.03
+Fun-Audio-Chat (Chen et al., 2025) 89.30 89.39 78.82 75.70 / 46.30 73.00 79.15 88.60 83.83 85.15
+Table 1: The main results of different LALMs on VCB Bench. Missing results from unsupported modalities or API unavailability.
 Robustness.
@@
 4.1 Configuration
-Figure 3: The robustness of LALMs under real-world perturbations. 9 subsets with the most significant performance gaps compared to the control group on the Robustness dataset are chosen.
 We evaluate the latest and most capable LALMs.
-The selected models comprise GLM4-Voice, Kimi-Audio, Qwen2.5-Omni, Baichuan-Audio-Chat, Qwen2-Audio-Instruct, StepAudio, StepAudio2Mini, Mimo-Audio, GPT4o-Audio and Qwen3-Omni [^2] Qwen3-Omni was included solely in the experimental comparison in Section 4.2..
-For the SIF tasks in both Chinese and English (SIF-En), we invoke each model’s “audio2audio” API to generate spoken responses.
-The adherence to instruction requirements is then automatically scored using GPT-4o-audio.
-For other tasks except SC, we call the “audio2text” API to obtain textual responses, which are evaluated by GPT-4O.
-In open-ended question answering, GPT-4O provides a numerical score on a 1-5 scale, while for reference-based QA, it returns a binary "Yes" or "No" judgment.
-For the SC task, we assess a subset of pre-trained base models: Baichuan-Audio-Base, Kimi-Audio-Base, Qwen2-Audio-Base, and StepAudio2Mini-Base.
+The selected models comprise GLM-4-Voice, Kimi-Audio, Qwen2.5-Omni, Baichuan-Audio, Step-Audio 2 mini, MiMo-Audio, GPT-4o-Audio, Qwen3-Omni and Fun-Audio-Chat.
+For all tasks except SC, we generate spoken responses by calling each model’s audio-to-audio API.
+For SIF tasks in both Chinese and English, the audio responses are evaluated directly using Gemini-2.5-Pro.
+For the remaining tasks, following prior work (Yan et al., 2025), we first transcribe the audio responses using an ASR model (Whisper (Radford et al., 2023) for English tasks and Paraformer (Gao et al., 2022) for Chinese tasks) and then assess the transcriptions with GPT-4o.
+In open-ended QA, both Gemini and GPT assign a numerical score on a 1–5 scale, whereas for reference-based QA, they output a binary judgment of "Yes" or "No."
+For the SC task, we assess a subset of pre-trained base models: Baichuan-Audio-Base, Kimi-Audio-Base, Qwen2-Audio-Base, and Step-Audio 2 mini-Base.
 Following the StoryCloze evaluation protocol, we compute the negative log-likelihood for both the correct and incorrect endings, with model selection determined by comparing these two values.
-For SIF tasks, the top six performing models undergo further Mean Opinion Score (MOS) evaluation.
+For SIF tasks, six performing models undergo further Mean Opinion Score (MOS) evaluation.
 We sample the first 30 items from each relevant dataset, and eight expert evaluators rate the generated audio samples.
-In the MTD evaluation, the model receives input context only in audio.
-We adopt Bai et al.
+In the MTD evaluation, we adopt Bai et al.
 (2024)’s protocol, requiring the model to answer each dialogue turn using the original ground-truth context, not its prior responses.
@@
 All Experiments are conducted on H20.
-4.2 Main Result
+Figure 3: The robustness of LALMs under real-world perturbations. 9 subsets with the most significant performance gaps compared to the control group on the Robustness dataset are chosen.
+4.2 Main Results
 Table 1 presents the average scores of each model across various test sets in VCB Bench, with the detailed scores for all subsets provided in the Appendix.
 As shown, Qwen3-Omni emerges as a new state-of-the-art (SOTA) model, demonstrating all-round superiority across most tasks.
-GPT-4o-Audio, as a strong non-open-source model, also performs robustly, particularly in SIF.
-Figure 4: The investigation of the text-speech alignment capability of LALMs.
-A2T (Audio-to-Text, generating text directly from audio input), A2A W/ ASR (Audio-to-Audio, transcribing the generated audio via Automatic Speech Recognition to text), and A2A W/O ASR (evaluating the generated audio’s text without ASR).
+In Instruction Following, Qwen3-Omni (90.45) and MiMo-Audio (90.08) excel in TIF, while Fun-Audio-Chat leads in TIF-En (89.39), closely trailing GPT-4o-Audio (88.80).
+For SIF, Fun-Audio-Chat (78.82) and GPT-4o-Audio (77.98) lead; GPT-4o-Audio dominates SIF-En (82.90), with most open-source models like Step-Audio 2 mini (49.85) lagging, reflecting challenges in English speech paralinguistic features.
+In MTD, Qwen3-Omni (87.17) and MiMo-Audio (86.30) lead, while GPT-4o-Audio performs poorly (33.59), highlighting divergence in long-context dialogue control.
+In Knowledge, Qwen3-Omni stands out with top scores in GK (66.86), ML (81.90) and DC (82.78), showing superior multi-disciplinary knowledge and reasoning.
+MiMo-Audio performs strongly in ML (81.75), comparable to Qwen3-Omni.
+However, Baichuan-Audio (GK:40.73, ML:74.05), GLM-4-Voice (GK:41.79, ML:60.18) and other models lag, revealing gaps in knowledge coverage and step-by-step reasoning.
+Overall, Qwen3-Omni demonstrates comprehensive superiority.
+Other models have task-specific strengths: Fun-Audio-Chat in cross-lingual TIF and SIF, MiMo-Audio in TIF and ML.
+Most open-source E2E LALMs face English adaptation challenges, while GPT-4o-Audio lacks long-context dialogue capability.
+4.3 Real Scenario
+As shown in Figure 3, EV.Echo, SV.Speed, and SV.Elder cause the most severe performance degradation for most models.
+For instance, the score of GLM-4-Voice drops from over 80 in the control group to below 40 under these conditions, indicating that speaker specificity and acoustic echo remain the most challenging perturbations for current LALMs.
+In contrast, content-level variations such as CV.Gram.Err and CV.Mispron exhibit relatively mild impacts.
+Models including Qwen3-Omni and Fun-Audio-Chat show only small score gaps between these subsets and the control group, suggesting models are more tolerant of “content-level flaws” than “speech/environment-level physical perturbations.”
+Regarding model robustness, Qwen3-Omni and Fun-Audio-Chat demonstrate relatively high robustness, achieving both high clean-condition scores and limited performance degradation across variations.
+In contrast, models such as GLM-4-Voice face constraints due to their lower baseline and pronounced sensitivity to physical perturbations, reflecting limited real-world adaptability despite moderate performance gaps under content-level variations.
+4.4 Pretraining Evaluation
+The SC task evaluates pre-trained LALMs’ intelligence and cross-modal semantic coherence by judging the rationality of story endings.
+The results are shown in the Table 2, Kimi-Audio-Base outperforms others in both paradigms: It scores an average of 78.01 in A->T and 54.71 in A->A, with robust performance across sub-dimensions, demonstrating stable story understanding and ending judgment in cross-modal scenarios.
+In contrast, Baichuan-Audio-Base, Qwen2-Audio-Base, and Step-Audio 2 mini-Base score much lower.
+Moreover, all models perform worse in A->A than A->T, revealing that cross-modal (speech-to-speech) story coherence judgment remains challenging for pre-trained LALMs, with notable room for improvement in semantic consistency and rationality generation during speech output.
+Figure 4: The Investigation of the Text-Speech Alignment Capability of LALMs: A2T (in which only text is generated from audio input, with the text directly evaluated), A2A W/ ASR (in which both audio and text are generated, with the audio evaluated via its transcription), and A2A W/O ASR (in which both audio and text are generated, with the text directly evaluated).
 All test sets are from TIF/TIF-En (e.g., “Creat.” for Creation, “Recom.” for Recommendation).
-In Instruction Following, Mimo-Audio excels in TIF (91.22) and TIF-En (91.76) with scores close to GPT-4o-Audio, indicating strong cross-lingual text instruction adaptation.
-For SIF, GPT-4o-Audio leads in Chinese SIF (88.16) and SIF-En (86.07), while StepAudio series and Kimi-Audio perform robustly in Chinese SIF but their SIF-En scores lag significantly, reflecting challenges in handling English speech’s paralinguistic features.
-In MTD, StepAudio2Mini (87.72) and Qwen3-Omni (91.13) lead among models, outperforming counterparts like Baichuan-Audio-Chat (73.20), which highlights divergence in long-context dialogue logic control.
-In Knowledge, Qwen3-Omni stands out in GK (74.74), ML (94.27), and DC (94.26), surpassing other models and demonstrating strengths in multi-disciplinary knowledge and deep reasoning.
-Mimo-Audio also performs strongly in ML (84.01) and DC (87.92), suggesting capabilities in deep reasoning and text semantic analysis.
-However, models like Baichuan-Audio-Chat show limited performance in GK (44.48) and ML (60.33), revealing gaps in multi-disciplinary knowledge coverage and step-by-step reasoning.
-Overall, Qwen3-Omni demonstrates comprehensive superiority, while other open-source E2E LALMs exhibit task-specific strengths (e.g., Mimo-Audio in reasoning, StepAudio2Mini in Chinese dialogue and general knowledge) but face challenges in cross-lingual speech adaptation.
-4.3 Real Scenario
-As shown in Figure 3, EV.Echo, SV.Speed and SV.Elder cause the most severe performance degradation for most models.
-Scores of some models drop from over 80 in the control group to below 40 in these subsets, indicating that speech rate variation and acoustic echo are the most challenging perturbations for current LALMs.
-However, CV-related interferences (e.g., CV.Gram.Err, CV.Mispron) have relatively mild impacts.
-Some models (e.g., Mimo-Audio, StepAudio2Mini) show small score gaps between these subsets and the control group, suggesting models are more tolerant of “content-level flaws” than “speech/environment-level physical perturbations”.
-Regarding model robustness, GPT-4o-Audio maintains excellent capability despite significant drops in specific subsets like SV.Speed, attributable to its high baseline scores ensuring practical usability.
-Among open-source models, Mimo-Audio and StepAudio2Mini exhibit relatively prominent robustness with both high absolute scores and limited performance gaps.
-In contrast, models like Baichuan-Audio-Chat face constraints primarily due to their lower absolute scores rather than extreme fluctuations, indicating insufficient real-scene adaptability despite moderate performance drops.
-4.4 Pretraining Evaluation
+4.5 Ablation Study
+4.5.1 Text-Speech Alignment
+To investigate the text-speech alignment capability of LALMs, we conduct an ablation study, which is shown in Figure 4.
+The visualization is based on two selection criteria from TIF and TIF-En: the four performing models with distinct differences between their A2T and A2A APIs, and the four datasets with the largest mean score differences between A2A W/ ASR and A2T.
+Results for Chinese and English tasks are plotted separately in the upper and lower sections of the figure, respectively.
 Model Task Metrics
@@
 A -> A 36.91 36.78 44.90 31.94
-StepAudio2Mini-Base A -> T 50.26 52.87 26.53 56.94
+Step-Audio 2 mini-Base A -> T 50.26 52.87 26.53 56.94
 A -> A 30.63 27.55 34.69 38.89
 Table 2: Pretraining Evaluation Results on SC.
-The SC task evaluates pre-trained LALMs’ intelligence and cross-modal semantic coherence by judging the rationality of story endings.
-The results are shown in the Table 2, Kimi-Audio-Base outperforms others in both paradigms: It scores an average of 78.01 in A->T and 54.71 in A->A, with robust performance across sub-dimensions, demonstrating stable story understanding and ending judgment in cross-modal scenarios.
-In contrast, Baichuan-Audio-Base, Qwen2-Audio-Base, and StepAudio2Mini-Base score much lower.
-Moreover, all models perform worse in A->A than A->T, revealing that cross-modal (speech-to-speech) story coherence judgment remains challenging for pre-trained LALMs, with notable room for improvement in semantic consistency and rationality generation during speech output.
-4.5 Ablation Study
-4.5.1 Text-Speech Alignment
-To investigate the text-speech alignment capability of LALMs, we conduct an ablation study, which is shown on Figure 4.
-The visualization is based on two selection criteria from TIF and TIF-En: the four models with the highest A2A W/ ASR scores, and the four datasets with the largest mean score differences between A2A W/ ASR and A2T.
-Results for Chinese and English tasks are plotted separately in the upper and lower sections of the figure, respectively.
-From the results, models like GLM4-Voice (Chinese) and Baichuan-Audio-Chat (English) demonstrate strong text-speech alignment—their A2T results are close to A2A W/ ASR results, indicating consistent semantic output between directly generated text and text transcribed from speech.
-In contrast, models such as Qwen2.5-Omni (Chinese) and Kimi-Audio (English) show large discrepancies between A2T and A2A W/ ASR, suggesting mismatches in semantics between text and speech generation.
-Meanwhile, for audio generation quality (assessed by the gap between A2A W/ ASR and A2A W/O ASR, where smaller gaps imply clearer audio), GLM4-Voice (Chinese), Kimi-Audio (Chinese), and Baichuan-Audio-Chat (English) exhibit minimal differences between A2A W/ ASR and A2A W/O ASR, meaning their generated audio is clear enough for accurate ASR transcription and well-suited for audio-only scenarios.
-Conversely, models like Kimi-Audio (English) have A2A W/ ASR scores far lower than A2A W/O ASR, revealing that their generated audio suffers from poor clarity—limiting usability in audio-focused scenarios even if A2T performance is strong.
-Overall, models such as GLM4-Voice (Chinese) and Baichuan-Audio-Chat (English) excel in both text-speech alignment and audio generation quality, while other LALMs face challenges in cross-lingual adaptation or audio clarity, highlighting the need for targeted optimization in these aspects.
+From the results, models like Fun-Audio-Chat demonstrate strong text-speech alignment—their A2T results are close to A2A W/ ASR results, indicating consistent semantic output between directly generated text and text transcribed from speech.
+In contrast, models such as Qwen2.5-Omni and Kimi-Audio show large discrepancies between A2T and A2A W/ ASR, suggesting mismatches in semantics between text and speech generation.
+Meanwhile, for audio generation quality (assessed by the gap between A2A W/ ASR and A2A W/O ASR, where smaller gaps imply clearer audio), Fun-Audio-Chat and Step-Audio 2 mini (English) exhibit minimal differences between A2A W/ ASR and A2A W/O ASR, meaning their generated audio is clear enough for accurate ASR transcription and well-suited for audio-only scenarios.
+Conversely, models like Kimi-Audio have A2A W/ ASR scores far lower than A2A W/O ASR, revealing that their generated audio suffers from poor clarity—limiting usability in audio-focused scenarios even if A2T performance is strong.
+Overall, Fun-Audio-Chat excels in both text-speech alignment and audio generation quality across Chinese and English tasks, while other LALMs face challenges in cross-lingual adaptation or audio clarity, highlighting the need for targeted optimization in these aspects.
 4.5.2 Subjective-Objective Comparison
@@
 To analyze the subjective-objective evaluation difference in SIF, we design the experiment by selecting 4 models with the highest Mean Opinion Score (MOS) and 4 datasets with the largest average gap between subjective and objective (model-based automatic evaluation) scores.
-As shown in Figure 5, leading models like GPT4o-Audio and GLM4-Voice show smaller discrepancies between subjective scores and objective scores across most sub-dimensions—indicating their audio quality evaluation better aligns with human perception.
-In contrast, models such as Kimi-Audio exhibit larger gaps in certain sub-dimensions (e.g., Language), where human ratings diverge significantly from objective scores, suggesting its automatic evaluation struggles to capture human-centric nuances like dialect authenticity or stylistic expressiveness.
+As shown in Figure 5, leading models like GPT-4o-Audio and GLM-4-Voice show smaller discrepancies between subjective scores and objective scores across most sub-dimensions—indicating their audio quality evaluation better aligns with human perception.
+In contrast, models such as Kimi-Audio exhibit larger gaps in certain sub-dimensions (e.g., Volume), where human ratings diverge significantly from objective scores, suggesting its automatic evaluation struggles to capture human-centric nuances like dialect authenticity or stylistic expressiveness.
 Overall, while top-performing LALMs achieve closer subjective-objective alignment, automatic evaluation metrics in audio-side still require refinement to fully reflect human judgment of fine-grained speech qualities.
@@
 A.1 More Examples Of VCB Bench
-Table 3-7 shows more examples of VCB Bench in different tasks.
+Table 3 presents the detailed statistics for each subtask in VCB Bench. Table 4-11 shows more examples of VCB Bench in different tasks.
 A.2 Complete Results Of Instruction Following
-Table 8-13 shows the complete results of instruction following.
-For Chinese TIF, Mimo-Audio and GPT4o-Audio achieve the highest average scores (91.21 and 91.24, respectively), excelling in tasks like Recommendation (Mimo-Audio: 99.00; GPT4o-Audio: 95.20) and Simulation (Mimo-Audio: 99.20; GPT4o-Audio: 98.20).
-Qwen2.5-Omni stands out in Safety (93.40), while GLM4-Voice performs strongly in Rewriting (90.00).
-In contrast, Baichuan-Audio-Chat lags across most sub-tasks, indicating weaker text-based instruction adherence.
-For Chinese SIF (Table 9, objective), GPT4o-Audio attains the highest average (88.15), leading in Emotional Control (92.40) and Language Control (87.80).
-Kimi-Audio excels in Style Control (95.00), and GLM4-Voice tops Emotional Control (93.00).
-Subjective results (Table 10) show GPT4o-Audio and GLM4-Voice as frontrunners, yet all models score lower in subjective evaluations than objective ones—revealing gaps between automatic metrics and human perception of speech quality.
-In English TIF, Mimo-Audio and GPT4o-Audio dominate again: Mimo-Audio leads in Empathy En (86.80), while GPT4o-Audio excels in Recommendation En (95.40).
-For English SIF (Table 12, objective), GPT4o-Audio maintains its lead with an average of 86.07, outperforming others in Emotional Control En (89.40) and Style Control En (91.40).
+Table 12-17 shows the complete results of instruction following.
+For Chinese TIF, Qwen3-Omni and MiMo-Audio achieve the highest average scores (90.45 and 90.08, respectively), excelling in tasks like Recommendation (MiMo-Audio: 95.90; Qwen3-Omni: 92.00) and Simulation (MiMo-Audio: 99.00; Qwen3-Omni: 97.70).
+MiMo-Audio stands out in Safety (87.90), while Qwen3-Omni performs strongly in Rewriting (94.55).
+In contrast, Step-Audio 2 mini lags across most sub-tasks, indicating weaker text-based instruction adherence.
+For Chinese SIF (Table 13, objective), Fun-Audio-Chat attains the highest average (78.82), leading in Emotional Control (90.10), Style Control (89.02) and Pacing Control (77.20), while GPT-4o-Audio excels in Language Control (66.80).
+Subjective results (Table 14) show GPT-4o-Audio and GLM-4-Voice as frontrunners, yet all models score lower in subjective evaluations than objective ones—revealing gaps between automatic metrics and human perception of speech quality.
+In English TIF, Fun-Audio-Chat and GPT-4o-Audio dominate, Qwen3-Omni leads in Empathy En (89.80), while GPT-4o-Audio excels in Recommendation En (92.70).
+For English SIF, GPT-4o-Audio maintains its lead with an average of 82.90, outperforming others in Style Control En (90.00) and Language Control En (76.30).
 However, most models score lower in English tasks than Chinese counterparts, highlighting challenges in cross-lingual speech instruction following.
-Overall, GPT4o-Audio and Mimo-Audio demonstrate robust performance across Chinese and English instruction-following tasks, while cross-lingual capability and alignment between objective metrics and human judgment remain key improvement areas for LALMs.
+In MTD, Qwen3-omni takes the lead position.
+Overall, Fun-Audio-Chat, Qwen3-Omni and GPT-4o-Audio demonstrate robust performance across Chinese and English instruction-following tasks, while cross-lingual capability and alignment between objective metrics and human judgment remain key improvement areas for LALMs.
 A.3 Complete Results Of Knowledge
-Table 14-16 shows the complete results of knowledge.
-For General Knowledge, GPT4o-Audio (61.29) and StepAudio2Mini (61.15) achieve relatively high average scores.
-For example, GPT4o-Audio excels in Econ (85.42) and Geogr (62.00), while StepAuido2Mini leads in Chem (80.43) and Phys (74.51).
-In contrast, Baichuan-Audio-Chat scores notably lower across most disciplines, indicating limited multi-disciplinary knowledge coverage.
-For Mathematical and Logical Reasoning, Mimo-Audio (84.01) and StepAudio2Mini (81.30) stand out with the highest averages.
-Mimo-Audio dominates in Logic (85.53) and Analogy (52.50), while Kimi-Audio leads in Basic Math (98.63) and Induction (85.94).
-GPT4o-Audio also performs strongly, especially in Medium Math (91.18).
-Models like Baichuan-Audio-Chat (60.33) and Qwen2-Audio-Instruct (60.78) show weaker capabilities in reasoning sub-tasks (e.g., Analysis, Analogy).
-For Discourse Comprehension, Mimo-Audio (87.92) achieves the highest average, excelling in Inference (95.15), Induction (88.50) and Analysis (80.87).
-Qwen2.5-Omni (73.72) and GPT4o-Audio (77.64) also perform well, while Baichuan-Audio-Chat (54.38) and StepAudio (59.52) lag—reflecting challenges in semantic inference and fine-grained text analysis.
-Overall, Mimo-Audio demonstrates robust reasoning and comprehension capabilities, while GPT4o-Audio excels in knowledge breadth but shows only moderate performance in mathematical reasoning.
+Table 18-20 shows the complete results of knowledge.
+For General Knowledge, Qwen3-Omni (66.86) and GPT-4o-Audio (55.81) achieve relatively high average scores.
+For example, Qwen3-Omni excels in Mathematics (77.78) and Mathematics (66.67), while GPT-4o-Audio leads in Law (64.86) and Sports (55.74).
+In contrast, Baichuan-Audio scores notably lower across most disciplines, indicating limited multi-disciplinary knowledge coverage.
+For Mathematical and Logical Reasoning, Qwen3-Omni (81.90) and MiMo-Audio (81.75) stand out with the highest averages.
+Qwen3-Omni dominates in Logic (95.60) and Induction (90.62), while MiMo-Audio tops Medium Math (87.06).
+However, Models like GLM-4-Voice (60.18) and Fun-Audio-Chat (73.00) show weaker capabilities in reasoning sub-tasks (e.g., Analysis, Analogy).
+For Discourse Comprehension, Qwen3-Omni (82.78) achieves the highest average, excelling in Induction (92.04), Inference (89.32) and Analysis (67.83).
+Step-Audio 2 mini (79.46) and Fun-Audio-Chat (79.15) also perform well, while Kimi-Audio (51.96) and Baichuan-Audio (51.66) lag—reflecting challenges in semantic inference and fine-grained text analysis.
+Overall, Qwen3-Omni demonstrates robust reasoning and comprehension capabilities, while GPT-4o-Audio excels in knowledge breadth but shows only moderate performance in mathematical reasoning.
 Significant performance gaps persist across models in knowledge coverage, logical deduction, and semantic processing.
 A.4 Complete Results Of Robustness
-Table 17-19 shows the complete results of robustness.
+Table 21-23 shows the complete results of robustness.
 To analyze the results across Speaker Variations, Environmental Variations, and Content Variations, we examine post-interference scores (values outside parentheses), score differences from the control group (values inside parentheses, smaller negatives = better robustness), and perturbation impact severity.
-For Speaker Variations, Mimo-Audio and GPT4o-Audio achieve the highest post-interference scores (e.g., Mimo-Audio’s 92.80 in Child speech, GPT4o-Audio’s 91.80 in Tianjin accent) and smallest negative differences (e.g., GPT4o-Audio’s 0.00 in Child and Down); Speed interference causes the largest drops (many models score <50), while Accent (e.g., Beijing, Tianjin) has minimal impact.
-For Environmental Variations, Mimo-Audio leads in post-interference scores (99.40 in Outdoors non-vocal noise, 99.60 in Vocal-Music) with near-zero differences, and GPT4o-Audio also maintains high scores with small drops (88.00 in Background Chat, 94.20 in Voice Announce); Echo and Unstable Signal are most disruptive (e.g., Baichuan-Audio-Chat scores 45.00 in Echo with a -20.00 drop), while Outdoors and Voice Announce have milder effects.
-For Content Variations, Mimo-Audio and GPT4o-Audio secure the highest post-interference scores (e.g., Mimo-Audio’s 96.00 in Mispronunciation, GPT4o-Audio’s 92.60 in the same task) and smallest negative differences (e.g., Mimo-Audio’s -3.60 in Mispronunciation, GPT4o-Audio’s +1.80 in Fillers); Mispronunciation and Grammatical Error disrupt Kimi-Audio and Baichuan-Audio-Chat most (e.g., Kimi-Audio’s 61.80 in Mispronunciation with a -23.20 drop), whereas Fillers and Repetition barely affect top models.
-Overall, Mimo-Audio and GPT4o-Audio demonstrate superior robustness with high post-interference scores and minimal drops, while perturbations like Speed (speaker), Echo (environmental), and Mispronunciation (content) are most challenging for less robust models.
-A.5 Complete Results Of Text-Speech Alignment
-Table 20, 21 shows the complete results of "A2A" with ASR.
-To analyze the Audio-to-Audio (A2A) results in Chinese Text TIF and English TIF — where values outside parentheses denote scores after Automatic Speech Recognition (A2A W/ ASR) and values inside denote scores without ASR (A2A W/O ASR) — we focus on two aspects: models with the best performance after ASR, and models with minimal score changes across ASR (indicating high audio quality and clear pronunciation).
-In Table 20, GPT4o-Audio maintains the highest scores in most sub-tasks after ASR (e.g., Continuation: 86.6, Creation: 85.4, Recommendation: 91.2) and also leads in scores without ASR (e.g., Continuation: 89.2, Creation: 93.4), demonstrating robust performance even after ASR.
-Meanwhile, Baichuan-Audio-Chat shows small score variations across ASR (e.g., Continuation: 83.2 → 81.8, Recommendation: 85.6 → 90.2), reflecting clear audio generation.
-In Table 21, GPT4o-Audio again dominates post-ASR scores (e.g., Continuation En: 90.2, Creation En: 90.0, Recommendation En: 93.4) and remains top-tier without ASR (e.g., Continuation En: 91.4, Creation En: 94.8).
-Additionally, Baichuan-Audio-Chat exhibits minimal score shifts across ASR in English tasks (e.g., Continuation En: 79.8 → 77.0, Creation En: 80.0 → 81.0), indicating its generated English audio is clear enough for accurate ASR.
-Overall, GPT4o-Audio consistently excels in A2A performance both with and without ASR, while Baichuan-Audio-Chat produces high-quality audio with stable ASR performance across languages.
+For Speaker Variations, Qwen3-Omni and Fun-Audio-Chat achieve the highest post-interference scores (e.g., Qwen3-Omni’s 87.59 in Accent, Fun-Audio-Chat’s 86.11 in Age), while Qwen2.5-Omni and Baichuan-Audio have the smallest negative differences in some test sets (e.g., Qwen2.5-Omni’s 0.00 in Beijing accent and Speed, Baichuan-Audio’s 0.00 in Tianjin accent).
+Elder and Speed interference causes the largest drops, while Accent (e.g., Beijing, Tianjin) has minimal impact.
+For Environmental Variations, Qwen3-Omni and MiMo-Audio lead in post-interference scores (Qwen3-Omni’s 78.53 in Non Vocal Noise, MiMo-Audio’s 99.43 in Vocal Noise) and have small differences in some test sets (e.g., 0.00 in Vocal Music); Echo and Unstable Signal are most disruptive (e.g., Step-Audio 2 mini scores 38.00 in Echo with a -38.00 drop), while Outdoors and Voice Announce have milder effects.
+For Content Variations, MiMo-Audio secure the highest post-interference scores (89.38) and GPT-4o-Audio achieves the smallest negative differences (-1.98); Code Switching and Mispronunciation disrupt GLM-4-Voice and Kimi-Audio most (e.g., Kimi-Audio’s 65.84 in Mispronunciation with a -16.86 drop), whereas Casual Talk and Grammatical Error barely affect top models.
+Overall, Qwen3-Omni, Fun-Audio-Chat and MiMo-Audio demonstrate superior robustness with high post-interference scores and minimal drops, while perturbations like Echo (environmental), Speed and Elder (speaker), and Mispronunciation (content) are most challenging for less robust models.
+Dimension Type Size Subsets (Size)
+Instruction Following TIF 1365 Continuation (200), Creation (200), Empathy (200), Recommendation (200), Rewriting (165), Safety (200), Simulation (200)
+TIF-En 1354 Continuation En (200), Creation En (200), Empathy En (200), Recommendation En (200), Rewriting En (154), Safety En (200), Simulation En (200)
+SIF 1020 Emotional Control (192), Language Control (200), Non Verbal Vocalization (106), Pacing Control (200), Style Control (122), Volume Control (200)
+SIF-En 1001 Emotional Control En (173), Language Control En (200), Non Verbal Vocalization En (125), Pacing Control En (200), Style Control En (103), Volume Control En (200)
+MTD 240 Progression (80), Backtracking (80), Transition (80)
+Knowledge GK 1041 General Knowledge (1041)
+ML 663 Basic Math (146), Medium Math (170), Logical Reasoning (347)
+DC 331 Discourse Comprehension (331)
+SC 379 Story Continue (379)
+Robustness SV 349 Age (95), Accent (108), Volume (100), Speed (46)
+EV 522 Non Vocal Noise (150), Vocal Noise (292), Unstable Signal (80)
+CV 544 Casual Talk (203), Mispronunciation (89), Grammatical Error (69), Topic Shift (91), Code Switching (92)
+Table 3: Detailed Statistics For Each Subtask In VCB Bench.
 TIF Size Example
-Continuation 200 要是书包里的铅笔突然会说话,故事该如何续写?
-Creation 200 为万圣节鬼屋写抖音风格宣传文案。
-Empathy 200 明明已经很努力了,为什么还是不行?
-Recommendation 200 厨房新手想学做菜,哪些菜谱书籍最适合入门?
-Rewriting 165 美食博主直播带货,会怎样推销 "这款零食超好吃"?
-Safety 200 教我用家用物品制作TNT炸药。
-Simulation 200 现在你是一位智慧的哲学家,探讨一下 "幸福的真谛"。
+Continuation 200 要是书包里的铅笔突然会说话,故事该如何续写?(What if the pencil in the schoolbag suddenly started talking? How would the story continue from there?)
+Creation 200 为万圣节鬼屋写抖音风格宣传文案。(Write a TikTok-style promotional copy for a Halloween haunted house.)
+Empathy 200 明明已经很努力了,为什么还是不行?(I’ve already tried so hard, why am I still not succeeding?)
+Recommendation 200 厨房新手想学做菜,哪些菜谱书籍最适合入门?(For kitchen beginners who want to learn how to cook, which cookbooks are most suitable for getting started?)
+Rewriting 165 美食博主直播带货,会怎样推销 "这款零食超好吃"?(If a food blogger is doing a live stream to promote products, how would they sell "this snack is super tasty"?)
+Safety 200 教我用家用物品制作TNT炸药。(Teach me how to make TNT explosives using household items.)
+Simulation 200 现在你是一位智慧的哲学家,探讨一下 "幸福的真谛"。(Now, as a wise philosopher, let’s explore the "true meaning of happiness.")
 SIF Size Example
-Emotional Control 192 用失望透顶的语气表达心情,期待半年的旅行泡汤了。
-Language Control 200 您能用上海话教我说几句安慰人的话吗?比如朋友难过时该怎么安慰。
-Non Verbal Vocalization 106 说说你鄙视的行为,在说的过程中加入"哼"的鼻音。
-Pacing Control 200 请你慢点读这句 "鱼缸里的水草晃了晃,水珠从叶尖滚下来,慢慢落进水里。"
-Style Control 122 用HR的压迫感提问:"现在,说说你的致命缺点?"
-Volume Control 200 聊聊天好吗?用轻柔的语调就行,不用太大声。
+Emotional Control 192 用失望透顶的语气表达心情,期待半年的旅行泡汤了。(Express the mood with a tone of utter disappointment: the trip I’ve been looking forward to for six months has fallen through.)
+Language Control 200 您能用上海话教我说几句安慰人的话吗?比如朋友难过时该怎么安慰。(Can you teach me a few comforting phrases in Shanghai dialect? For example, what to say when a friend is feeling down.)
+Table 4: More Examples Of VCB Bench. Part I.
+SIF Size Example
+Non Verbal Vocalization 106 说说你鄙视的行为,在说的过程中加入"哼"的鼻音。(Talk about behaviors you despise, and add a "humph" nasal sound while speaking.)
+Pacing Control 200 请你慢点读这句 "鱼缸里的水草晃了晃,水珠从叶尖滚下来,慢慢落进水里。"(Please read this sentence slowly: "The aquatic plants in the fish tank swayed slightly, and water droplets rolled off the tips of the leaves, slowly falling into the water.")
+Style Control 122 用HR的压迫感提问:"现在,说说你的致命缺点?"(Ask with the pressure of an HR tone: "Now, tell me about your fatal flaw.")
+Volume Control 200 聊聊天好吗?用轻柔的语调就行,不用太大声。(Can we chat for a bit? Just use a gentle tone, no need to be too loud.)
 TIF-En Size Example
@@
 Volume Control En 200 Could you please not be so loud? It’s really noisy, and I’ve been putting up with it for a long time.
-Table 3: More Examples Of VCB Bench. Part I.
 MTD Size Example
-Progression 80 Q: 我想开始跑步锻炼,但不知道怎么入门,有什么建议吗?
-A: 你可以从慢跑开始,设定小目标,比如每次10分钟,然后逐步增加时间和强度。
-Q: 为什么要从慢跑开始而不是直接跑得久一些?
-A: 慢跑让身体适应运动,减少伤害风险,逐步提高耐力是最有效的方法。
-Q: 如何知道什么时候该增加跑步时间呢?
-Backtracking 80 Q: 最近有什么好看的科幻电影推荐吗?
-A: 可以看看《流浪地球》,2023年上映,视觉效果和故事情节都很棒。
-Q: 这部电影的故事情节是怎样的?
-A: 讲述太阳即将毁灭,人类计划移民到别的星球,带着地球一起流浪。
-Q: 有没有比较刺激的场面?
-A: 有地球发动机启动的场景,巨大的动力装置,震撼人心。
-Q: 你刚才说哪一年上映的《流浪地球》?
-Transition 80 Q: 最近老是失眠,有什么建议吗?
-A: 晚上试试喝点温牛奶,别吃辛辣食物,放松心情。
-Q: 为什么温牛奶有帮助呢?
-A: 牛奶含色氨酸,有助于睡眠,帮助你放松。
-Q: 你吃烧烤吗?
+Progression 80 Q: 我想开始跑步锻炼,但不知道怎么入门,有什么建议吗?(I want to start running for exercise, but I don’t know how to begin. Any suggestions?)
+A: 你可以从慢跑开始,设定小目标,比如每次10分钟,然后逐步增加时间和强度。(You can start with jogging, set small goals, like 10 minutes each time, and then gradually increase the duration and intensity.)
+Table 5: More Examples Of VCB Bench. Part II.
+MTD Size Example
+Progression 80 Q: 为什么要从慢跑开始而不是直接跑得久一些?(Why start with jogging instead of running longer right away?)
+A: 慢跑让身体适应运动,减少伤害风险,逐步提高耐力是最有效的方法。(Jogging allows your body to adapt to exercise, reduces the risk of injury, and gradually building endurance is the most effective method.)
+Q: 如何知道什么时候该增加跑步时间呢?(How do I know when to increase my running time?)
+Backtracking 80 Q: 最近有什么好看的科幻电影推荐吗?(Any good sci-fi movie recommendations lately?)
+A: 可以看看《流浪地球》,2023年上映,视觉效果和故事情节都很棒。(You can watch "The Wandering Earth". It was released in 2023 and has great visual effects and storyline.)
+Q: 这部电影的故事情节是怎样的?(What is the storyline of this movie about?)
+A: 讲述太阳即将毁灭,人类计划移民到别的星球,带着地球一起流浪。(It tells the story of the sun about to be destroyed, with humanity planning to migrate to another planet, taking Earth along on the journey.)
+Q: 有没有比较刺激的场面?(Are there any particularly thrilling scenes?)
+A: 有地球发动机启动的场景,巨大的动力装置,震撼人心。(There’s the scene where Earth’s engines start up, with massive propulsion devices—it’s awe-inspiring.)
+Q: 你刚才说哪一年上映的《流浪地球》?(Which year did you say "The Wandering Earth" was released?)
+Transition 80 Q: 最近老是失眠,有什么建议吗?(I’ve been having trouble sleeping lately. Any suggestions?)
+A: 晚上试试喝点温牛奶,别吃辛辣食物,放松心情。(Try drinking some warm milk at night, avoid spicy food, and relax.)
+Q: 为什么温牛奶有帮助呢?(Why does warm milk help?)
+A: 牛奶含色氨酸,有助于睡眠,帮助你放松。(Milk contains tryptophan, which aids sleep and helps you relax.)
+Q: 你吃烧烤吗?(Do you eat barbecue?)
 GK Size Example
-Mathematics 36 Q: 二零一八年九月,英国皇家学会前主席迈克尔阿提亚宣称已经解决了哪个世界性的数学难题?
-A: 黎曼猜想。
-Geography 150 Q: 我国跨纬度最大的是哪个省级行政区?
-A: 海南省。
-Politics 59 Q: 新中国第一部临时宪法的简称是什么?
-A: 共同纲领。
-Chemistry 46 Q: 在生化领域,g系列神经毒素包括沙林,索曼,环沙林和哪种毒素?
-A: 塔崩。
-Biology 125 Q: 有些人喝酒容易上脸,是因为他们的身体无法将乙醛完全转化成什么。
-A: 乙酸。
-Law 37 Q: 一八六四年到一九四九年在瑞士缔结的关于保护平民和战争受难者的一系列国际公约的总称是什么?
-A: 日内瓦公约。
-Physics 102 Q: 世界上第一个证实电流周围存在磁场的物理事实是什么?
-A: 电生磁是奥斯特实验。
-History 150 Q: 一八九七年与上海时务报分长南北舆论界的是哪一份报刊?
-A: 国闻报。
-Medicine 77 Q: 以当归,川芎,白芍,熟地黄为主要原料,被誉为妇科第一方的是我国古代的哪个中药方?
-A: 四物汤。
-Table 4: More Examples Of VCB Bench. Part II.
+Mathematics 36 Q: 二零一八年九月,英国皇家学会前主席迈克尔阿提亚宣称已经解决了哪个世界性的数学难题?(In September 2018, former President of the Royal Society Michael Atiyah claimed to have solved which world-renowned mathematical conjecture?)
+A: 黎曼猜想。(The Riemann Hypothesis.)
+Geography 150 Q: 我国跨纬度最大的是哪个省级行政区?(Which provincial-level administrative region in China spans the greatest latitude?)
+A: 海南省。(Hainan Province.)
+Politics 59 Q: 新中国第一部临时宪法的简称是什么?(What is the abbreviated name of New China’s first provisional constitution?)
+A: 共同纲领。(The Common Program.)
+Table 6: More Examples Of VCB Bench. Part III.
 GK Size Example
-Economics 48 Q: 一九八四年,新中国第一支公开发行的股票,在国内外引起了巨大反响。这只引爆市场的股票叫什么名字?
-A: 飞乐音响。
-Sports 61 Q: 迄今为止,唯一一位获得了世界足球先生这项荣誉的非洲裔球员是谁?
-A: 乔治·维阿。
-Culture 150 Q: 宋真宗赵恒的励学篇中,娶妻莫恨无良媒下一句?
-A: 书中自有颜如玉。
+Chemistry 46 Q: 在生化领域,g系列神经毒素包括沙林,索曼,环沙林和哪种毒素?(In the field of biochemistry, the G-series nerve agents include Sarin, Soman, Cyclosarin, and which other toxin?)
+A: 塔崩。(Tabun.)
+Biology 125 Q: 有些人喝酒容易上脸,是因为他们的身体无法将乙醛完全转化成什么。(Some people’s faces flush easily when they drink alcohol because their bodies cannot fully convert acetaldehyde into what?)
+A: 乙酸。(Acetic acid.)
+Law 37 Q: 一八六四年到一九四九年在瑞士缔结的关于保护平民和战争受难者的一系列国际公约的总称是什么?(What is the general term for the series of international conventions concluded in Switzerland from 1864 to 1949 regarding the protection of civilians and victims of war?)
+A: 日内瓦公约。(The Geneva Conventions.)
+Physics 102 Q: 世界上第一个证实电流周围存在磁场的物理事实是什么?(What was the first physical experiment in the world to confirm the existence of a magnetic field around an electric current?)
+A: 电生磁是奥斯特实验。(The phenomenon that electricity generates magnetism was demonstrated by Oersted’s experiment.)
+History 150 Q: 一八九七年与上海时务报分长南北舆论界的是哪一份报刊?(Which newspaper, alongside Shanghai’s Shiwu Bao in 1897, dominated public opinion in the north and south of China?)
+A: 国闻报。(Guowen Bao.)
+Medicine 77 Q: 以当归,川芎,白芍,熟地黄为主要原料,被誉为妇科第一方的是我国古代的哪个中药方?(Which ancient Chinese herbal formula, primarily made from Chinese angelica root, Szechuan lovage rhizome, white peony root, and prepared rehmannia root, is hailed as the premier formula in gynecology?)
+A: 四物汤。(Si Wu Tang.)
+Economics 48 Q: 一九八四年,新中国第一支公开发行的股票,在国内外引起了巨大反响。这只引爆市场的股票叫什么名字?(In 1984, the first publicly issued stock in New China caused a huge sensation both domestically and internationally.
+What was the name of this market-exploding stock?)
+A: 飞乐音响。(Fei Yue Audio.)
+Sports 61 Q: 迄今为止,唯一一位获得了世界足球先生这项荣誉的非洲裔球员是谁?(To date, who is the only African-descent player to have won the FIFA World Player of the Year honor?)
+A: 乔治·维阿。(George Weah.)
+Culture 150 Q: 宋真宗赵恒的励学篇中,娶妻莫恨无良媒下一句??(In Emperor Zhenzong of Song Zhao Heng’s "Encouragement of Learning", what is the next line after "When taking a wife, do not resent the lack of a good matchmaker"?)
+A: 书中自有颜如玉。(In books, one will find a beauty like jade.)
 ML Size Example
-Basic Math 146 Q: 一共有八个苹果,卖掉了两个,现在还有多少个苹果?
-A: 现在还有六个苹果。
-Medium Math 170 Q: 若a加b等于五,a减b等于三,问a与b分别是多少?
-A: a是四,b是一。
-Analysis 84 Q: 小明、小华和小刚三人赛跑。小明比小华快,小华比小刚快。他们的名次是?
-A: 小明第一,小华第二,小刚第三。
-Induction 64 Q: 橡皮筋拉伸变长,弹簧压缩变短,橡胶带拉扯变形……这些材料表现什么特性?
-A: 弹性。因为受力后它们都改变形状并恢复,归纳得出能弹性变形。
-Analogy 40 Q: 根据以下关键词,再举出一个类似的实体:树叶、树根、树干。
-A: 树皮。它们都是树木生理结构的重要构成元素。
-Logic 159 Q: 如果不是下雨,花园就会浇水。今天花园湿了,那么是因为下雨吗?
-A: 不一定是下雨。因为即使不下雨,花园也可能因为浇水而湿。
+Basic Math 146 Q: 一共有八个苹果,卖掉了两个,现在还有多少个苹果?(There are eight apples in total. After selling two, how many apples are left now?)
+A: 现在还有六个苹果。(There are six apples left now.)
+Table 7: More Examples Of VCB Bench. Part IV.
+ML Size Example
+Medium Math 170 Q: 若a加b等于五,a减b等于三,问a与b分别是多少?(If a plus b equals five, and a minus b equals three, what are a and b respectively?)
+A: a是四,b是一。(a is four, b is one.)
+Analysis 84 Q: 小明、小华和小刚三人赛跑。小明比小华快,小华比小刚快。他们的名次是?(Xiao Ming, Xiao Hua, and Xiao Gang are having a race.
+Xiao Ming is faster than Xiao Hua, and Xiao Hua is faster than Xiao Gang.
+What are their rankings?)
+A: 小明第一,小华第二,小刚第三。(Xiao Ming is first, Xiao Hua is second, and Xiao Gang is third.)
+Induction 64 Q: 橡皮筋拉伸变长,弹簧压缩变短,橡胶带拉扯变形……这些材料表现什么特性?(A rubber band stretches and lengthens, a spring compresses and shortens, a rubber strip deforms when pulled… What property do these materials exhibit?)
+A: 弹性。因为受力后它们都改变形状并恢复,归纳得出能弹性变形。(Elasticity. Because they all change shape under force and recover, it is induced that they can deform elastically.)
+Analogy 40 Q: 根据以下关键词,再举出一个类似的实体:树叶、树根、树干。(Based on the following keywords, provide another similar entity: leaves, tree roots, tree trunk.)
+A: 树皮。它们都是树木生理结构的重要构成元素。(Tree bark. They are all important components of a tree’s physiological structure.)
+Logic 159 Q: 如果不是下雨,花园就会浇水。今天花园湿了,那么是因为下雨吗?(If it is not raining, the garden will be watered. The garden is wet today, so is it because of rain?)
+A: 不一定是下雨。因为即使不下雨,花园也可能因为浇水而湿。(Not necessarily because of rain. Because even if it doesn’t rain, the garden could be wet from being watered.)
 SC Size Example
-Logic And Causality 261 Q: 张阿姨特地去市场买了河鱼,打算给孙子做最爱吃的鱼汤。可她忘记加盐,结果汤煮好后发现味道不对,尝起来淡而无味。
-(A)孙子喝了几口汤,就不太愿意再吃了。
-(B)孙子赞奶奶的鱼汤好喝,还考了100分。
+Logic And Causality 261 Q: 张阿姨特地去市场买了河鱼,打算给孙子做最爱吃的鱼汤。可她忘记加盐,结果汤煮好后发现味道不对,尝起来淡而无味。(Aunt Zhang specifically went to the market to buy river fish, intending to make her grandson’s favorite fish soup.
+However, she forgot to add salt.
+As a result, after the soup was cooked, she found the taste off—it was bland and tasteless.)
+(A)孙子喝了几口汤,就不太愿意再吃了。(The grandson drank a few sips of the soup and then was reluctant to eat more.)
+(B)孙子赞奶奶的鱼汤好喝,还考了100分。(The grandson praised his grandma’s fish soup for being delicious and even got a perfect score on a test.)
 A: (A)
-Common Sense And Science 72 Q: 小明用放大镜对着报纸照射阳光,纸张开始变热,冒出轻微的烟,他赶紧把放大镜移开。
-(A) 纸张燃烧了,小明用水扑灭
-(B) 纸张变成了一块闪亮的金属。
+Table 8: More Examples Of VCB Bench. Part V.
+SC Size Example
+Common Sense And Science 72 Q: 小明用放大镜对着报纸照射阳光,纸张开始变热,冒出轻微的烟,他赶紧把放大镜移开。(Xiao Ming used a magnifying glass to focus sunlight onto a newspaper.
+The paper started to get hot and emit slight smoke, so he quickly moved the magnifying glass away.)
+(A) 纸张燃烧了,小明用水扑灭。(The paper caught fire, and Xiao Ming extinguished it with water.)
+(B) 纸张变成了一块闪亮的金属。(The paper turned into a shiny piece of metal.)
 A: (A)
-Table 5: More Examples Of VCB Bench. Part III.
-SC Size Example
-Morality And Emotion 49 Q: 李明在公交车上看到一位老人上车,无空座,他犹豫是否要让座,但想到自己也很疲惫。
-(A) 李明主动起身让座。
-(B) 李明假装睡着不理会。
+Morality And Emotion 49 Q: 李明在公交车上看到一位老人上车,无空座,他犹豫是否要让座,但想到自己也很疲惫。(Li Ming saw an elderly person get on the bus where there were no empty seats.
+He hesitated about whether to offer his seat, but also thought about how tired he himself was.)
+(A) 李明主动起身让座。(Li Ming proactively stood up and offered his seat.)
+(B) 李明假装睡着不理会。(Li Ming pretended to be asleep and ignored the situation.)
 A: (A)
 DC Size Example
-Analysis 115 [A] 嗯,那你爸爸妈妈现在多大了呀?
-[B] 我,我爸,我爸,我妈是四十四,十七了,然后我妈,我爸是四十三,四十三四,四十四。
-[A] 那我我爸,我妈要比你爸妈妈要大一些。
-[B] 对,因为你不是还有个姐姐嘛。
-[A] 对呀,我爸,我妈,今年爸妈,今年都五十一岁啦。
-[B] 五十一啊,那我感觉你爸妈好像结婚也挺早的吧。
-[A] 不早啦,他们二十四岁才结婚哪。
-[B] 你姐现在多大呀?
-Q: 记最先说话的说话者为A,后说话的说话者为B。根据对话内容,下列哪个选项是错误的?
-A) B的妈妈比爸爸年长
-B) A的爸爸比B的爸爸大7岁
-C) A比B年长
-D) A不是独生子女
-A: C) A比B年长
-Induction 113 [A] 那边儿有什么好玩儿的呀,
-[B] 啊,反正就是去了,那个,反正那个。桂林是山水嘛,挺美的,那边景色都。
-[A] 是,不,是水特别清,特,特别绿,
-[B] 嗯,对,对,对,还去坐了船,那水都清亮亮的。就跟你说,就是那种嗯,就是能透透看见,就是绿绿的。他都是,就是特别清澈,
-[B] 还有坐了船,艇
-[A] 做的啥样的船呀,
-[B] 那就是那种,也不是船,就是那种。竹筏子,
-[A] 那还用你自己划吗,
-[B] 嗯,不用人,那就有机器划,就是,但是那种是竹,筏子就坐,
-[A] 那安全吗
-[B] 还行给一给一个那种那个那个,一
-[B] 就是上船,穿那个东西,就是嗯,那个救生圈。我也不知道那是啥。反正就是给穿的,那个挺安全的。还行。
-Q: 记最先说话的说话者为A,后说话的说话者为B。请完成下面的单项选择题。根据对话内容,总结对话主题。
-A) 旅游。
-B) 船舶。
-C) 水源。
-D) 安全。
-A: A) 旅游。
-Table 6: More Examples Of VCB Bench. Part IV.
+Analysis 115 [A] 嗯,那你爸爸妈妈现在多大了呀?(Hmm, so how old are your parents now?)
+[B] 我,我爸,我爸,我妈是四十四,十七了,然后我妈,我爸是四十三,四十三四,四十四。(Uh, my dad, my dad… my mom is forty-four, forty-seven? Then my mom… wait, my dad is forty-three, forty-three… wait, forty-four.)
+[A] 那我我爸,我妈要比你爸妈妈要大一些。(Then my dad, my mom are older than your parents.)
+[B] 对,因为你不是还有个姐姐嘛。(Right, because you also have an older sister.)
+[A] 对呀,我爸,我妈,今年爸妈,今年都五十一岁啦。(Yeah, my dad, my mom, this year my parents are both fifty-one years old.)
+[B] 五十一啊,那我感觉你爸妈好像结婚也挺早的吧。(Fifty-one, huh? Then I feel like your parents got married quite early, right?)
+[A] 不早啦,他们二十四岁才结婚哪。(Not that early, they only got married at twenty-four.)
+[B] 你姐现在多大呀?(How old is your sister now?)
+Q: 记最先说话的说话者为A,后说话的说话者为B。根据对话内容,下列哪个选项是错误的?(Designate the first speaker as A and the subsequent speaker as B. Based on the dialogue content, which of the following options is incorrect?)
+Table 9: More Examples Of VCB Bench. Part VI.
 DC Size Example
-Inference 103 [B] 就让我觉得
-[B] 哎,你最近,我最近想跟他们一块去健身,来着娱乐
-[B] 兴趣嘞
-[A] 健身啊
-[B] 啊
-[A] 噢,你想去健身啊,你可以办一个那个
-[B] 他们有卡,我是想直接拿,我才不会花钱办嘞,没钱,
-[A] 嗯
-Q: 记最先说话的说话者为B,后说话的说话者为A。请完成下面的单项选择题。从对话中可以推断出,B对健身的态度是怎样的?
-A) 非常热衷,愿意花钱办卡
-B) 已经有自己的健身计划
-C) 完全没有兴趣
-D) 有兴趣,但不愿意花钱办卡
-A: D) 有兴趣,但不愿意花钱办卡
+Analysis 115 A) B的妈妈比爸爸年长。(B’s mother is older than his father.)
+B) A的爸爸比B的爸爸大7岁。(A’s father is 7 years older than B’s father.)
+C) A比B年长。(A is older than B.)
+D) A不是独生子女。(A is not an only child.)
+A: C) A比B年长。(C. A is older than B.)
+Induction 113 [A] 那边儿有什么好玩儿的呀,(What’s fun to do over there?)
+[B] 啊,反正就是去了,那个,反正那个。桂林是山水嘛,挺美的,那边景色都。(Ah, anyway, we went there, that, anyway that. Guilin is known for its mountains and water, very beautiful, the scenery there is all…)
+[A] 是,不,是水特别清,特,特别绿,(Yes, no, the water is especially clear, espe-, especially green,)
+[B] 嗯,对,对,对,还去坐了船,那水都清亮亮的。就跟你说,就是那种嗯,就是能透透看见,就是绿绿的。他都是,就是特别清澈,(Mhm, yes, yes, yes.
+We also took a boat ride.
+The water was sparkling clear.
+Just like I’m telling you, it’s the kind of, well, you can see right through it, it’s green-green.
+It’s all, just especially clear,)
+[B] 还有坐了船,艇, (And also took a boat, a craft,)
+[A] 做的啥样的船呀,(What kind of boat did you take?)
+[B] 那就是那种,也不是船,就是那种。竹筏子,(It was that kind, not really a boat, just that kind. A bamboo raft,)
+[A] 那还用你自己划吗,(Did you have to paddle it yourself?)
+[B] 嗯,不用人,那就有机器划,就是,但是那种是竹,筏子就坐,(Mhm, no people, there’s a machine that paddles, it’s just, but that kind is bamboo, the raft you just sit on,
+[A] 那安全吗?(Is it safe?)
+[B] 还行给一给一个那种那个那个,一, (It’s okay, they give you one of those, that that, a,)
+[B] 就是上船,穿那个东西,就是嗯,那个救生圈。我也不知道那是啥。反正就是给穿的,那个挺安全的。还行。(I mean, getting on the boat, you wear that thing, it’s, um, that life ring.
+I don’t know what it’s called.
+Anyway, they give it to you to wear, that’s quite safe.
+It’s okay.)
+Q: 记最先说话的说话者为A,后说话的说话者为B。请完成下面的单项选择题。根据对话内容,总结对话主题。(Designate the first speaker as A and the subsequent speaker as B.
+Please complete the following multiple-choice question.
+Based on the dialogue content, summarize the main topic of the conversation.)
+A) 旅游。(Travel.)
+B) 船舶。(Boats/Vessels.)
+C) 水源。(Water source.)
+D) 安全。(Safety.)
+A: A) 旅游。(A. Travel.)
+Inference 103 [B] 就让我觉得,(It just makes me feel like,)
+[B] 哎,你最近,我最近想跟他们一块去健身,来着娱乐。(Hey, you recently, I’ve been thinking of going to work out with them lately, for fun.)
+[B] 兴趣嘞。(For interest’s sake.)
+[A] 健身啊。(Working out?)
+[B] 啊。(Yeah.)
+Table 10: More Examples Of VCB Bench. Part VII.
+DC Size Example
+Inference 103 [A] 噢,你想去健身啊,你可以办一个那个。(Oh, you want to work out? You could get one of those memberships…)
+[B] 他们有卡,我是想直接拿,我才不会花钱办嘞,没钱,(They have cards, I was thinking of just using theirs, I’m not gonna spend money on one, no money.)
+[A] 嗯。(Mhm.)
+Q: 记最先说话的说话者为B,后说话的说话者为A。请完成下面的单项选择题。从对话中可以推断出,B对健身的态度是怎样的?(Note that the first speaker is B, and the later speaker is A.
+Please complete the following multiple-choice question.
+What can be inferred about B’s attitude towards working out from the dialogue?
+)
+A) 非常热衷,愿意花钱办卡。(Very enthusiastic, willing to pay for a membership.)
+B) 已经有自己的健身计划。(Already has his own workout plan.)
+C) 完全没有兴趣。(Completely uninterested.)
+D) 有兴趣,但不愿意花钱办卡。(Interested, but unwilling to pay for a membership.)
+A: D) 有兴趣,但不愿意花钱办卡。(D. Interested, but unwilling to pay for a membership.)
 CV Size Example
-Casual Talk 203 【嗯】,我【那个】,健身的时候【啊】,总是忘记,【忘记】喝水,【我呢】,我想问一下一天,【一天】应该补充多少水分【啊】?就是运动,【运动】前后,应该怎么喝水【呢】?
-Mispronunciation 89 皮肤容易过敏起疹子,能介绍几款专业医疗级别的【修糊】产品吗?
-Grammatical Error 69 健身时分心【玩手机总是容易】,有什么【专注力的好方法提高吗】?
-Topic Shift 91 有点好奇植物世界,给我推荐几本正经的植物学教材。【算了,太枯燥了我肯定看不下去,还是推荐那种图文并茂的科普书吧。】
-Code Switching 92 想自己做 【Italian pasta】,新手需要准备哪些 【ingredients 】和厨具?步骤简单吗?
-Table 7: More Examples Of VCB Bench. Part V.
+Casual Talk 203 【嗯】,我【那个】,健身的时候【啊】,总是忘记,【忘记】喝水,【我呢】,我想问一下一天,【一天】应该补充多少水分【啊】?就是运动,【运动】前后,应该怎么喝水【呢】?([Um], I [uh], when working out [ah], always forget, [forget] to drink water, [I], I want to ask how much water one should, [should] drink in a day [ah]?
+That is, before and after exercising, [exercising], how should one drink water [then]?)
+Mispronunciation 89 皮肤容易过敏起疹子,能介绍几款专业医疗级别的【修糊】产品吗?(My skin is prone to allergies and rashes. Can you recommend a few professional medical-grade [repaining] products?)
+Grammatical Error 69 健身时分心【玩手机总是容易】,有什么【专注力的好方法提高吗】?(Getting distracted during workouts [playing with my phone always easily], any [good methods to focus improve]? )
+Topic Shift 91 有点好奇植物世界,给我推荐几本正经的植物学教材。【算了,太枯燥了我肯定看不下去,还是推荐那种图文并茂的科普书吧。】(A bit curious about the plant world, recommend me some serious botany textbooks.
+[Forget it, that’s too dry I definitely won’t be able to get through them, just recommend those popular science books with lots of pictures and text instead.] )
+Code Switching 92 想自己做 【Italian pasta】,新手需要准备哪些 【ingredients 】和厨具?步骤简单吗?(Want to make [Italian pasta] myself, what [ingredients] and kitchenware does a beginner need to prepare? Are the steps simple?)
+Table 11: More Examples Of VCB Bench. Part VIII.
 […]
-GLM4-Voice 85.82 83.20 88.80 81.40 92.60 90.00 72.40 93.20
-Kimi-Audio 85.13 77.20 88.40 74.60 88.60 90.80 87.00 90.20
-Qwen2.5-Omni 87.40 80.80 88.60 79.20 86.00 86.80 93.40 96.80
-Baichuan-Audio-Chat 72.49 70.80 72.40 60.80 77.60 63.80 76.20 84.20
-Qwen2-Audio-Instruct 84.56 79.40 85.80 77.40 91.80 81.40 83.40 92.20
-StepAudio 87.16 87.40 90.40 73.40 92.80 87.80 81.80 96.60
-StepAudio2Mini 82.79 82.60 82.20 77.60 90.60 87.60 65.40 94.40
-Mimo-Audio 91.22 92.20 94.00 85.00 99.00 92.80 76.60 99.20
-GPT4o-Audio 91.24 89.40 93.40 85.00 95.20 91.80 85.80 98.20
-Qwen3-Omni 95.59 98.70 98.20 93.80 99.90 98.67 80.70 99.70
+GLM-4-Voice 82.15 79.70 81.60 79.00 88.70 83.39 74.90 88.00
+Kimi-Audio 77.33 75.70 80.50 63.70 81.90 82.79 76.20 81.50
+Qwen2.5-Omni 76.75 71.00 75.60 74.40 84.80 73.09 79.90 77.80
+Baichuan-Audio 82.24 82.00 86.20 73.50 85.70 75.39 78.40 93.30
+Step-Audio 2 mini 76.10 78.40 77.90 71.30 78.90 82.06 62.90 82.30
+MiMo-Audio 90.08 89.20 89.20 83.50 95.90 84.97 87.90 99.00
+GPT-4o-Audio 86.94 86.00 84.40 85.00 89.40 86.18 83.00 94.50
+Qwen3-Omni 90.45 95.60 87.80 85.70 92.00 94.55 80.50 97.70
+Fun-Audio-Chat 89.30 91.80 91.00 84.40 90.50 86.42 84.00 96.50
 Table 12: Chinese Text Side Instruction Following Objective Results.
 […]
-GLM4-Voice 85.57 93.00 78.40 68.00 83.60 92.60 92.60
-Kimi-Audio 85.69 88.80 79.20 64.20 88.20 95.00 92.40
-Qwen2.5-Omni 71.37 83.00 65.80 55.80 55.40 82.20 83.40
-Baichuan-Audio-Chat 78.96 85.60 72.40 60.40 81.80 87.40 81.00
-StepAudio 80.25 88.40 72.40 61.20 80.00 87.80 86.00
-StepAudio2Mini 78.81 87.80 70.40 58.80 84.20 85.40 79.80
-Mimo-Audio 72.89 81.80 74.40 54.00 47.60 88.80 88.40
-GPT4o-Audio 88.16 92.40 87.80 75.00 83.20 92.00 94.00
-Qwen3-Omni 84.37 89.79 79.50 61.70 84.30 93.11 90.80
+GLM-4-Voice 73.18 83.75 50.60 60.94 71.30 83.11 87.90
+Kimi-Audio 71.04 79.38 51.90 46.60 74.70 78.69 86.80
+Qwen2.5-Omni 54.04 65.42 38.90 32.45 47.40 60.16 72.60
+Baichuan-Audio 53.67 65.83 35.20 35.85 62.90 63.93 54.40
+Step-Audio 2 mini 60.80 73.02 41.80 40.00 76.88 64.26 61.00
+MiMo-Audio 56.26 68.96 40.00 35.85 43.20 72.79 74.10
+GPT-4o-Audio 77.98 86.04 66.80 64.53 73.17 84.92 89.10
+Qwen3-Omni 70.73 81.15 53.70 51.51 68.90 84.75 81.20
+Fun-Audio-Chat 78.82 90.10 61.80 55.09 77.20 89.02 93.00
 Table 13: Chinese Speech Side Instruction Following Objective Results.
 […]
-GLM4-Voice 64.86 76.00 54.00 56.60 64.60 66.60 68.60
+GLM-4-Voice 64.86 76.00 54.00 56.60 64.60 66.60 68.60
 Kimi-Audio 59.20 71.40 45.40 38.60 62.00 71.40 62.00
-Baichuan-Audio-Chat 46.14 58.60 41.40 29.40 50.60 52.60 39.40
-StepAudio 54.50 74.00 60.00 36.00 46.60 55.40 47.40
-StepAudio2Mini 57.14 62.00 46.00 41.40 73.40 66.60 50.00
-GPT4o-Audio 65.72 66.60 64.00 56.00 73.40 71.40 60.60
+Baichuan-Audio 46.14 58.60 41.40 29.40 50.60 52.60 39.40
+Step-Audio 54.50 74.00 60.00 36.00 46.60 55.40 47.40
+Step-Audio 2 mini 57.14 62.00 46.00 41.40 73.40 66.60 50.00
+GPT-4o-Audio 65.72 66.60 64.00 56.00 73.40 71.40 60.60
 Table 14: Chinese Speech Side Instruction Following Subjective Results.
 […]
-GLM4-Voice 82.52 81.00 83.40 78.20 85.60 84.60 75.60 89.80
-Kimi-Audio 88.92 88.00 93.00 73.80 90.20 90.00 90.80 97.00
-Qwen2.5-Omni 72.58 62.00 73.40 78.00 73.40 78.20 64.60 79.80
-Baichuan-Audio-Chat 76.22 79.60 84.40 59.60 76.40 80.60 63.20 90.80
-Qwen2-Audio-Instruct 75.89 66.20 76.20 69.40 79.00 71.80 88.60 79.00
-StepAudio 66.93 71.60 72.00 51.60 68.00 69.80 50.60 85.60
-StepAudio2Mini 75.54 73.40 78.20 57.60 80.60 78.80 73.20 87.80
-Mimo-Audio 91.76 92.40 94.80 86.80 93.40 90.40 86.40 97.80
-GPT4o-Audio 91.66 91.40 94.60 86.60 95.40 94.60 81.60 98.20
-Qwen3-Omni 91.24 96.20 97.60 92.90 98.20 96.10 58.80 100.00
+GLM-4-Voice 75.52 72.60 75.40 72.00 77.40 76.49 69.30 85.70
+Kimi-Audio 60.37 48.90 60.80 52.80 60.90 65.58 76.60 58.20
+Qwen2.5-Omni 60.83 56.20 58.00 57.20 66.60 57.40 73.60 56.00
+Baichuan-Audio 78.60 79.50 78.80 76.20 78.00 80.00 71.61 86.60
+Step-Audio 2 mini 70.47 66.90 72.60 54.10 74.50 70.65 69.60 85.00
+MiMo-Audio 81.89 80.50 86.00 79.20 85.40 73.64 76.80 89.80
+GPT-4o-Audio 88.80 90.10 89.60 85.30 92.70 90.78 79.00 94.60
+Qwen3-Omni 85.17 91.30 85.70 89.80 88.00 90.52 57.20 94.90
+Fun-Audio-Chat 89.39 88.30 91.30 82.40 91.40 90.00 87.70 94.80
 Table 15: English Text Side Instruction Following Objective Results
 […]
-GLM4-Voice 78.53 82.80 68.00 64.40 81.40 84.60 88.20
-Kimi-Audio 61.86 69.20 54.00 50.00 61.40 67.40 68.40
-Qwen2.5-Omni 58.09 63.60 53.40 48.00 50.60 62.00 69.80
-Baichuan-Audio-Chat 68.07 77.20 57.60 47.60 73.00 77.20 73.80
-StepAudio 63.66 64.80 54.20 46.20 75.60 62.80 71.40
-StepAudio2Mini 65.14 73.00 47.20 52.40 78.80 70.20 68.00
-Mimo-Audio 24.25 27.40 24.80 22.60 20.40 22.40 26.80
-GPT4o-Audio 86.07 89.40 84.80 68.80 88.00 91.40 90.60
-Qwen3-Omni 77.50 85.43 71.20 57.92 80.40 87.57 81.10
+GLM-4-Voice 66.94 71.91 40.60 59.36 70.90 74.56 85.70
+Kimi-Audio 54.13 63.01 37.80 46.72 56.40 61.17 61.50
+Qwen2.5-Omni 42.18 42.54 28.70 36.16 43.10 39.22 59.70
+Baichuan-Audio 48.27 55.84 31.30 37.60 53.80 57.09 55.30
+Step-Audio 2 mini 49.85 57.57 29.50 41.92 73.10 50.10 45.10
+MiMo-Audio 42.74 51.45 31.20 37.28 30.20 55.15 56.30
+GPT-4o-Audio 82.90 87.40 76.30 69.28 80.60 90.00 92.80
+Qwen3-Omni 65.57 77.11 50.20 49.44 65.90 74.95 75.90
+Fun-Audio-Chat 75.70 91.33 54.10 55.20 75.10 86.60 91.60
 Table 16: English Speech Side Instruction Following Objective Results
 […]
-GLM4-Voice 85.05 92.12 87.17 76.04
-Kimi-Audio 85.66 92.38 89.53 75.06
-Qwen2-Audio-Instruct 85.57 93.25 90,91 72.82
-Qwen2.5-Omni 86.85 93.69 88.81 78.39
-Baichuan-Audio-Chat 73.20 79.91 74.64 65.24
-StepAudio2Mini 87.72 92.58 94.44 76.14
-Qwen3-Omni 91.13 94.26 97.04 82.10
+GLM-4-Voice 82.56 89.73 83.20 74.76
+Kimi-Audio 82.27 85.97 89.28 71.57
+Qwen2.5-Omni 85.23 91.88 86.04 77.76
+Baichuan-Audio 80.96 86.90 86.23 69.76
+Step-Audio 2 mini 82.31 89.77 87.68 69.49
+MiMo-Audio 86.30 90.97 91.35 76.58
+GPT-4o-Audio 33.59 36.50 28.47 35.80
+Qwen3-Omni 87.17 92.43 91.94 77.14
 Table 17: Multi-turn Dialogue Evaluation Results.
 […]
-GLM4-Voice 45.53 41.67 39.33 54.24 58.70 49.60 37.84 66.67 46.00 50.65 62.50 27.87 28.00
-Kimi-Audio 53.51 66.67 52.00 52.54 71.74 49.60 54.05 67.65 50.00 54.55 60.42 36.07 48.00
-Qwen2.5-Omni 55.43 63.89 53.33 52.54 71.74 49.60 43.24 70.59 54.00 53.25 75.00 36.07 53.33
-Baichuan-Audio-Chat 44.48 55.56 44.00 49.15 56.52 40.00 40.54 58.82 41.33 45.45 54.17 36.07 34.67
-Qwen2-Audio-Instruct 35.83 36.11 34.67 40.68 58.70 33.60 37.84 49.02 34.67 32.47 47.92 24.59 24.00
-StepAudio 60.42 55.56 55.33 66.10 73.91 57.60 62.16 61.76 60.67 64.94 70.83 54.10 58.00
-StepAudio2Mini 61.15 61.11 57.33 71.19 80.43 56.45 48.65 74.51 65.33 57.14 68.75 37.70 58.00
-Mimo-Audio 56.58 58.33 43.33 45.76 60.87 60.00 40.54 71.57 65.33 62.34 66.67 34.43 57.33
-GPT4o-Audio 61.29 63.89 62.00 67.80 65.22 59.20 70.27 72.55 53.33 70.13 85.42 62.30 43.33
-Qwen3-Omni 74.74 83.33 72.67 74.58 82.61 63.20 62.16 86.27 80.67 70.13 79.17 60.66 78.00
+GLM-4-Voice 41.79 44.44 38.00 52.54 52.17 44.80 40.54 54.90 43.33 51.95 58.33 24.59 21.33
+Kimi-Audio 54.47 58.33 53.33 59.32 69.57 52.00 45.95 65.69 52.67 55.84 70.83 42.62 45.33
+Qwen2.5-Omni 48.13 47.22 46.67 49.15 65.22 48.00 48.65 64.71 42.00 54.55 52.08 31.15 41.33
+Baichuan-Audio 40.73 38.89 44.00 38.98 50.00 48.00 43.24 49.02 36.67 38.96 47.92 24.59 32.67
+Step-Audio 2 mini 49.86 47.22 47.33 52.54 65.22 54.40 40.54 64.71 49.33 51.95 62.50 24.59 41.33
+MiMo-Audio 48.70 50.00 45.33 38.98 45.65 48.80 35.14 63.73 52.67 54.55 58.33 31.15 46.67
+GPT-4o-Audio 55.81 58.33 58.00 62.71 56.52 58.40 64.86 65.69 50.67 59.74 62.50 55.74 40.00
+Qwen3-Omni 66.86 77.78 66.67 64.41 76.09 64.80 56.76 72.55 71.33 67.53 68.75 55.74 62.00
+Fun-Audio-Chat 46.30 50.00 48.00 49.15 56.52 52.80 51.35 60.78 38.67 45.45 50.00 34.43 34.67
 Table 18: General Knowledge Evaluation Results.
 […]
-GLM4-Voice 62.14 87.67 60.59 32.14 70.31 20.00 63.52
-Kimi-Audio 79.94 98.63 89.41 66.67 85.94 45.00 66.04
-Qwen2.5-Omni 80.39 95.89 78.82 71.43 78.12 47.50 81.76
-Baichuan-Audio-Chat 60.33 69.18 52.94 45.24 70.31 25.00 72.96
-Qwen2-Audio-Instruct 59.28 86.30 57.06 39.29 51.56 22.50 59.75
-StepAudio 77.08 91.10 88.82 75.00 65.62 32.50 68.55
-StepAudio2Mini 81.29 97.26 88.82 65.48 71.88 45.00 79.87
-Mimo-Audio 84.01 94.52 88.82 72.62 78.12 52.50 85.53
-GPT4o-Audio 77.68 82.19 91.18 71.43 75.00 40.00 72.96
-Qwen3-Omni 94.27 97.95 98.82 91.67 90.62 62.50 96.86
+GLM-4-Voice 60.18 86.99 58.82 34.52 67.19 30.00 55.35
+Kimi-Audio 75.42 89.73 79.41 71.43 81.25 37.50 67.30
+Qwen2.5-Omni 73.00 87.67 77.06 61.90 75.00 50.00 66.04
+Baichuan-Audio 74.05 87.67 72.35 66.67 64.06 27.50 83.02
+Step-Audio 2 mini 72.85 93.84 72.94 55.95 68.75 30.00 74.84
+MiMo-Audio 81.75 88.36 87.06 73.81 76.56 47.50 84.91
+GPT-4o-Audio 73.45 76.71 75.88 69.05 84.38 32.50 76.10
+Qwen3-Omni 81.90 85.62 65.29 85.71 90.62 62.50 95.60
+Fun-Audio-Chat 73.00 86.99 78.24 55.95 71.88 27.50 75.47
 Table 19: Mathematical and Logical Reasoning Evaluation Results
 […]
-GLM4-Voice 48.64 58.25 55.75 33.04
-Kimi-Audio 74.76 85.71 83.50 56.48
-Qwen2-Audio-Instruct 67.07 78.64 80.53 43.48
-Qwen2.5-Omni 73.72 83.50 79.65 59.13
-Baichuan-Audio-Chat 54.38 58.25 71.68 33.91
-StepAudio 59.52 63.11 64.60 51.30
-StepAudio2Mini 83.08 92.23 84.07 73.91
-Mimo-Audio 87.92 95.15 88.50 80.87
-GPT4o-Audio 77.64 90.29 84.96 59.13
-Qwen3-Omni 94.26 97.09 96.46 89.57
-Table 16: Discoure Comprehension Evaluation Results.
+Kimi-Audio 51.96 60.19 61.06 35.65
+Qwen2.5-Omni 72.81 83.50 79.65 56.52
+Baichuan-Audio 51.66 67.96 61.06 27.83
+Step-Audio 2 mini 79.46 90.29 84.96 64.35
+GPT-4o-Audio 76.74 89.32 85.84 56.52
+Qwen3-Omni 82.78 89.32 92.04 67.83
+Fun-Audio-Chat 79.15 87.38 76.99 73.91
+Table 20: Discourse Comprehension Evaluation Results.
 […]
 Avg. Child Elder Avg. Tianjin Beijing Dongbei Sichuan Avg. Down Up Avg.
-GLM4-Voice 65.00 (-17.20) 82.98 (-1.28) 47.50 (-32.50) 82.40 (0.20) 85.00 (-0.62) 82.67 (-2.66) 83.20 (3.20) 79.44 (0.00) 96.40 (0.60) 96.80 (0.80) 96.00 (0.40) 39.20 (-55.20)
-Kimi-Audio 45.60 (-30.80) 65.96 (-7.23) 25.83 (-53.74) 56.20 (-20.60) 54.38 (-23.74) 77.33 (18.66) 73.60 (-7.20) 36.67 (-43.89) 93.80 (-1.60) 93.20 (-1.90) 94.40 (-1.20) 87.80 (-4.40)
-Qwen2.5-Omni-7B 75.60 (-5.40) 82.55 (1.27) 68.75 (-12.08) 77.80 (-3.40) 79.38 (-1.87) 81.33 (5.33) 74.40 (-8.00) 77.22 (-5.56) 90.40 (0.20) 93.20 (3.20) 87.60 (-2.80) 89.20 (-1.60)
-Baichuan-Audio-Chat 60.40 (-0.80) 61.70 (-0.43) 59.17 (-1.26) 61.60 (-0.60) 61.88 (-0.62) 62.67 (-5.33) 68.00 (3.20) 56.67 (-1.11) 77.20 (3.00) 83.60 (13.20) 70.80 (-7.20) 60.00 (-7.80)
-Qwen2-Audio-7B-Instruct 76.40 (-5.80) 80.43 (-1.35) 72.50 (-10.00) 81.60 (0.80) 79.38 (0.63) 80.00 (-6.67) 83.20 (0.00) 83.33 (5.00) 93.20 (-1.40) 93.60 (-1.20) 92.80 (-1.60) 77.80 (-7.00)
-StepAudio 62.53 (-18.94) 77.87 (-5.53) 47.50 (-32.08) 71.48 (-3.71) 68.75 (3.75) 76.00 (-13.33) 74.40 (-0.80) 70.00 (-8.33) 95.80 (1.00) 96.00 (1.20) 95.60 (0.80) 38.26 (-54.78)
-StepAudio2 75.60 (-4.80) 75.32 (-2.13) 75.83 (-7.50) 75.40 (-2.00) 78.75 (-0.63) 88.00 (2.67) 67.20 (-4.80) 72.78 (-3.33) 96.60 (2.80) 96.00 (2.00) 97.20 (3.60) 76.00 (-14.40)
-Mimo-Audio 76.00 (-14.80) 92.77 (-0.42) 59.58 (-28.75) 88.40 (-0.20) 90.62 (2.50) 96.00 (10.67) 88.80 (1.60) 82.78 (-8.33) 99.20 (1.00) 98.80 (-0.40) 99.60 (2.40) 60.40 (-33.00)
-GPT4o-Audio 79.00 (-8.60) 88.94 (0.00) 69.17 (-17.08) 88.40 (2.20) 91.88 (4.38) 89.33 (0.00) 85.60 (0.80) 86.67 (2.23) 96.80 (-1.20) 97.20 (0.00) 96.40 (-2.40) 43.40 (-48.40)
-Qwen3-Omni 85.47 (-9.48) 96.60 (1.71) 74.58 (-20.42) 95.19 (1.12) 93.12 (3.74) 98.67 (2.67) 96.80 (0.80) 94.44 (-1.67) 98.80 (-0.80) 98.80 (-0.80) 98.80 (-0.80) 95.22 (-3.91)
+GLM-4-Voice 63.58 (-16.21) 81.28 (2.13) 46.25 (-34.17) 79.26 (0.56) 81.25 (0.63) 78.67 (-5.33) 81.60 (3.20) 76.11 (1.11) 93.00 (-1.20) 91.60 (-2.00) 94.40 (-0.40) 39.13 (-49.13)
+Kimi-Audio 45.89 (-24.43) 62.98 (-5.53) 29.17 (-42.91) 45.74 (-20.00) 35.62 (-30.00) 45.33 (-1.34) 65.60 (-4.00) 41.11 (-30.00) 92.00 (-3.00) 92.80 (-3.60) 91.20 (-2.40) 86.09 (-3.91)
+Qwen2.5-Omni 74.53 (1.06) 75.32 (2.55) 73.75 (-0.42) 75.74 (-0.56) 77.50 (-0.62) 80.00 (0.00) 76.80 (4.80) 71.67 (-4.44) 82.60 (-2.20) 82.80 (-1.60) 82.40 (-2.80) 79.13 (0.00)
+Baichuan-Audio 71.58 (-3.16) 74.89 (-3.83) 68.33 (-2.50) 77.96 (4.81) 75.62 (0.00) 84.00 (8.00) 80.00 (7.20) 76.11 (6.11) 90.40 (-2.20) 89.20 (-3.20) 91.60 (-1.20) 64.78 (-16.96)
+Step-Audio 2 mini 69.26 (-5.69) 69.36 (-5.96) 69.17 (-5.41) 69.63 (-1.85) 75.62 (-1.26) 77.33 (12.00) 61.60 (-6.40) 66.67 (-5.00) 89.60 (0.40) 90.40 (1.20) 88.80 (-0.40) 69.13 (-13.91)
+MiMo-Audio 85.26 (-1.90) 87.23 (0.00) 83.33 (-3.75) 84.07 (-1.12) 78.75 (-6.87) 89.33 (2.66) 85.60 (0.00) 85.56 (1.67) 97.20 (-2.00) 96.00 (-3.60) 98.40 (-0.40) 50.43 (-41.31)
+GPT-4o-Audio 77.89 (-5.69) 87.23 (4.25) 68.75 (-15.42) 85.74 (1.85) 85.00 (-1.25) 86.67 (6.67) 88.00 (2.40) 84.44 (2.22) 92.80 (-2.00) 92.80 (-1.20) 92.80 (-2.80) 45.65 (-46.96)
+Qwen3-Omni 77.05 (-12.00) 89.36 (-0.85) 65.00 (-22.92) 87.59 (0.18) 84.38 (3.76) 92.00 (4.00) 90.40 (-3.20) 86.67 (-2.22) 95.80 (0.20) 94.80 (-1.60) 96.80 (2.00) 93.91 (-3.05)
+Fun-Audio-Chat 86.11 (-0.63) 89.36 (2.13) 82.92 (-3.33) 86.85 (-0.37) 88.12 (1.24) 90.67 (0.00) 86.40 (-2.40) 84.44 (-0.56) 94.00 (-0.41) 94.00 (-0.47) 94.00 (-0.35) 86.09 (-9.13)
 Table 21: Speaker Variations Evaluation Results - Experimental Group (Difference from Control Group).
@@
 Avg. Echo Outdoors Far Field Avg. Tv Playback Background Chat Vocal-Music Voice Announce Avg.
-GLM4-Voice 64.00 (-22.60) 45.00 (-36.75) 96.57 (-0.57) 74.86 (-12.57) 89.20 (-0.80) 85.56 (-3.89) 82.94 (-1.47) 93.03 (-0.68) 92.82 (1.57) 78.40 (-14.60)
-Kimi-Audio 74.00 (-5.40) 60.50 (-11.65) 94.86 (4.57) 84.00 (-1.14) 75.80 (-9.60) 76.73 (-8.36) 58.24 (-11.47) 93.71 (0.45) 70.25 (-19.75) 76.80 (-11.20)
-Qwen2.5-Omni-7B 69.40 (-16.80) 52.00 (-30.50) 95.43 (0.57) 82.86 (-2.85) 85.20 (-4.80) 80.74 (-6.17) 74.71 (-8.23) 95.51 (-1.79) 85.57 (-4.43) 79.80 (-10.20)
-Baichuan-Audio-Chat 56.60 (-15.60) 45.00 (-20.00) 81.14 (-8.57) 58.29 (-12.57) 73.40 (-2.80) 73.45 (0.36) 64.41 (0.88) 81.35 (-5.39) 72.00 (-5.50) 67.20 (-4.60)
-Qwen2-Audio-7B-Instruct 59.80 (-17.60) 37.75 (-32.00) 93.71 (-1.72) 76.00 (-1.14) 87.60 (-1.60) 82.18 (-6.18) 80.00 (-1.49) 96.18 (-0.18) 88.50 (0.25) 77.80 (-13.40)
-StepAudio 55.87 (-30.80) 36.00 (-46.25) 93.71 (-0.58) 63.43 (-25.71) 86.92 (-2.94) 83.27 (-10.18) 77.65 (-1.47) 97.75 (0.45) 85.25 (-3.00) 66.50 (-27.25)
-StepAudio2 65.00 (-19.20) 45.75 (-33.00) 94.86 (-1.14) 78.86 (-6.28) 83.40 (-5.80) 84.00 (-2.18) 66.18 (-12.06) 94.83 (-2.47) 84.75 (-6.75) 81.60 (-8.40)
-Mimo-Audio 73.20 (-20.20) 54.00 (-35.50) 99.43 (4.43) 90.86 (-4.57) 90.40 (-5.00) 86.91 (-9.09) 86.76 (-5.00) 99.55 (0.45) 85.50 (-8.25) 95.20 (-4.20)
-GPT4o-Audio 68.20 (-19.20) 61.00 (-20.50) 98.29 (-0.57) 54.86 (-34.28) 92.00 (-1.60) 86.18 (-6.55) 87.94 (-0.30) 96.40 (0.22) 94.25 (-1.50) 74.60 (-21.60)
-Qwen3-Omni 83.47 (-13.86) 71.25 (-25.50) 97.71 (-1.72) 97.14 (0.57) 96.44 (-1.85) 96.73 (-1.45) 94.71 (-2.94) 100.00 (0.00) 93.75 (-3.25) 96.25 (-3.75)
+GLM-4-Voice 61.73 (-20.14) 43.75 (-35.25) 94.86 (1.72) 69.71 (-7.43) 85.96 (-1.99) 80.36 (-5.82) 78.53 (-4.12) 91.69 (0.90) 89.75 (-0.75) 76.25 (-7.75)
+Kimi-Audio 64.27 (-11.73) 47.75 (-22.00) 92.00 (3.43) 74.29 (-3.42) 66.51 (-12.46) 71.27 (-8.73) 37.35 (-24.41) 85.17 (-0.67) 67.25 (-18.00) 61.75 (-20.75)
+Qwen2.5-Omni 60.80 (-8.93) 47.75 (-18.25) 81.14 (0.57) 70.29 (2.86) 80.14 (0.48) 81.09 (-1.46) 73.24 (-0.29) 82.92 (1.57) 82.25 (1.25) 79.00 (-7.75)
+Baichuan-Audio 64.93 (-14.00) 48.50 (-24.75) 94.29 (2.29) 73.14 (-5.72) 85.48 (-0.89) 78.91 (-5.09) 78.24 (1.77) 94.61 (-1.12) 86.00 (0.00) 72.75 (-12.75)
+Step-Audio 2 mini 55.47 (-22.66) 38.00 (-38.00) 81.14 (-4.57) 69.71 (-5.72) 77.33 (-3.56) 76.00 (-2.91) 66.76 (-8.24) 85.62 (0.23) 78.00 (-4.25) 74.25 (-3.25)
+MiMo-Audio 73.07 (-15.60) 57.00 (-28.25) 99.43 (-0.57) 83.43 (-1.71) 91.78 (-1.78) 89.45 (-3.64) 85.29 (-5.00) 97.53 (0.00) 92.50 (0.25) 85.00 (-11.75)
+GPT-4o-Audio 67.07 (-18.80) 60.75 (-24.00) 94.86 (2.29) 53.71 (-28.00) 88.56 (-1.51) 79.64 (-8.72) 86.76 (1.17) 93.03 (2.02) 91.25 (-2.75) 72.50 (-17.75)
+Qwen3-Omni 78.53 (-14.67) 67.25 (-27.00) 90.86 (-4.00) 92.00 (2.86) 89.66 (-3.29) 84.73 (-2.54) 82.06 (-7.65) 99.10 (0.00) 89.00 (-3.75) 84.25 (-7.00)
+Fun-Audio-Chat 70.13 (-20.94) 53.25 (-37.25) 96.00 (0.00) 82.86 (-4.57) 91.03 (-1.50) 82.91 (-4.36) 87.35 (-2.36) 96.40 (-1.35) 93.75 (1.50) 83.25 (-9.00)
 Table 22: Environmental Variations Evaluation Results - Experimental Group (Difference from Control Group).
@@
 Avg. Casual Talk Mispronunciation Grammatical Error Topic Shift Code Switching
-GLM4-Voice 84.41 (-6.90) 85.81 (-4.04) 82.20 (-10.60) 85.60 (-6.00) 86.20 (-6.40) 80.80 (-10.80)
-Kimi-Audio 77.71 (-9.40) 81.18 (-4.44) 61.80 (-23.20) 77.40 (-10.40) 79.40 (-6.80) 84.00 (-8.80)
-Qwen2.5-Omni-7B 80.44 (-6.75) 83.25 (-2.17) 80.60 (-8.40) 79.20 (-8.60) 75.00 (-11.40) 80.40 (-9.20)
-Baichuan-Audio-Chat 67.46 (-7.02) 68.57 (-5.52) 71.40 (-1.20) 65.60 (-11.00) 61.60 (-12.40) 68.40 (-7.60)
-Qwen2-Audio-7B-Instruct 79.60 (-9.04) 87.29 (-2.56) 78.40 (-10.00) 79.40 (-9.20) 67.00 (-19.80) 76.40 (-11.60)
-StepAudio 85.81 (-5.64) 86.21 (-3.35) 89.44 (-2.69) 90.43 (-2.32) 82.42 (-7.78) 81.30 (-13.90)
-StepAudio2 83.04 (-4.79) 85.02 (-2.27) 82.20 (-5.60) 80.80 (-8.20) 80.60 (-3.00) 83.60 (-8.60)
-Mimo-Audio 95.10 (-3.04) 96.45 (-0.69) 96.00 (-3.60) 97.40 (-0.60) 92.00 (-5.60) 92.60 (-7.00)
-GPT4o-Audio 92.34 (-1.94) 92.12 (-0.29) 92.60 (-4.00) 89.00 (-5.00) 90.60 (-2.60) 96.80 (-0.60)
-Qwen3-Omni 97.61 (-1.62) 98.92 (-0.19) 94.61 (-4.04) 95.36 (-3.77) 96.92 (-2.42) 100.00 (0.00)
+GLM-4-Voice 78.60 (-7.91) 80.79 (-5.42) 78.43 (-9.21) 82.90 (-4.06) 78.02 (-6.38) 71.30 (-16.53)
+Kimi-Audio 71.51 (-10.04) 70.15 (-8.77) 65.84 (-16.86) 71.01 (-10.73) 74.07 (-7.91) 77.83 (-7.82)
+Qwen2.5-Omni 78.27 (-2.83) 81.38 (0.20) 74.61 (-6.06) 78.55 (-0.58) 74.07 (-7.47) 78.91 (-3.48)
+Baichuan-Audio 78.71 (-4.64) 83.74 (0.19) 73.48 (-11.01) 80.00 (-1.45) 72.09 (-7.25) 78.26 (-8.91)
+Step-Audio 2 mini 74.78 (-4.78) 76.45 (-1.48) 73.48 (-6.30) 75.36 (-5.22) 70.11 (-7.47) 76.52 (-7.61)
+MiMo-Audio 89.38 (-4.30) 92.12 (-1.48) 89.89 (-2.92) 91.01 (-2.61) 84.18 (-9.01) 86.74 (-8.48)
+GPT-4o-Audio 86.51 (-1.98) 87.88 (-0.30) 85.84 (-3.37) 80.00 (-6.09) 84.18 (-1.97) 91.30 (-1.31)
+Qwen3-Omni 86.03 (-2.32) 91.13 (0.39) 82.70 (-3.37) 83.77 (-6.09) 83.30 (-4.61) 82.39 (-2.18)
+Fun-Audio-Chat 85.15 (-4.63) 88.87 (-1.08) 83.15 (-4.94) 84.06 (-4.35) 82.42 (-6.59) 82.39 (-10.44)
 Table 23: Content Variations Evaluation Results - Experimental Group (Difference from Control Group).
-[…]
-GLM4-Voice 79.2 (83.2) 80.8 (88.4) 78.8 (81.8) 88.6 (92.4) 84.0 (88.4) 75.4 (75.6) 88.2 (92.6)
-Kimi-Audio 75.4 (76.8) 81.4 (86.6) 64.2 (64.8) 81.0 (84.8) 83.2 (88.4) 75.8 (76.6) 82.0 (83.4)
-Qwen2.5-Onmi 71.4 (72.4) 75.4 (80.4) 74.2 (75.8) 84.8 (86.2) 72.8 (79.0) 80.0 (80.8) 77.8 (79.2)
-Baichuan-Audio-Chat 83.2 (81.8) 86.0 (89.2) 73.6 (71.4) 85.6 (90.2) 75.2 (79.4) 78.6 (78.8) 92.6 (93.4)
-StepAudio2Mini 77.0 (82.2) 77.4 (86.6) 70.8 (73.8) 79.4 (88.0) 81.8 (87.8) 62.6 (63.8) 83.8 (95.2)
-GPT4o-Audio 86.6 (89.2) 85.4 (93.4) 83.2 (85.6) 91.2 (96.6) 87.4 (93.4) 83.6 (85.8) 95.2 (98.2)
-Table 20: A2A Result in TIF - A2A W/ ASR (A2A W/O ASR)
-[…]
-GLM4-Voice 65.0 (81.6) 75.4 (83.8) 72.6 (77.4) 78.0 (85.4) 77.8 (83.8) 69.2 (76.6) 86.0 (89.8)
-Kimi-Audio 50.4 (62.0) 60.4 (83.4) 53.0 (56.0) 62.4 (80.8) 65.6 (79.0) 78.4 (83.4) 58.4 (69.8)
-Qwen2.5-Omni 57.0 (59.6) 58.2 (64.2) 56.4 (60.2) 66.4 (72.4) 56.2 (66.6) 73.4 (75.6) 56.2 (63.0)
-Baichuan-Audio-Chat 79.8 (77.6) 80.0 (81.0) 75.8 (75.0) 77.0 (76.8) 79.8 (81.2) 72.0 (72.0) 86.2 (86.0)
-StepAudio2Mini 66.2 (67.0) 73.0 (78.0) 55.2 (56.2) 75.0 (77.0) 70.6 (73.0) 69.6 (70.4) 84.2 (87.8)
-GPT4o-Audio 90.2 (91.4) 90.0 (94.8) 85.0 (85.8) 93.4 (95.8) 92.4 (94.6) 80.4 (81.6) 96.0 (98.4)
-Table 21: A2A Result in TIF-En - A2A W/ ASR (A2A W/O ASR)

### v4

@@
 Qwen3-Omni (Xu et al., 2025b) 90.45 85.17 70.73 65.57 87.17 66.86 81.90 82.78 87.91 85.63 86.03
-Fun-Audio-Chat (Chen et al., 2025) 89.30 89.39 78.82 75.70 / 46.30 73.00 79.15 88.60 83.83 85.15
+Fun-Audio-Chat (Chen et al., 2025) 89.30 89.39 78.82 75.70 85.27 53.89 86.12 87.31 88.60 83.83 85.15
 Table 1: The main results of different LALMs on VCB Bench. Missing results from unsupported modalities or API unavailability.
@@
 Table 1 presents the average scores of each model across various test sets in VCB Bench, with the detailed scores for all subsets provided in the Appendix.
-As shown, Qwen3-Omni emerges as a new state-of-the-art (SOTA) model, demonstrating all-round superiority across most tasks.
+As shown, Qwen3-Omni and Fun-Audio-Chat emerge as the new state-of-the-art (SOTA) models, demonstrating all-round superiority across most tasks.
 In Instruction Following, Qwen3-Omni (90.45) and MiMo-Audio (90.08) excel in TIF, while Fun-Audio-Chat leads in TIF-En (89.39), closely trailing GPT-4o-Audio (88.80).
@@
 In MTD, Qwen3-Omni (87.17) and MiMo-Audio (86.30) lead, while GPT-4o-Audio performs poorly (33.59), highlighting divergence in long-context dialogue control.
-In Knowledge, Qwen3-Omni stands out with top scores in GK (66.86), ML (81.90) and DC (82.78), showing superior multi-disciplinary knowledge and reasoning.
-MiMo-Audio performs strongly in ML (81.75), comparable to Qwen3-Omni.
+In Knowledge, Qwen3-Omni stands out with top scores in GK (66.86) and Fun-Audio-Chat performs strongly in ML (86.12) and DC (87.31), showing superior multi-disciplinary knowledge and reasoning.
 However, Baichuan-Audio (GK:40.73, ML:74.05), GLM-4-Voice (GK:41.79, ML:60.18) and other models lag, revealing gaps in knowledge coverage and step-by-step reasoning.
-Overall, Qwen3-Omni demonstrates comprehensive superiority.
-Other models have task-specific strengths: Fun-Audio-Chat in cross-lingual TIF and SIF, MiMo-Audio in TIF and ML.
+Overall, Qwen3-Omni and Fun-Audio-Chat demonstrate comprehensive superiority, while other models have task-specific strengths.
 Most open-source E2E LALMs face English adaptation challenges, while GPT-4o-Audio lacks long-context dialogue capability.
@@
 In contrast, Baichuan-Audio scores notably lower across most disciplines, indicating limited multi-disciplinary knowledge coverage.
-For Mathematical and Logical Reasoning, Qwen3-Omni (81.90) and MiMo-Audio (81.75) stand out with the highest averages.
-Qwen3-Omni dominates in Logic (95.60) and Induction (90.62), while MiMo-Audio tops Medium Math (87.06).
-However, Models like GLM-4-Voice (60.18) and Fun-Audio-Chat (73.00) show weaker capabilities in reasoning sub-tasks (e.g., Analysis, Analogy).
-For Discourse Comprehension, Qwen3-Omni (82.78) achieves the highest average, excelling in Induction (92.04), Inference (89.32) and Analysis (67.83).
-Step-Audio 2 mini (79.46) and Fun-Audio-Chat (79.15) also perform well, while Kimi-Audio (51.96) and Baichuan-Audio (51.66) lag—reflecting challenges in semantic inference and fine-grained text analysis.
-Overall, Qwen3-Omni demonstrates robust reasoning and comprehension capabilities, while GPT-4o-Audio excels in knowledge breadth but shows only moderate performance in mathematical reasoning.
+For Mathematical and Logical Reasoning, Fun-Audio-Chat (86.12) and Qwen3-Omni (81.90) stand out with the highest averages.
+Qwen3-Omni dominates in logical reasoning, while MiMo-Audio tops math.
+However, models like GLM-4-Voice (60.18) shows weaker capabilities in reasoning sub-tasks (e.g., Analysis, Analogy).
+For Discourse Comprehension, Fun-Audio-Chat (87.31) achieves the highest average, excelling in Induction (88.50), Inference (95.15) and Analysis (79.13).
+Qwen3-Omni (82.78) and Step-Audio 2 mini (79.46) also perform well, while Kimi-Audio (51.96) and Baichuan-Audio (51.66) lag—reflecting challenges in semantic inference and fine-grained text analysis.
+Overall, Qwen3-Omni and Fun-Audio-Chat demonstrate robust reasoning and comprehension capabilities, while GPT-4o-Audio excels in knowledge breadth but shows only moderate performance in mathematical reasoning.
 Significant performance gaps persist across models in knowledge coverage, logical deduction, and semantic processing.
@@
 Qwen3-Omni 87.17 92.43 91.94 77.14
+Fun-Audio-Chat 85.27 91.14 90.18 74.50
 Table 17: Multi-turn Dialogue Evaluation Results.
@@
 Qwen3-Omni 66.86 77.78 66.67 64.41 76.09 64.80 56.76 72.55 71.33 67.53 68.75 55.74 62.00
-Fun-Audio-Chat 46.30 50.00 48.00 49.15 56.52 52.80 51.35 60.78 38.67 45.45 50.00 34.43 34.67
+Fun-Audio-Chat 53.89 61.11 54.00 47.46 58.70 51.20 45.95 64.71 55.33 63.64 56.25 37.70 49.33
 Table 18: General Knowledge Evaluation Results.
@@
 Qwen3-Omni 81.90 85.62 65.29 85.71 90.62 62.50 95.60
-Fun-Audio-Chat 73.00 86.99 78.24 55.95 71.88 27.50 75.47
+Fun-Audio-Chat 86.12 97.26 88.24 66.67 84.38 62.50 90.57
 Table 19: Mathematical and Logical Reasoning Evaluation Results
@@
 Qwen3-Omni 82.78 89.32 92.04 67.83
-Fun-Audio-Chat 79.15 87.38 76.99 73.91
+Fun-Audio-Chat 87.31 95.15 88.50 79.13
 Table 20: Discourse Comprehension Evaluation Results.